### v1

# Leveraging data-driven weather models for improving numerical weather prediction skill through large-scale spectral nudging

Abstract
Operational meteorological forecasting has long relied on physics-based numerical weather prediction (NWP) models. Recently, this landscape has been disrupted by the advent of data-driven artificial intelligence (AI)-based weather models, which offer tremendous computational performance and competitive forecasting skill. However, data-driven models for medium-range forecasting generally suffer from major limitations, including low effective resolution and a narrow range of predicted variables. This study illustrates the relative strengths and weaknesses of these competing paradigms using the GEM (Global Environmental Multiscale) and GraphCast models to represent physics-based and AI-based approaches, respectively. By analyzing global predictions from these two models against observations and analyses in both physical and spectral spaces, this study demonstrates that GraphCast-predicted large scales outperform GEM, particularly for longer lead times. Building on this insight, a hybrid NWP-AI system is proposed, wherein GEM-predicted large-scale state variables are spectrally nudged toward GraphCast predictions, while allowing GEM to freely generate fine-scale details critical for weather extremes. Results indicate that this hybrid approach is capable of leveraging the strengths of GraphCast to enhance the prediction skill of the GEM model. Importantly, trajectories of tropical cyclones are predicted with enhanced accuracy without significant changes in intensity. Furthermore, this new hybrid system ensures that meteorologists have access to a complete set of forecast variables, including those relevant for high-impact weather events.
1 Introduction
State-of-the-art numerical weather prediction (NWP) models generally include some form of a dynamical core that solves the nonlinear partial differential equations governing the resolved-scale dynamical and thermodynamical processes within the atmosphere. Solutions from the dynamical core are subsequently combined with tendencies from parameterized physics schemes, representing the unresolved and subgrid-scale processes that are not explicitly accounted for by the dynamical equations. Although statistical approaches have been explored in the past — particularly for downscaling purposes (Yu et al., 2006; Busuioc et al., 2008; Cheng et al., 2014) — physics-based NWP models have long been the foundational approach for operational meteorological forecasting. The recent emergence of data-driven models inspired by artificial intelligence (AI) has, however, started to seriously challenge this well established paradigm.
Keisler (2022) presented an important breakthrough demonstrating considerable potential for weather forecasting with data-driven models. This was quickly followed by Pangu-Weather from Huawei (Bi et al., 2022, 2023), GraphCast from Google-DeepMind (Lam et al., 2022, 2023) and several other models (e.g., Pathak et al. 2022; Chen et al. 2023a, b). In general, these models rely on some form of deep neural network architecture. Like any other application of AI, data-driven weather models necessitate a substantial volume of high quality training data. As a result, all currently available AI-based deterministic global weather simulators are trained on the ERA5 reanalyses (Hersbach et al., 2020) from the European Centre for Medium-Range Weather Forecasts (ECMWF), which is undoubtedly the most comprehensive resource available. The neural network weights within these models are specifically trained to make predictions that closely emulate ERA5.
One important aspect of these AI-based weather emulators that set them apart from the classical NWP models is their computational performance. They can speed up the production of multi-day forecasts by orders of magnitude while using a fraction of the computational resources. Furthermore, many of these models are not only comparable to the NWP models with respect to their prediction accuracy, in terms of a number of the performance metrics, they can even surpass the skill of the IFS (Integrated Forecasting System; Lang et al. 2023) from ECMWF, which is widely accepted as the benchmark within the NWP community (Bi et al., 2023; Lam et al., 2023; Chen et al., 2023a).
Despite their advantages, AI-based weather emulators also have limitations. A major weakness of most of the currently available data-driven models is considerable fine-scale smoothing (Bi et al., 2022; Lam et al., 2022). Notably, data-driven models based on generative AI, e.g., diffusion-based GenCast model from Google-DeepMind (Price et al., 2024), can address the fine-scale smoothing problem to some extent. However, as of now, such approaches generally require considerable increase in computational cost, thereby diminishing one of the principal advantages of the AI models.
Improving the nominal horizontal resolution of the AI model inferences (i.e., forecasts) is also challenging. The general construction of these models requires higher-resolution training data to enable these models to make inferences at higher resolutions. Since the ERA5 reanalysis is only available at 0.25°$\degree$ resolution, this restriction means that current global AI models cannot be trained for kilometer-scale forecasting. Another common limitation of the current generation of AI models is the limited range of predicted variables. At present, these models can only be trained to emulate fields that are present in the analyses dataset, although training to predict non-analyzed variables (e.g., precipitation in ERA5) would generally lead to suboptimal inference. Furthermore, adding the full suite of physics-related prognostic variables required for operational NWP, such as categorization of cloud and precipitation types, would be cost-prohibitive with current AI model architectures. Time resolution of inferences is another constraint for most of the current AI models, which make predictions at 6-hour intervals, far coarser than physics-based NWP models. Pangu-Weather (Bi et al., 2022) provides separate models for 1-, 3-, 6-, and 24-hour predictions, but the models with shorter lead times suffer from larger error accumulation. In summary, despite their strengths, the aforementioned limitations make it implausible for the AI models to replace the classical NWP approach in operational weather forecasting.
The GEM (Global Environmental Multiscale) model (Girard et al., 2014), which is employed by Environment and Climate Change Canada (ECCC) for operational weather forecasting, relies entirely on the physics-based NWP approach. The principal objective of this study is to determine whether inferences from a state-of-the-art AI-based weather emulator can be leveraged through large-scale spectral nudging to enhance the predictability of the GEM model, particularly for global weather forecasting. Spectral nudging can be seen as an indirect suboptimal data assimilation method (von Storch et al., 2000). Although the concept of spectral nudging was first proposed to control spatial computational modes in a limited-area model (LAM; Waldron et al. 1996), it is primarily used to improve dynamical downscaling of coarse-resolution atmospheric model simulations over high-resolution limited area domains (Leduc and Laprise, 2009; Husain et al., 2014). Spectral nudging has also been applied for global atmospheric hindcasting (Schubert-Frisius et al., 2017).
The ultimate goal of this study is to develop a hybrid NWP-AI system that combines the strengths of both approaches while addressing the inherent limitations of the AI-based models. The AI model chosen for this study is the GraphCast model from Google-DeepMind. Based on internal testing at ECCC, among the different freely available open-source AI-based models, GraphCast demonstrates the best accuracy with respect to most of the metrics.
To achieve the most optimal fusion of GraphCast and GEM, it is essential to ascertain the relative strengths and limitations of these models through objective evaluation against observations and analyses. A detailed comparative analysis of these models is also critical for devising the most optimal spectral nudging configuration, and therefore, constitutes a significant component of this study.
Spectral nudging has so far been unfeasible for real-time forecasting applications due to the lack of availability of timely and accurate reference data. The emergence of AI models like GraphCast has unlocked the potential for applying spectral nudging in real-time operational weather forecasting. In this regard, application of spectral nudging for enhancing real-time NWP represents an important novelty proposed by this study. Although this paper illustrates the hybridization concept with a specific combination of the AI and NWP models, the proposed strategy is expected to be broadly applicable to any other combination as long as the AI model offers some added value over the NWP model.
This article is organized as follows. Sections 2 and 3 provide some important background information regarding the GEM and GraphCast models, respectively. Details pertaining to the relative performances of these models are documented in section 4. Section 5 covers the concept of spectral nudging, its implementation in GEM as well as the optimal nudging configuration used in this study. Detailed evaluations of the spectral nudging-based hybrid NWP-AI system are presented in 6. Finally, section 7 summarizes the conclusions of this study and outlines potential directions for future work.
2 The NWP System
2.1 The GEM Model
The dynamical core of the GEM model solves the elastic Euler system of equations. These equations are first transformed from regular height coordinate to some form of a terrain-following coordinate (TFC) in the vertical, denoted by $\zeta$ in the model. All current operational GEM-based NWP systems employ a log-hydrostatic-pressure-type TFC (Girard et al., 2014; Husain et al., 2021).
The nonhydrostatic version of the dynamical core involves eight prognostic variables: horizontal velocity components ($u$ and $v$), vertical velocity ($w$), virtual temperature ($T_{v}$), geopotential ($\phi$), vertical motion with respect to the TFC ($\dot{\zeta}$), hydrostatic surface pressure ($\pi_{s}$), and nonhydrostatic pressure perturbation ($q=\ln\frac{p}{\pi}$, where $p$ and $\pi$ denote the complete and hydrostatic pressure, respectively).
The advection terms in the model are treated with semi-Lagrangian approach while a two-time-level iterative implicit Crank-Nicolson type scheme is employed for temporal discretization. Spatial discretization relies on staggered grids: an Arakawa C grid in the horizontal and a Charney-Phillips grid in the vertical. Further details in this regard can be found in the existing literature (Girard et al., 2014; Husain and Girard, 2017). Global simulations in GEM utilize a Yin-Yang grid system, where two orthogonal and overlapping LAM grids are combined to construct the global grid (Qaddouri and Lee, 2011).
Solutions from the GEM dynamical core are combined with tendencies from parameterized physics schemes that represent unresolved (or partially resolved) physical processes, such as radiation, convection, boundary layer turbulence, surface-atmosphere interactions, condensation, etc. The parameterization schemes are executed in a sequential manner. All operational GEM-based NWP systems rely on a splitting-type mechanism for coupling the parameterized physics tendencies with the resolved dynamics. In this approach, physics-induced tendencies are applied as grid-point adjustments to the dynamics solution to obtain the complete solution for a model time step (Husain et al., 2019).
2.2 Global Deterministic Prediction System (GDPS)
The version of the GDPS used in this study is based on its current operational configuration. It has a horizontal grid spacing of approximately 15 km. In the vertical, GDPS employs 84 prognostic vertical levels with the topmost level at approximately 0.1 hPa. The first momentum and thermodynamical levels are approximately 20 m and 10 m away from the surface, respectively. It uses a time-step length of 450 s.
In terms of the underlying physics parameterizations, GDPS has undergone a major modernization in the recent past (McTaggart-Cowan et al., 2019). This includes updated versions of schemes for radiative transfer, blocking, orography gravity waves and non-orographic gravity waves.
Although operational GDPS is coupled with the NEMO ocean model (Smith et al., 2018), all the experiments for this study were conducted with atmosphere-only configurations. As the impact of coupling is usually modest for 10-day forecasts (McTaggart-Cowan et al., 2019), conclusions drawn from this study are expected to generally hold in the presence of atmosphere-ocean coupling.
3 The AI-based Prediction System
3.1 The GraphCast Model
The GraphCast (GC) model from Google-DeepMind is an AI-based weather emulator which has been trained to emulate ECMWF’s ERA5 reanalyses. It requires the state of the atmosphere at times $t$ and $t-\Delta t_{GC}$ to produce a forecast valid at $t+\Delta t_{GC}$, where $\Delta t_{GC}$=6 hr. Although the model can be employed auto-regressively $N$ times to generate predictions valid at $t+N\Delta t_{GC}$, it is trained to minimize error for up to 12 recursive forecast steps (from 6 hr to 3 days) by considering ERA5 as the ground truth. The objective of such a training is to produce forecasts with reasonable skill for up to day 10.
At its core, GC employs graph neural networks (GNNs) with an “encoder-processor-decoder” configuration (Lam et al., 2022, 2023). The input and output states of GC are represented over a 0.25°$\degree$ latitude-longitude global grid. A major advantage of GNNs is the possibility of having arbitrary range of spatial interactions. GraphCast takes advantage of this GNN feature by introducing a multi-mesh architecture based on iteratively-refined icosahedral grids (from level-0 to level-6 refinements) within the processor component. The encoder of GC maps data from the input 0.25°$\degree$ latitude-longitude grid to the internal multi-mesh of the processor, whereas the decoder brings back information from the processor multi-mesh to the output grid.
The internal multi-mesh of GC processor is spatially homogeneous. The highest resolution (level-6) mesh has a nominal resolution of 1°$\degree$ with 40,962 nodes, while the nodes of the coarse-resolution meshes are subsets of these nodes. This facilitates multi-scale graph edges that connect the highest-resolution mesh nodes to the nodes associated with the coarse-resolution meshes to achieve multi-mesh message passing. Such an approach, therefore, leads to multi-scale interactions ranging from synoptic to mesoscales.
The atmospheric state predicted by GC is represented by 6 variables defined on multiple pressure levels (temperature, $u-v$ components of wind, geopotential, specific humidity, and vertical wind speed) and 5 surface variables (2-m temperature, 10-m wind components, mean sea-level pressure, and total precipitation). In addition to the predicted variables, the input system of GC requires time-dependent forcing in the form of total incident solar radiation at the top of the atmosphere along with information regarding the local time of the day and progression of year. These time-dependent forcing terms are computed analytically. Furthermore, the input system requires some stationary input features like the land-sea mask, surface geopotential as well as information on the local longitude and latitude of its input grid.
Pre-trained weights of GC (learned features of the GNNs) are available from Google-DeepMind for two different sets of pressure levels. The version with higher vertical resolution has 37 pressure levels with the top at 1 hPa, while the coarser-resolution version has only 13 levels with the top at 50 hPa (Lam et al., 2022). It is important to note that, in addition to training on ERA5, the 13-level GC was also fine-tuned by DeepMind on initial conditions for ECMWF’s operational IFS system, from 2016 to 2021.
3.2 GraphCast-based Global Prediction System
An experimental GC-based global prediction system has been developed at ECCC. Both the 37-level and 13-level versions of GC are available. This system takes the GDPS operational analyses (Caron and Buehner, 2022) as its input. However, the GNN weights have not been subjected to any fine-tuning at ECCC. All GC predictions utilized in this study are based on this experimental system.
Based on the evaluation of GC predictions with the GDPS analyses as input, the 13-level version of GC (GC-13L) outperforms the 37-level version (GC-37L) for most variables, at most levels and over most regions of the world. This is presumably due to the fact that stratospheric specific humidity in the GDPS analyses generally differs from ERA5 by an order of magnitude creating large imbalances when 37-level GC (with multiple levels in the stratosphere) is initialized with the GDPS analyses. The previously mentioned additional fine-tuning of GC-13L by DeepMind could also have contributed to its superior performance. This study is, therefore, entirely based on the GC-13L and, henceforth, GC is used to refer to GC-13L.
4 Relative performances of the GDPS and GC forecasts
4.1 Experimental setup
Ten-day forecasts from GC and GDPS are initialized using the GDPS analysis every 36 h, leading to 60 cases for boreal winter 2022 (from 01 December 2021 at $0$ UTC to 27 February 2022 at 1200 UTC) and 62 cases for boreal summer 2022 (from 01 June 2022 at $0$ UTC to 31 August 2022 at 1200 UTC). This version of GDPS, which includes data assimilation and forecasting components, became operational in June 2024 (ECCC, 2024). Henceforth, the GDPS experiment based on this configuration will be referred to as GDPS-CTL.
4.2 Verification with global power spectra
Although GC employs a 0.25°$\degree$ latitude-longitude resolution grid, its predictions are subject to a smoothing effect, which increases with lead time. This is generally assumed to be associated with learning to minimize the mean square error (MSE) (e.g., Keisler (2022), Lam et al. (2023)). The smoothing effect is well indicated by departures of 120-h GC forecasts’ global 250-hPa kinetic energy and 500-hPa temperature variance spectra from those of GDPS-CTL forecasts and the analyses (Fig. 1). GraphCast’s spectra are steeper, leading to a notable variance deficiency at wavenumbers as low as 30, corresponding to lower synoptic and sub-synoptic scales ($<$1500 km). However, both GDPS-CTL forecasts and analyses much more closely follow the $k^{-3}$-law, pertaining to the 3-dimensional quasi-geostrophic flow (Charney, 1971; Skamarock et al., 2014), which indicates that GC slopes are not realistic. Hence, the effective resolution of GC, defined as the highest wave number at which the atmospheric variability is still adequately represented, is considerably lower than its nominal resolution. Conversely, the effective resolution of the classical NWP models is typically about 6$\sim$8 times their nominal resolution, i.e., grid spacing (Skamarock, 2004). With the current operational configuration, the nominal resolution for GDPS-CTL is 0.1375°$\degree$, leading to an effective resolution of around 90-100 km.
Figure 1: Global (a) kinetic energy power spectrum of 250-hPa horizontal wind and (b) variance spectrum of 500-hPa temperature after 120 h of integration for the GDPS analyses (black), GDPS-CTL (red) and GC (red), averaged over 60 cases of boreal winter 2022. The dashed line indicates the -3 slope.
It is important to note that the stationary effects of orography, land-sea contrast and large latitudinal climatological variations tend to be much less affected by smoothing in GC than the less predictable transient anomalies. It is therefore useful to introduce the transient spectral amplitude ratio, defined as follows:
$\gamma=\left(\frac{\overline{\sigma_{x^{\prime}x^{\prime}}}}{\overline{\sigma_{y^{\prime}y^{\prime}}}}\right)^{1/2},$ (1)
where $\sigma_{x^{\prime}x^{\prime}}$ and $\sigma_{y^{\prime}y^{\prime}}$ are the power spectra of the forecast’ and analysis’ transient-eddy components, respectively, and the overbar denotes averaging over the cases for a given season. The forecast transient eddies are defined as $x^{\prime}=x-\overline{x}$, and similarly for the analysis. Figure 2 shows the evolution of $\gamma$ with lead time for GC and GDPS-CTL forecasts, against the GDPS analysis. While $\gamma$ remains close to 1 for all lead times for GDPS-CTL, indicating that its effective resolution does not change during the integration, for GC it rapidly decreases, leading to an effective resolution as low as 1000 km for 24-h forecasts. The smoothing effect further reduces the effective resolution to 2500$\sim$2700 km before it saturates around forecast days 3–5. It is worth noting here that GC still resolves scales smaller than these limits, but only partly ($10<\gamma<$ 1). While designing any strategy to leverage GC predictions to improve GEM, it would be prudent to not force the GEM model to follow GC for these variance-deficient scales in order to not adversely affect GEM’s effective resolution. This will be revisited in Section 5, where the concept of spectral nudging and its implementation in GEM are discussed.
Figure 2: Spectral (a, b) amplitude ratio $\gamma$ and (c, d) coherence $\rho$ for the global (a, c) kinetic energy spectrum of 500-hPa horizontal wind transient-eddy deviations and (b, d) variance spectrum of 500-hPa temperature transient-eddy deviations, or GDPS-CTL (blue) and for GC (red), after (dark to light) 24, 72, 120 and 240 h of integration. Spectra are averaged over 60 cases of boreal winter 2022.
Figure 2 also displays the spectral coherence defined as follows:
$\rho=\frac{\overline{\sigma_{x^{\prime}y^{\prime}}}}{\left(\overline{\sigma_{x^{\prime}x^{\prime}}}\;\overline{\sigma_{y^{\prime}y^{\prime}}}\right)^{1/2}}\;,$ (2)
where $\sigma_{x^{\prime}y^{\prime}}$ is the transient-eddy co-spectrum between the GDPS-CTL or GC forecasts and the GDPS analyses. Since the same GEM model is used to produce both the GDPS forecasts and backgrounds for the analysis steps, they may have common systematic errors, especially at shorter lead times. This may unfairly penalize GC, which is trained on ECMWF-ERA5 reanalyses. Despite that, Fig. 2 clearly indicates an improved spectral coherence with the GDPS analyses for a large range of scales – from synoptic to planetary, especially at longer lead times. This suggests that forcing GEM toward GC has potential for improving the predictive skill of its transient-eddy anomalies.
4.3 Verification against radiosondes
At ECCC, model developers often rely on forecast verification against ECMWF-HRES analysis (Lang et al., 2023) to avoid the problem associated with own-analysis verification (see, e.g., Casati et al. (2008), Privé et al. (2021)). This also helps to get around the issue of inhomogeneity of the observing network. Since GC was primarily trained with the ERA5 reanalysis and then fine-tuned on the ECMWF-HRES analysis, this likely introduces some dependency on the ECMWF analysis systems. To ensure that the dataset used to verify the forecasts is completely independent from both models as well as the GDPS data assimilation system, it was opted to rely mainly on observations from the global radiosondes network. Observations of zonal wind ($U$), geopotential height ($Z$), temperature ($T$) and dewpoint depression ($T–T_{d}$) at mandatory pressure levels between 1000 and 100 hPa will therefore be used as reference for verification.
Two flavors of the verification against radiosonde observations are performed: (1) verification at full resolution, with the GDPS-CTL and GC forecasts at their native resolution, and (2) verification at low resolution –- corresponding to the effective resolution of GC – wherein both the GDPS-CTL and GC forecasts are filtered to remove the variability at scales not adequately resolved by GC. The fact that GC has a much lower effective resolution than GDPS-CTL would otherwise provide an unfair advantage to GC by reducing the so-called “double penalty” for misplacing the predicted patterns. Low-resolution verification against radiosondes also allows the study of how suitable GC forecasts are for large-scale spectral nudging in GEM.
The filter applied here is a spectral filter based on spherical harmonics. It approximately removes scales smaller than 1000 km (total wavenumber 40) and fully preserves scales larger than 2000 km (wavenumber 20). These cutoff wavenumbers well reflect the effective resolution of GC at forecast days 3 to 5 (See Figs. 2a and b). Note also that the same filter configuration is applied to GC and GDPS-CTL (more details on this filter can be found in Appendix A).
Figure 3: Verification for 120-h forecasts from GC (red) and GDPS-CTL (blue) at their full (top row: a–e) and low (bottom row: e–h) resolutions against global radiosonde observations, averaged over 60 cases in boreal winter. The error standard deviation (STDE; solid lines) and bias (mean forecast minus observation; dashed lines) are shown for (a,e) zonal wind, (b,f) geopotential heights, (c,g) temperature and (d,h) dewpoint depression. Red and blue markers denote statistically significant differences in favor of GC and GDPS-CTL, respectively. Significance is computed using the $t$-test for bias and the $F$-test for STDE. No marker at a pressure level implies that the null hypothesis, stating that the statistics of the two samples are the same, cannot be rejected based on the 90th percentile. A low pass filter (Appendix A) is applied on both GC and GDPS-CTL to obtain the low-resolution fields (e–h).
4.3.1 At full resolution
Comparing the forecasts at their native resolution to the radiosonde observations reveals that for most variables and lead times, and for levels between 1000 and 200 hPa, GC forecasts are much closer to the observations than GDPS-CTL. Vertical profiles of error averages (bias) oriented as forecast minus observation and error standard deviations (STDE), calculated over the global domain for 120-h forecasts for boreal winter, highlight the typical verification differences between GC and GDPS-CTL (Fig. 3a–d). The differences between GC and GDPS-CTL in terms of biases are mixed, except for temperature during boreal winter, where clear improvements can be seen with GC. In terms of STDE, however, GC clearly shows large improvements over GDPS-CTL at almost every level, and particularly for zonal wind and dewpoint depression. Above 200 hPa, however, GC suffers from increased errors, with biases and STDE being comparable or worse than those in GDPS-CTL.
To summarize the scores at different lead times, variables and vertical levels a single forecast quality index based on the fractional root-mean-square error (RMSE) change is defined as follows:
$f_{RMSE}=\left[\frac{RMSE(x_{1})}{RMSE(x_{2})}-1\right]\times 100\%,$ (3)
with $x_{1}$ and $x_{2}$ here representing GDPS-CTL and GC, respectively. Positive values of $f_{RMSE}$ thus imply improved forecasts from GC measured by the RMSE against the radiosonde observations, while negative values represent the opposite. The forecast quality index $f_{RMSE}$ was computed every 24 h for each selected variable and level. The results for each variable were then vertically averaged for three layers: low-level (from 1000 to 850 hPa), mid-level (from 700 to 300 hPa) and high-level layer (from 250 hPa to 100 hPa).
Table 1: Latitude and longitude bounds for the 6 subdomains considered in Figures 5, 11 and 13.
[…]
Southern Hemisphere 90S 20S - -
Tropics 20S 20N - -
North America 25N 85N 170W 40W
Europe 25N 70N 10W 28E
Asia 25N 60N 65E 145E
[…]
Results for boreal winter (see Fig. 4a) show that $f_{RMSE}$ is quite constant with lead times for each variable and layer. The largest improvements from GC are obtained in the low-level layer for all variables and decrease with altitude, turning into deteriorations in the high-level layer for geopotential heights and temperatures. GraphCast geopotential heights are also degraded with respect to GDPS-CTL in the mid-level layer at short lead times. Results for boreal summer are roughly similar (not shown). The pressure weighting approach adopted for the training of GC (see Fig. 6 in Lam et al. (2022)) is likely responsible for this vertical variation in performances. The differences in the upper-level layer between ECMWF-ERA5 and GDPS analyses used for GC initial conditions (e.g., different bias correction approaches for radiances, different data assimilated, distinct biases in the IFS and GEM model, etc.) could also explain some of the degradations observed in the upper-level layer when GC is initialized with the GDPS analysis. However, GC initialized with ECMWF-ERA5 has also been shown to increase the RMSE with respect to ECMWF-HRES above the tropopause when compared to analyses (see, e.g., Fig. 19 in Lam et al. (2022)).
The region-wise variations in relative forecast performances can be examined by computing, as above, the forecast quality index $f_{RMSE}$ averaged over all variables and lead times for each vertical layer but, this time, for the 6 subdomains listed in Table 1 with their respective latitude and longitude limits. In low- and mid-level layers, GC forecasts show smaller RMSE in every subdomain in both seasons (Figs. 5a and c). In the upper-level layer, GDPS-CTL performs better during each hemisphere’s respective winter. However, GC’s poor performance during winter over the northern hemisphere is primarily originating from Europe. The largest improvements from GC in the mid- and high-level layers are found over the Tropics, which is consistent with the results of Lam et al. (2022) (see their Fig. 19).
Figure 4: Changes in the forecast quality index (Eq. 3) against global radiosonde observations, averaged over 3 layers and 60 cases during boreal winter 2022, corresponding to forecasts at (a) high resolution and (b) low resolution. Upward-pointing red triangles indicate a reduction of the RMSE in GC forecasts with respect to GDPS-CTL, whereas downward-pointing blue triangles indicate the opposite. The size of the triangles varies linearly up to a value of 5.0%. When the index is greater than this threshold, the size is kept constant, but the rounded values are shown. Triangles are color-filled if the significances computed using an $F$-test exceeds the 90th percentile. The values on the left-hand side show the averaged indices over all the lead times and all the variables for each layer.
Figure 5: Heatmap of the changes in the forecast quality index (Eq. 3) against global radiosonde observations for various geographical domains and atmospheric layers for (a, b) 60 cases of boreal winter 2022 and (c, d) 62 cases of boreal summer 2022. Verifications are presented for forecasts (a, c) at their full resolution and (b, d) after applying a low pass filter (Appendix A) on both GC and GDPS-CTL. Red shadings indicate a reduction of the RMSE by GC with respect to GDPS-CTL, while blue shadings indicate the opposite.
4.3.2 At low resolution
At low resolution, the forecast improvements in terms of RMSE from GC over GDPS-CTL are significantly reduced for all variables, levels, lead times, seasons and geographical domains (Figs. 3e–h and 4). This reduction in improvements at low resolution is particularly striking in the mid-level layer over the first 72 h (compare Figs 4a and b), leading to growing forecast improvements with time from GC that saturate or peak near 120 h (Fig. 5b) compared to a relatively steady improvement at all lead times found with the full resolution (Fig. 4a). This suggests that the improvements with GC for the first 72 h at full resolution largely originate from partial smoothing of the less-accurately predicted fine scales that reduces the double penalty.
Degradations in GC, seen at full resolution over Northern Hemisphere in boreal winter (mostly coming from Europe) and Southern hemisphere in boreal summer, are enhanced at low resolution (compare Figs 5a, c, and 5b, d). There are also notable degradations in the mid- and low-level layers over Asia in summer (Fig. 6d). The average reduction factor for $f_{RMSE}$ from full- to low-resolution verification is roughly between 2 to 3 (compare the left and right columns in Fig. 5), which confirms that the lower effective resolution of GC provides a significant advantage over the GDPS-CTL. Nevertheless, large scales of GC forecasts are still, on average, significantly closer to the radiosonde observations than those from GDPS-CTL in the mid- and low-level layers. This suggests that nudging these large-scale components of key GEM state variables toward GC could result in substantial improvements over GDPS-CTL.
5 Spectral nudging in GEM
5.1 The concept
Spectral nudging operates by gradually directing the model-predicted atmospheric large scales towards those found in some reference data which is considered to have a more accurate representation of those scales. This approach fundamentally differs from grid nudging or indiscriminate nudging, as it only targets a predefined range of scales. Even though an NWP model may employ grid-based spatial discretizations, the nudging increments at a given model level are computed by decomposing the model predictions and the reference fields in a spectral space, retaining only the target scales. Hence, this technique is referred to as “spectral nudging”.
In limited-area configuration, such as regional climate models, the nudging reference is usually derived from the same data that provides the lateral boundary conditions (LBCs). For extended-range integrations – spanning weeks to months or even years – the LBCs alone are insufficient to prevent the model-predicted large scales from drifting significantly. Consequently, spectral nudging becomes essential (Leduc and Laprise, 2009). For multi-month kilometer-scale downscaling applications, the evolving surface fields may also exhibit unacceptable deviations, necessitating nudging toward a reliable reference, as discussed by Husain et al. (2014).
5.2 Implementation in GEM
With the split-adjustment-type dynamics-physics coupling in GEM, the adiabatic dynamical core first solves the prognostic dynamical equations to obtain an intermediate state of the atmosphere. In the absence of spectral nudging, this interim dynamics solution serves as the input for the physical parameterization schemes, which compute the physics-induced tendencies. These tendencies are then coupled with the dynamics solution to obtain the complete model solution for a given time step. However, when spectral nudging is applied, the large scales in the solutions for a selected set of prognostic variables from GEM dynamics are first nudged toward the reference (here, from GC). Subsequently, the nudged solutions are fed to the physics schemes to compute the physics-induced tendencies before the eventual dynamics-physics coupling, following the regular model procedure.
In mathematical form, the nudging step at a given model vertical level, i.e., a constant-$\zeta$ surface, can be expressed as
$F_{nudge}=F_{GEM}+\omega\left[F_{GC}-F_{GEM}\right]_{LS},$ (4)
where $F_{GEM}$ is a prognostic variable predicted by GEM dynamics, $F_{GC}$ is the corresponding prediction from GC valid at the same time, $\omega$ is the nudging relaxation factor such that $(0\leq\omega\leq 1)$, and $F_{nudge}$ is the nudged solution. The subscript $LS$ in the above equation refers to some user-defined large scales targeted by the nudging mechanism. Through simple rearrangement of the terms on the right-hand-side of Eq.4, it can be shown that the large scales in $F_{nudge}$ are indeed a weighted average of those from $F_{GEM}$ and $F_{GC}$ (with $\omega$ being the weighting parameter), whereas the small scales from $F_{GEM}$ are retained entirely in $F_{nudge}$.
As shown by Husain et al. (2014), $\omega$ can vary with model vertical levels and in time, and is given by
$\omega=\frac{\beta(\zeta)}{\tau(t)}\Delta t,$ (5)
where $\beta(\zeta)$ defines the nudging vertical profile with respect the $\zeta$-coordinate, $\tau(t)$ represents the nudging relaxation time scale , and $\Delta t$ denotes the model time-step length (in the same unit as $\tau$). A vertically-variable $\beta$ allows for different nudging strengths at different model levels. Conversely, a time-dependent $\tau$ permits changing nudging strengths with time. In general, a large value of $\tau$ implies weak nudging, and vice versa.
To nudge a prognostic variable $F$ at any model level, it is first essential to spectrally decompose $(F_{GC}-F_{GEM})$ so that a filter can be applied in the spectral space to retain only the desirable scales. Spherical harmonics-based spectral decomposition would be the ideal choice for such a filter. However, this would require multiple transformations of model solutions between the model’s Yin-Yang grids and an intermediate global Gaussian grid, resulting in significant increase in the computational cost. As the present study is primarily of a proof-of-concept nature, in order to keep the approach simple, the spectral filter for nudging was instead chosen to be based on discrete cosine transform (DCT) (Denis et al., 2002), and the filtering is performed separately on the Yin and Yang grids. Additionally, nudging is also separately applied to the Yin and Yang grid solutions after every dynamics sub-step.
Large scales that are retained with spectral nudging are defined by users through configurable parameters. Instead of an abrupt cut-off for the target scales, which may lead to Gibbs phenomena (Sardeshmukh and Hoskins, 1984), a soft cut-off is employed within the DCT-based spectral filter. With this method, the spectral filter relies on two different cut-off scale parameters denoted by $\lambda_{LS}$ and $\lambda_{SS}$ such that scales larger than $\lambda_{LS}$ are fully retained while those smaller than $\lambda_{SS}$ are completely eliminated. A partial filtering, similar to the one utilized by Husain et al. (2014), is applied for the scales in between $\lambda_{LS}$ and $\lambda_{SS}$. Further information on the DCT-based filter and its response is provided in Appendix B.
5.3 Optimal configuration
A series of systematic sensitivity experiments was carried out to identify an optimal spectral nudging configuration for GDPS. All subsequent results presented in the following section are based on this optimal configuration which, from this point onward, will be referred to as GDPS-SN. It is important to note that spectral nudging is currently only applied within the forecasting component of GDPS, i.e., GDPS-CTL, GDPS-SN and GC are all initialized with the same analyses.
Some key aspects of this optimal setup are discussed below.
5.3.1 Nudged variables
Nudging is only applied to the $u$-$v$ components of wind and virtual temperature. While nudging specific humidity can improve the global bias of temperature and humidity in the boundary layer, the computational cost outweighs the benefits. Moreover, nudging specific humidity tends to negatively impact tropical cyclone intensity during the tested period, making it less desirable.
5.3.2 Nudging vertical profile
Although different vertical profiles, defined by $\beta(\zeta)$, have been explored, a plateau-shaped profile, as shown in Fig. 6, is found to yield the best results. This profile involves no nudging in the boundary layer (below 850 hPa) and the stratosphere (above 250 hPa). Relatively weaker performance of GC in the stratosphere – as has been mentioned earlier – is the reason for no stratospheric nudging. Within the boundary layer, nudging is avoided to prevent potential negative impacts arising from inconsistencies between the surface forcing in GDPS and GC training data (ERA5). Furthermore, GC-13 does not have sufficient vertical resolution in the boundary layer to provide a useful reference.
[…]
5.3.3 Nudging length scales
It has been already shown that the fine-scales in GC suffer from considerable smoothing, worsening at longer lead times (see Figs. 1 and 2). The implementation of spectral nudging in GEM allows for time-varying cut-off scales for nudging. However, the improvements with such an approach over a stationary cut-off scale are found to be negligible. Therefore, based on the comparison of spectra between GDPS and GC, the DCT-based filter for the optimal configuration utilized in this study is configured with $\lambda_{LS}$=2750 km and $\lambda_{SS}$=2250 km. This implies that scales larger than 2750 km are entirely retained, while those smaller than 2250 km are completely eliminated. The scales in between are subjected to partial filtering (see Fig. 20). This choice of nudging length scales corresponds to the effective resolution of GC beyond day 3 (see Figs. 1 and 2).
5.3.4 Nudging relaxation time
The nudging relaxation time, $\tau$, is analogous to the $e$-folding time as it determines the time-rate of decay in the large-scale differences between the model and the driving data in the absence of any other process. Arguably, $\tau$ is one of the most important configurable parameters for spectral nudging. However, selecting an optimal $\tau$ is not straightforward. First, it is important to note that GC inferences are only available every 6 hours and become increasingly inaccurate with longer lead times (as demonstrated by spectral coherence in Fig. 2). During model integration, when a valid GC inference is not available at certain time steps, an estimate is approximated through linear interpolation between two consecutive GC inferences around the time of interest. In this regard, Omrani et al. (2012) have argued that $\tau$ should not be less than $\tau_{a}$ which denotes the time interval between consecutive driving fields (here, GC inferences). Strong nudging with $\tau<\tau_{a}$ can lead to erroneous evolution of nudged fields. Furthermore, sensitivity experiments carried out during this study revealed considerable deterioration of model’s skill over complex terrain with $\tau<\tau_{a}$ (not shown). As a result, $\tau_{a}$ may safely be considered the lower bound for $\tau$.
It is also crucial to ascertain an acceptable upper bound for $\tau$. Intuitively, as the accuracy of reference nudging data worsens with increasing forecast lead time, it might seem prudent to employ a large $\tau$. However, sensitivity experiments have revealed that too large of a nudging relaxation time ($\tau>24$ hr) can result in excessive undesirable smoothing in the nudged fields around the cut-off scales developing at day 3 or 4. Eventually, this smoothing affects all scales by day 10 of the forecast (not shown). This increased smoothing with overly large $\tau$, is caused by the averaging effect resulting from blending two forecasts, with $\tau=48h$ maximizing the effect. Based on the insights derived from the sensitivity tests, a nudging relaxation time of 12 hr is chosen to be optimal.
6 Impact of large-scale spectral nudging
6.1 Verification with global power spectra
The power spectra of the GDPS-SN forecast variables remain approximately the same as in the GDPS-CTL (Fig. 7), which essentially demonstrates that the new hybrid system can adequately resolve the full range of scales present in the GDPS-CTL, without significant smoothing. The resulting spectral amplitude ratio of transient-eddy anomalies, $\gamma$ (Eq. 1), also has values very close to one for all prognostic variables, lead times and spatial scales for both GDPS-CTL and GDPS-SN. This provides a solid proof-of-concept for the spectral nudging approach (Fig. 8). The only exception is a 10-percent reduction of $\gamma$ for scales between 2000 and 4000 km at lead times approaching day 10 (wavenumbers 10-20 in Figs. 8a, b). This reduction is explained by GC having some variance deficiency at the corresponding range of scales (2000-2750 km) for longer lead times (beyond day 4). Sensitivity tests with time-varying cut-off length scales that more strictly follow GC’s time-evolving effective resolution, showed that this issue could be avoided, but at the expense of significantly reduced GDPS-SN forecast improvements with respect to the GDPS-CTL. This is caused by a weaker constraint being imposed on GDPS-SN towards GC at synoptic scales by such a configuration.
Comparison of the spectral coherence (Eq. 2) between the forecasts and GDPS analysis shows an improved skill of GDPS-SN over GDPS-CTL (Figs. 8c, d). These improvements are somewhat weaker than those seen with GC at shorter lead times (up to day 5), but become comparable beyond day 5 (Figs 2c, d). The scales relevant for improvements at shorter lead times are smaller than those included in the spectral nudging configuration. Nudging the smaller scales would therefore bring the coherence on par with GC but it would lead to variance deficiency, as discussed earlier.
Figure 7: Same as in Fig. 1, but for the analysis, GDPS-CTL and GDPS-SN.
Figure 8: Same as in Fig. 2, but for GDPS-SN and GDPS-CTL.
An illustrative example of the value added by the new hybrid system over GC is provided in Fig. 9. Large-scale features of transient-eddy anomalies in GDPS-SN tend to be better (but not fully) aligned with GC than with GDPS-CTL, while containing similar level of fine-scale details as in GDPS-CTL. Figure 10 suggests that large-scale components of GDPS-SN forecast fields are preconditioned with the piloting AI model. Therefore, further improvements to its predictions have potential for similarly improving GDPS-SN.
Figure 9: 10-day forecast of 700-hPa transient-eddy specific humidity anomalies, valid 22 February 2022 at 12z for (a) GDPS-CTL, (b) GC and (c) GDPS-SN.
6.2 Verification against radiosondes
Comparison of forecasts at full resolution against the global radiosonde observations clearly indicates that GDPS-SN forecasts are closer to the observations than GDPS-CTL forecasts for basically all variables, lead times and levels, and over both seasons (Fig. 10). The improvements gradually grow in the first 72 h and are remarkably constant with altitude, which contrasts with the strong vertical variations observed with GC in Section 4. This demonstrates that large-scale spectral nudging, despite being only applied in the mid-level layer for winds and temperature, is sufficient to impart similar improvements at both upper- and lower-level layers that are not constrained by spectral nudging. Compared to boreal winter (Fig. 10a), the improvements from GDPS-SN are reduced by a factor of about 2 during boreal summer (Fig. 10b), which is not surprising given that the benefits from GC over GDPS-CTL at the large scales are smaller by about the same factor for zonal wind and temperature in the mid-level (not shown).
In terms of region-wise performances, the GDPS-SN has smaller RMSE than GDPS-CTL in every subdomain except over Asia in the mid- and upper-levels during boreal summer (Figs. 11c and d), which is relatively consistent with the degradations from GC over the GDPS-CTL at the large scales previously noted for the mid-levels over that part of globe during this season (see Fig. 5d). Unlike GC, the improvements from GDPS-SN over GDPS-CTL are virtually identical when only the large-scale components of the forecast are compared to the radiosonde observations (compare the left and right columns in Fig. 11 and, again, in Fig. 6). This does not necessarily imply that the improvements only originate from the large scales. It rather implies that the relative RMSE reductions with GDPS-SN are of similar level for both the large and small scales. In other words, as the small scales are generally conditioned by the large scales, these results suggest that the fine scales predicted by GDPS-SN benefit from fractional improvements that are on par with its large-scale skill enhancement.
Figure 10: Same as in Fig. 4, except that upward-pointing red triangles indicate a reduction of the RMSE in GDPS-SN with respect to the GDPS-CTL, whereas downward-pointing blue triangles indicate the opposite. Only verification using forecasts at full resolution is shown.
Figure 11: Same as in Fig. 5, except that red shadings indicate a reduction of the RMSE in GDPS-SN with respect to GDPS-CTL, while blue shadings indicate the opposite.
Figure 12: Same as in Fig. 4, except that upward-pointing red triangles indicate a reduction of the RMSE in GDPS-SN with respect to GC, whereas downward-pointing blue triangles indicate the opposite. Only verification using forecasts at low resolution is shown but for both seasons.
It is also interesting to compare GDPS-SN to GC. For the sake of fair comparison, this will only be done at low resolution. Predictions from GDPS-SN are closer to the radiosonde observations than GC in the upper and mid-level layers, except for dewpoint depression. Conversely, for the low-level layer, GC is closer to the observations, except for geopotential height (Fig. 12). The degradations with respect to dewpoint depression, also observed in the mid-level layer where spectral nudging is active, suggest that further gains could likely be achieved by nudging the humidity field in GEM towards GC. The results also suggest that extending spectral nudging to the boundary layer could be beneficial. This, however, would necessitate an AI model with more vertical levels in the boundary layer and fine-tuned to emulate the GDPS analyses rather than ECMWF-ERA5, for improved consistency with the GEM model’s lower-boundary forcings.
The small but statistically significant GDPS-SN improvements over GC in the mid-level layer for temperature and zonal winds, two variables that are nudged toward GC, that grow with the integration time, are somewhat surprising (Fig. 12). This could be due to a positive feedback mechanism between the large and small scales in GEM. The averaging effect caused by the blending of two forecasts due to a long relaxation time, as previously discussed in Section 5, may also be partly responsible for these improvements. Answering this question is, however, deemed outside the scope of this paper.
Finally, the overall region-wise comparison of GDPS-SN and GC forecasts at low resolution (Fig. 13) shows that the results that were observed over the globe hold also relatively well for each of the subdomains. The fact that spectral nudging is applied only in the mid-level layer targeting only the synoptic scales, it considerably limits the ability of GDPS-SN to fully leverage the strengths of GC for seasonal periods and over regions where mesoscale near-surface conditions play a major role in determining the weather. This may explain the degradations seen over the tropics.
Figure 13: Same as in Fig. 5, except that red shadings indicate a reduction of the RMSE in GDPS-SN with respect to GC, whereas blue shadings indicate the opposite. Only verification using forecasts at low resolution is shown.
6.3 Verification against ECMWF analyses
The comparison of GDPS-SN and GDPS-CTL forecasts against ECMWF-HRES analyses leads to very much similar results as the comparison against radiosonde observations, with general improvements for all variables, lead times and pressure levels (not shown). Figure 14 provides an example of improvements in anomaly correlation coefficient (ACC). For 500-hPa geopotential heights, the ACC improvements over the Northern Hemisphere region at forecast day 7 are equivalent to an 18-h increase in predictability during boreal winter and to an 8-h increase in summer. The corresponding ACC improvements over the Southern Hemisphere region are about 2 hours during austral winter and even exceed 24 hours during austral summer. There are also notable improvements in zonal wind ACC in the tropics that are equivalent to 24 hours or more (e.g., see Fig. 14b,e). The strength of GC in improving the ACC around day 5 (and beyond) is thus well leveraged by the hybrid system.
Figure 14: Anomaly correlation coefficient against the ECMWF-HRES analyses with GDPS-CTL (blue) and GDPS-SN (red) for (a,d) northern hemisphere 500-hPa geopotential height, (b,e) tropical 850-hPa zonal wind, and (c,f) southern hemisphere 500-hPa geopotential height. Results are presented for 60 cases of boreal winter 2022 (left column: a–c) and 62 cases of boreal summer 2022 (right column: d–f). No marker at a pressure level implies that the null hypothesis, stating that the statistics of the two samples are the same, cannot be rejected based on the 95th percentile.
6.4 Verification against surface observations
For verification at the surface level, combined SYNOP, METAR and Surface Weather and marine OBservations (SWOB; available only over Canada and distributed by ECCC) data were used for surface pressure, screen-level temperature, dewpoint depression and anemometer-level wind speed. For precipitation, 24-h accumulation ground observations are subject to collection and quality control by the CAnadian Precipitation Analysis system (CaPA; Lespinas et al. (2015)). Any data pertaining to stations with an altitude difference larger than 100 m with respect to GDPS orography, along with any observed wind speed below 1.5 $\text{m s}^{-1}$, are excluded.
Figure 15: Heatmap of the changes in the forecast quality index (Eq. 3) against combined SYNOP, SWOB and METAR observations for surface pressure ($P_{s}$), screen-level temperature ($T$), dewpoint depression ($T-T_{d}$) and anemometer-level wind speed ($W$) in various geographical domains for (a) 60 cases of boreal winter 2022 and (b) 62 cases of boreal summer 2022. Red shadings indicate a reduction of the RMSE in GDPS-SN with respect to GDPS-CTL, while blue shadings indicate the opposite.
Figure 16: Q-Q plots of day-5 GC (black), GDPS-CTL (blue) and GDPS-SN (red) forecasts against combined SYNOP, SWOB and METAR for: (a, b) anemometer-level wind speeds and (c, d) screen-level temperatures over North America for (a, c) 30 cases of boreal winter valid at (a) 00 and (c) 12 UTC, and (b, d) 31 cases of boreal summer valid 00 UTC. Local standard time over North America is mostly between UTC-5 and UTC-8 h. Diamond symbols denote selected percentiles with their values printed. Shadings denote 95% confidence bands for the inverse cumulative distribution functions, based on the Kolmogorov-Smirnov statistic.
Comparison of GDPS-SN and GDPS-CTL using the forecast quality index (Eq. 3) is summarized in Fig. 15. Guidance by GDPS-SN is generally improved for all variables and regions. The improvements are, however, notably smaller than in the upper air, which is particularly the case for the tropics, where achieving considerable improvements appears to be difficult without introducing spectral nudging in the boundary layer. The same could be concluded for the screen-level thermodynamic variables and wind. On the other hand, the spectral nudging in the mid-level layer has the largest positive impact on the surface pressure via redistribution of upper-air mass. The improvements in boreal summer are notably smaller than in boreal winter, consistent with the upper-air verification.
The ability of the GPDS-SN to represent high-impact weather, such as high wind speeds and extreme temperatures, is not significantly different from the GDPS-CTL but, at the same time, it is substantially better than GC, which provides a strong example of the value added by the hybrid system over GC (Fig. 16). For example, the distribution of late-afternoon wind speeds over North America provided by GC shows a considerable shift toward lower values, which increases at high percentiles of the distribution, resulting in a -4 $\text{m s}^{-1}$ bias at the 99.9th percentile (Figs. 16 a, b). GDPS-CTL also has a negative bias but affecting only higher percentiles and, more importantly, GDPS-SN closely follows GDPS-CTL, with little or no sign of deceleration at extreme wind speeds for both seasons. GraphCast also suffers a 2-4 K cold bias at mid-to-high temperature percentiles over North America in summer, which has no impact on the GDPS-SN; its distribution aligns very well with the observations. GraphCast’s poor performance with respect to the extremes can be partly explained by its lack of well-resolved fine scales that are generally critical for representing the extremes. Furthermore, inconsistencies in the surface forcing between ERA5 and GDPS analyses may also be a contributing factor for the observed underestimation of temperature extremes by GC. As the spectral nudging configuration presented in this study only targets synoptic scales and the boundary layer is allowed to evolve freely, GDPS-SN is not adversely affected by GC’s limitations regarding the extremes.
Spectral nudging also significantly improves boreal winter precipitation forecast guidance over North America for a large range of thresholds from 1 to 50 mm day-1, mostly by reducing the false alarm ratio, with no reduction of probability of detection, which results in a reduced moist frequency bias and an improved equitable threat score (Fig. 17). These improvements seem to be mostly generated via an improved prediction of synoptic-scale flow. There is, however, not much of an impact of spectral nudging on summer precipitation over North America. Improving precipitation in summer, when weak synoptic forcings are prevalent, may necessitate the application of spectral nudging in the boundary layer.
Figure 17: Quantitative precipitation forecast verification for day 5 assessed by 24-hr precipitation threshold (accumulations between 108 and 132 h of integration) with the (a) equitable threat score, (b) frequency bias, (c) probability of detection, and (d) false alarm ratio. Results from the 60 GDPS-CTL (blue) and GDPS-SN (red) cases of boreal winter are compared to ground observations over North America used by CaPA. Red and blue line markers denote statistically significant differences in favor of GDPS-SN and GDPS-CTL, respectively. Significance is computed by bootstrapping 3-day data blocks consisting of forecast-observation pairs from all stations of a given region. No marker at a threshold level implies that the null hypothesis, stating that the statistics of the two samples are the same, cannot be rejected based on the 90th percentile.
6.5 Tropical cyclone evaluation
Evaluation of the impact of spectral nudging on the tropical cyclone (TC) forecast guidance is performed using the data from the International Best Track Archive for Climate Stewardship (IBTrACS; Knapp et al. 2010). For the results presented in this section, GDPS TC forecasts are matched with observations, by excluding from consideration any storm that is not detected by both GDPS and IBTrACS systems.
Spectral nudging improves TC trajectories by reducing the errors resulting from GDPS-CTL’s tendency of predicting storms that move too slowly and veer too much to the right (with respect to its observed track), especially at lead times beyond day 5 (Figs. 18a and b). Conversely, there is little or no significant impact of spectral nudging on the TC intensity, as measured by the minimum central pressure and maximum sustained wind speed (Figs. 18c and d). Both GDPS-CTL and GDPS-SN have a similar week intensity bias, which is a well-known weakness of the GEM model (see McTaggart-Cowan et al. (2024), for more details). Severe TC weak-intensity biases are typical of state-of-the-art deterministic AI-based forecast models (e.g., Bouallègue et al. (2024)) due to the smoothing effect caused by learning to minimize the MSE. However, AI-based models, including GC, are generally capable of predicting the TC trajectories with enhanced accuracy (Lam et al., 2022). Therefore, the improvements in the predicted trajectories by GDPS-SN show that the proposed hybrid system is capable of leveraging the well-documented strength of the AI-based models, without sacrificing TC intensity. Furthermore, these results suggest that the weak-intensity bias for TCs could be reduced in future by focusing on improving the physics-based NWP component of the hybrid system.
Figure 18: Tropical cyclone tracking performance comparison of the GDPS-CTL (blue) and GDPS-SN (red) against IBTrACS data in three northern hemisphere basins (West Atlantic, East and West Pacific) for 62 cases of boreal summer 2022: (a) along-track and (b) cross-track mean position errors, (c) mean central pressure error and (d) mean maximum wind speed error. Negative along-track errors imply cyclones move too slowly and positive cross-track errors imply cyclones veer to much to the right with respect to the observed trajectory.
6.6 Computational cost of spectral nudging
The current implementation of spectral nudging in GEM with the presented optimal configuration leads to an increase of computational cost by approximately 25%. Due to the proof-of-concept nature of this study, no optimization of the informatics code has been performed. Therefore, the cost could be considerably reduced through future code optimization, especially for the spectral filter. Further improvements could be attained by considering time-varying nudging increments, where nudging would only be applied at time-steps close to the inference times of GC. These considerations however lie beyond the scope of this study.
7 Summary and Future Work
The recent emergence of AI models has begun to disrupt the operational weather forecasting paradigm which has long been dominated by the physics-based NWP models. Naturally, serious questions about the future role of AI and NWP models in operational forecasting are arising. For instance, are AI models capable of adding any reliable and meaningful value for weather services, or will NWP models eventually disappear from the forecasting landscape and be replaced by AI models? This study attempts to address these pressing questions by choosing the GEM and GC models to represent the two forecasting approaches.
First, by comparing and contrasting the GEM and GC models, their relative strengths and weaknesses are identified. Skill verification against observations and GDPS analyses reveals that GC suffers from excessive fine-scale smoothing, affecting scales as large as 2500 km beyond day 3 of a 10-day forecast. However, despite smoothing, the results demonstrate that large-scale features (wavelength$>$2500 km) in GC exhibit better skill compared to GEM, which is evidenced by increased spectral coherence relative to the GDPS analyses and reduced STDE against radiosonde observations.
Inspired by the improved large-scale skill of GC, a hybrid NWP-AI system – namely, GDPS-SN – has been developed to produce global forecasts where GEM-predicted large scales in horizontal wind and temperature are spectrally nudged toward inferences from GC. This hybrid system is capable of generating real-time forecasts with accuracy that surpasses ECCC’s operational GPDS. For example, RMSE of the 500-hPa geopotential height is reduced by 5-10%, with an overall predictability improvement of 12-18 hours, peaking at day 7 of a 10-day forecast. Notably, trajectories of tropical cyclones are predicted with enhanced accuracy without significant changes in intensity. Regarding precipitation, statistically significant improvements are obtained in equitable threat score, frequency bias and false alarm ratio, particularly in non-summer regimes. Skill enhancements with GDPS-SN are generally found to be more remarkable in winter when large scales apparently play a more critical role in determining weather at the regional level.
In addition to enhancing the skill and predictability of GEM, the proposed hybrid NWP-AI system is shown to generate the full spectra of fine scales that is missing in GC inferences. Furthermore, the new system is capable of producing all the usual forecast variables that operational meteorologists rely on for forecasting weather in general and high-impact events in particular.
Comparisons against near-surface observations have revealed that – unlike GC, which has a limited ability to capture the extremes in wind speed and temperature – weather extremes predicted by GDPS-SN closely follow the operational GDPS-CTL. This implies that the fine scales predicted by GPDS-SN not only possess similar level of energy as GDPS-CTL, they are also comparable in terms of skill.
Other aspects of near-surface predictions from GDPS-SN, however, show relatively modest improvement (e.g., screen-level scores). Lack of vertical resolution in the boundary layer in GC-13 and inconsistencies in surface forcing (such as orography) between ERA5 and GDPS analyses were the reasons why nudging in the boundary layer has not been considered in this study. This largely explains the less-than-remarkable near-surface performance of GDPS-SN. However, work is currently in progress at ECCC to fine-tune GC to emulate the operational GDPS analyses. This fine-tuned version is expected to enhance GC’s skill with the 37-level version, particularly in the stratosphere and the boundary layer. Any such improvement could be easily leveraged by the hybrid system in the near future. Importantly, a fine-tuned GC would allow for nudging in the boundary layer which is expected to result in further enhancement of near-surface skill with GDPS-SN.
Finally, results from the proposed hybrid NWP-AI system imply that, rather than viewing AI and NWP models as two competing paradigms, it is likely more prudent to consider these forecasting methods as complementary. A well-designed fusion of these two approaches can significantly mitigate their individual limitations while allowing for harnessing their respective strengths to provide better meteorological guidance. Therefore, instead of focusing on efforts to replace one with the other, future research should prioritize improving both NWP and AI models.
Acknowledgements.
[…]
$f_{n}=\exp\left[-\left(\frac{n(n+1)}{n_{o}(n_{o}+1)}\right)^{r}\right],$ (6)
where $n$ denotes the total wavenumber, $n_{o}$ is the cut-off total wavenumber and the exponent $r$ is a non-dimensional parameter related to the sharpness of the filter response. Based on the spectral comparison of the GDPS and GC predictions, the filter is configured by setting $n_{0}=30$ and $r=4$. As shown in Fig. 19, the resulting filtered fields fully retain amplitudes for scales associated with approximately $n<20$ (wavelength$>$2000 km) and fully removes scales corresponding to $n>40$ (wavelength$<$1000 km). The physical fields reconstructed from the filtered spectra leads to the desired filtered fields.
Figure 19: Response of the spherical-harmonics based spectral filter with $n_{0}=30$ and $r=4$ (see Eq. 6).
[…]
$f_{F}(m,n)=\begin{cases}0.0,&\text{if }\hat{\alpha}>\lambda_{LS}/\lambda_{SS}\\ \left[\cos\left(\frac{\pi}{2}\frac{\alpha\lambda_{LS}/(2\Delta)-1}{\lambda_{LS}/\lambda_{SS}-1}\right)\right]^{2},&\text{if }1.0<\hat{\alpha}\leq\lambda_{LS}/\lambda_{SS}\\ 1.0,&\text{if }\hat{\alpha}\leq 1.0\end{cases}$ (7)
where $\Delta$ denotes the model grid spacing, $\alpha$ is the normalized two-dimensional wavenumber given by $\alpha=\sqrt{\frac{m^{2}}{N_{i}^{2}}+\frac{n^{2}}{N_{j}^{2}}}$ associated with each of the Yin-Yang sub-domains of size $(N_{i}\times N_{j})$, and $\hat{\alpha}=\alpha\lambda_{LS}/(2\Delta)$.
Figures 20a and 20b illustrate the response of this DCT-based filter for a configuration based on $\lambda_{LS}=1500$ km and $\lambda_{SS}=1000$ km. The figures demonstrate that the DCT-based filter is capable of targeting the desired scales for the individual Yin-Yang sub-domains. The response of the filter over the globe was determined by inspecting the spectral variance ratio of the filtered and unfiltered fields over a global Gaussian grid with respect to a spherical harmonics-based decomposition. The corresponding results are presented in Fig. 20c. The figure reveals that, even though DCT is primarily designed for LAMs where curvature effects of spherical surfaces are neglected, its application on global Yin-Yang grid does not result in any major issue as the filter is found to result in an acceptable response.
Figure 20: Response of the DCT-based filter (see Eq. 7) with respect to the (a) Yin grid, (b) Yang grid, and (c) a global Gaussian grid.

### v2

@@
 Operational meteorological forecasting has long relied on physics-based numerical weather prediction (NWP) models.
-Recently, this landscape has been disrupted by the advent of data-driven artificial intelligence (AI)-based weather models, which offer tremendous computational performance and competitive forecasting skill.
+Recently, this landscape is facing disruption by the advent of data-driven artificial intelligence (AI)-based weather models, which offer tremendous computational performance and competitive forecasting skill.
 However, data-driven models for medium-range forecasting generally suffer from major limitations, including low effective resolution and a narrow range of predicted variables.
@@
 State-of-the-art numerical weather prediction (NWP) models generally include some form of a dynamical core that solves the nonlinear partial differential equations governing the resolved-scale dynamical and thermodynamical processes within the atmosphere.
-Solutions from the dynamical core are subsequently combined with tendencies from parameterized physics schemes, representing the unresolved and subgrid-scale processes that are not explicitly accounted for by the dynamical equations.
+Solutions from the dynamical core are subsequently combined with tendencies from parameterized physics schemes, representing diabatic, frictional and subgrid-scale processes that are not explicitly accounted for by the dynamical equations.
 Although statistical approaches have been explored in the past — particularly for downscaling purposes (Yu et al., 2006; Busuioc et al., 2008; Cheng et al., 2014) — physics-based NWP models have long been the foundational approach for operational meteorological forecasting.
@@
 Keisler (2022) presented an important breakthrough demonstrating considerable potential for weather forecasting with data-driven models.
-This was quickly followed by Pangu-Weather from Huawei (Bi et al., 2022, 2023), GraphCast from Google-DeepMind (Lam et al., 2022, 2023) and several other models (e.g., Pathak et al.
+This was quickly followed by Pangu-Weather from Huawei (Bi et al., 2022, 2023), GraphCast from Google DeepMind (Lam et al., 2022, 2023) and several other models (e.g., Pathak et al.
 2022; Chen et al.
@@
 They can speed up the production of multi-day forecasts by orders of magnitude while using a fraction of the computational resources.
-Furthermore, many of these models are not only comparable to the NWP models with respect to their prediction accuracy, in terms of a number of the performance metrics, they can even surpass the skill of the IFS (Integrated Forecasting System; Lang et al.
+Furthermore, many of these models are not only comparable to the NWP models with respect to their prediction accuracy, in terms of a number of the performance metrics, they can even surpass the skill of the Integrated Forecasting System (IFS; Lang et al.
 2023) from ECMWF, which is widely accepted as the benchmark within the NWP community (Bi et al., 2023; Lam et al., 2023; Chen et al., 2023a).
@@
 A major weakness of most of the currently available data-driven models is considerable fine-scale smoothing (Bi et al., 2022; Lam et al., 2022).
-Notably, data-driven models based on generative AI, e.g., diffusion-based GenCast model from Google-DeepMind (Price et al., 2024), can address the fine-scale smoothing problem to some extent.
+Data-driven models based on generative AI, e.g., diffusion-based GenCast model from Google DeepMind (Price et al., 2024), can address the fine-scale smoothing problem to some extent.
 However, as of now, such approaches generally require considerable increase in computational cost, thereby diminishing one of the principal advantages of the AI models.
@@
 Another common limitation of the current generation of AI models is the limited range of predicted variables.
-At present, these models can only be trained to emulate fields that are present in the analyses dataset, although training to predict non-analyzed variables (e.g., precipitation in ERA5) would generally lead to suboptimal inference.
+At present, these models can only be trained to emulate fields that are present in the analyses dataset.
+However, training to predict non-analyzed variables present in the analyses (e.g., precipitation in ERA5) would generally lead to suboptimal inference.
 Furthermore, adding the full suite of physics-related prognostic variables required for operational NWP, such as categorization of cloud and precipitation types, would be cost-prohibitive with current AI model architectures.
 Time resolution of inferences is another constraint for most of the current AI models, which make predictions at 6-hour intervals, far coarser than physics-based NWP models.
-Pangu-Weather (Bi et al., 2022) provides separate models for 1-, 3-, 6-, and 24-hour predictions, but the models with shorter lead times suffer from larger error accumulation.
+Pangu-Weather (Bi et al., 2022) provides separate models for 1-, 3-, 6-, and 24-hour predictions, but the models with shorter forecasting steps suffer from larger error accumulation.
 In summary, despite their strengths, the aforementioned limitations make it implausible for the AI models to replace the classical NWP approach in operational weather forecasting.
-The GEM (Global Environmental Multiscale) model (Girard et al., 2014), which is employed by Environment and Climate Change Canada (ECCC) for operational weather forecasting, relies entirely on the physics-based NWP approach.
-The principal objective of this study is to determine whether inferences from a state-of-the-art AI-based weather emulator can be leveraged through large-scale spectral nudging to enhance the predictability of the GEM model, particularly for global weather forecasting.
+The ultimate goal of this study is to develop a hybrid NWP-AI system for real-time forecasting applications that combines the strengths of the physics-based NWP approach with those of the AI-based approach while addressing the inherent limitations of the AI models.
+The physics-based model chosen for this study is the GEM (Global Environmental Multiscale) model (Girard et al., 2014), which is employed by Environment and Climate Change Canada (ECCC) for operational weather forecasting.
+The AI model in this study is represented by GraphCast from Google DeepMind.
+Based on internal testing at ECCC, among the different freely available open-source AI models, GraphCast demonstrates the best accuracy with respect to most of the metrics.
+The study aims to determine whether large-scale spectral nudging of GEM predictions toward inferences from a state-of-the-art AI-based weather emulator, such as GraphCast, can improve GEM’s predictability and offer a feasible approach for developing a suitable hybrid NWP-AI system.
 Spectral nudging can be seen as an indirect suboptimal data assimilation method (von Storch et al., 2000).
@@
 1996), it is primarily used to improve dynamical downscaling of coarse-resolution atmospheric model simulations over high-resolution limited area domains (Leduc and Laprise, 2009; Husain et al., 2014).
-Spectral nudging has also been applied for global atmospheric hindcasting (Schubert-Frisius et al., 2017).
-The ultimate goal of this study is to develop a hybrid NWP-AI system that combines the strengths of both approaches while addressing the inherent limitations of the AI-based models.
-The AI model chosen for this study is the GraphCast model from Google-DeepMind.
-Based on internal testing at ECCC, among the different freely available open-source AI-based models, GraphCast demonstrates the best accuracy with respect to most of the metrics.
+Spectral nudging has also been explored for global atmospheric hindcasting (Schubert-Frisius et al., 2017).
+Spectral nudging has so far been unfeasible for real-time forecasting applications due to the absence of timely and accurate reference data.
+The emergence of AI models, like GraphCast, has unlocked the potential for applying spectral nudging in real-time operational weather forecasting.
+In this context, application of spectral nudging for enhancing real-time NWP represents an important novelty proposed by this study.
+Although the hybridization concept is illustrated with a specific combination of AI and NWP models, the proposed strategy is expected to be broadly applicable to any other combination, provided that the AI model offers added value over the NWP model.
 To achieve the most optimal fusion of GraphCast and GEM, it is essential to ascertain the relative strengths and limitations of these models through objective evaluation against observations and analyses.
 A detailed comparative analysis of these models is also critical for devising the most optimal spectral nudging configuration, and therefore, constitutes a significant component of this study.
-Spectral nudging has so far been unfeasible for real-time forecasting applications due to the lack of availability of timely and accurate reference data.
-The emergence of AI models like GraphCast has unlocked the potential for applying spectral nudging in real-time operational weather forecasting.
-In this regard, application of spectral nudging for enhancing real-time NWP represents an important novelty proposed by this study.
-Although this paper illustrates the hybridization concept with a specific combination of the AI and NWP models, the proposed strategy is expected to be broadly applicable to any other combination as long as the AI model offers some added value over the NWP model.
 This article is organized as follows.
@@
 Detailed evaluations of the spectral nudging-based hybrid NWP-AI system are presented in 6.
-Finally, section 7 summarizes the conclusions of this study and outlines potential directions for future work.
+Finally, section 7 summarizes the main conclusions of this study and outlines potential directions for future work.
 2 The NWP System
 2.1 The GEM Model
-The dynamical core of the GEM model solves the elastic Euler system of equations.
-These equations are first transformed from regular height coordinate to some form of a terrain-following coordinate (TFC) in the vertical, denoted by $\zeta$ in the model.
-All current operational GEM-based NWP systems employ a log-hydrostatic-pressure-type TFC (Girard et al., 2014; Husain et al., 2021).
-The nonhydrostatic version of the dynamical core involves eight prognostic variables: horizontal velocity components ($u$ and $v$), vertical velocity ($w$), virtual temperature ($T_{v}$), geopotential ($\phi$), vertical motion with respect to the TFC ($\dot{\zeta}$), hydrostatic surface pressure ($\pi_{s}$), and nonhydrostatic pressure perturbation ($q=\ln\frac{p}{\pi}$, where $p$ and $\pi$ denote the complete and hydrostatic pressure, respectively).
-The advection terms in the model are treated with semi-Lagrangian approach while a two-time-level iterative implicit Crank-Nicolson type scheme is employed for temporal discretization.
-Spatial discretization relies on staggered grids: an Arakawa C grid in the horizontal and a Charney-Phillips grid in the vertical.
-Further details in this regard can be found in the existing literature (Girard et al., 2014; Husain and Girard, 2017).
-Global simulations in GEM utilize a Yin-Yang grid system, where two orthogonal and overlapping LAM grids are combined to construct the global grid (Qaddouri and Lee, 2011).
-Solutions from the GEM dynamical core are combined with tendencies from parameterized physics schemes that represent unresolved (or partially resolved) physical processes, such as radiation, convection, boundary layer turbulence, surface-atmosphere interactions, condensation, etc.
+The dynamical core of the GEM model solves the elastic Euler system of equations using an iteratively implicit semi-Lagrangian approach (Girard et al., 2014; Husain and Girard, 2017).
+The equations are first transformed from regular height coordinate to some form of a terrain-following coordinate in the vertical, denoted by $\zeta$ in the model, which for ECCC’s current operational NWP systems, is based on log-hydrostatic-pressure (Husain et al., 2021).
+The global domain is represented as a pair of overlapping limited-area grids in a Yin-Yang configuration (Qaddouri and Lee, 2011).
+Solutions from the GEM dynamical core are combined with tendencies from parameterized physics schemes (McTaggart-Cowan et al., 2019) that represent the diabatic, frictional and subgrid-scale processes.
 The parameterization schemes are executed in a sequential manner.
-All operational GEM-based NWP systems rely on a splitting-type mechanism for coupling the parameterized physics tendencies with the resolved dynamics.
-In this approach, physics-induced tendencies are applied as grid-point adjustments to the dynamics solution to obtain the complete solution for a model time step (Husain et al., 2019).
+Sequential coupling (Gross et al., 2018), also referred to as split-type coupling (Husain et al., 2019), is employed to combine the parameterized physics tendencies with the resolved dynamics to obtain the complete solution for a model time step.
 2.2 Global Deterministic Prediction System (GDPS)
-The version of the GDPS used in this study is based on its current operational configuration.
-It has a horizontal grid spacing of approximately 15 km.
-In the vertical, GDPS employs 84 prognostic vertical levels with the topmost level at approximately 0.1 hPa.
-The first momentum and thermodynamical levels are approximately 20 m and 10 m away from the surface, respectively.
-It uses a time-step length of 450 s.
-In terms of the underlying physics parameterizations, GDPS has undergone a major modernization in the recent past (McTaggart-Cowan et al., 2019).
-This includes updated versions of schemes for radiative transfer, blocking, orography gravity waves and non-orographic gravity waves.
+The version of the GDPS used in this study is based on its current operational configuration (McTaggart-Cowan et al., 2019; ECCC, 2024).
+It has a horizontal grid spacing of approximately 15 km along with a time-step length of 450 s.
+In the vertical, GDPS employs 84 prognostic levels with the top approximately at 0.1 hPa.
 Although operational GDPS is coupled with the NEMO ocean model (Smith et al., 2018), all the experiments for this study were conducted with atmosphere-only configurations.
-As the impact of coupling is usually modest for 10-day forecasts (McTaggart-Cowan et al., 2019), conclusions drawn from this study are expected to generally hold in the presence of atmosphere-ocean coupling.
+This simplification reduces both complexity and computational cost, and in the past, has been found to provide reliable guidance for the behavior of the full operational system.
 3 The AI-based Prediction System
 3.1 The GraphCast Model
-The GraphCast (GC) model from Google-DeepMind is an AI-based weather emulator which has been trained to emulate ECMWF’s ERA5 reanalyses.
-It requires the state of the atmosphere at times $t$ and $t-\Delta t_{GC}$ to produce a forecast valid at $t+\Delta t_{GC}$, where $\Delta t_{GC}$=6 hr.
-Although the model can be employed auto-regressively $N$ times to generate predictions valid at $t+N\Delta t_{GC}$, it is trained to minimize error for up to 12 recursive forecast steps (from 6 hr to 3 days) by considering ERA5 as the ground truth.
+The GraphCast (GC) model from Google DeepMind is based on AI, and like other global AI weather models, has been trained to emulate ECMWF’s ERA5 reanalyses (Lam et al., 2022, 2023).
+It requires the state of the atmosphere at times $t$ and $t-\Delta t_{GC}$ to produce a forecast valid at $t+\Delta t_{GC}$, where $t$ is the initialization time and $\Delta t_{GC}$=6 hr.
+Although the model can be employed auto-regressively $N$ times to generate predictions valid at $t+N\Delta t_{GC}$, it is trained to minimize the mean squared error (weighted by vertical levels) for up to 12 recursive forecast steps (from 6 hr to 3 days) by considering ERA5 as the ground truth.
 The objective of such a training is to produce forecasts with reasonable skill for up to day 10.
@@
 These time-dependent forcing terms are computed analytically.
-Furthermore, the input system requires some stationary input features like the land-sea mask, surface geopotential as well as information on the local longitude and latitude of its input grid.
-Pre-trained weights of GC (learned features of the GNNs) are available from Google-DeepMind for two different sets of pressure levels.
+Furthermore, the input system requires some stationary input features, such as land-sea mask, surface geopotential as well as information on the local longitude and latitude of its input grid.
+Pre-trained weights of GC (learned features of the GNNs) are available from Google DeepMind for two different sets of pressure levels.
 The version with higher vertical resolution has 37 pressure levels with the top at 1 hPa, while the coarser-resolution version has only 13 levels with the top at 50 hPa (Lam et al., 2022).
@@
 All GC predictions utilized in this study are based on this experimental system.
-Based on the evaluation of GC predictions with the GDPS analyses as input, the 13-level version of GC (GC-13L) outperforms the 37-level version (GC-37L) for most variables, at most levels and over most regions of the world.
-This is presumably due to the fact that stratospheric specific humidity in the GDPS analyses generally differs from ERA5 by an order of magnitude creating large imbalances when 37-level GC (with multiple levels in the stratosphere) is initialized with the GDPS analyses.
+Based on the evaluation of GC predictions with the GDPS analyses as input, the 13-level version of GC outperforms the 37-level version for most variables, at most levels and over most regions of the world.
+Although a detailed analysis of this behavior is beyond the scope of this work, it is believed to be related to systematic differences in stratospheric moisture values between the initializing GDPS analyses and ERA5 reanalyses used to train GC.
 The previously mentioned additional fine-tuning of GC-13L by DeepMind could also have contributed to its superior performance.
-This study is, therefore, entirely based on the GC-13L and, henceforth, GC is used to refer to GC-13L.
+This study is therefore, entirely based on the 13-level version of the system and the acronym ‘GC’ is henceforth used to refer to this version.
 4 Relative performances of the GDPS and GC forecasts
-4.1 Experimental setup
 Ten-day forecasts from GC and GDPS are initialized using the GDPS analysis every 36 h, leading to 60 cases for boreal winter 2022 (from 01 December 2021 at $0$ UTC to 27 February 2022 at 1200 UTC) and 62 cases for boreal summer 2022 (from 01 June 2022 at $0$ UTC to 31 August 2022 at 1200 UTC).
-This version of GDPS, which includes data assimilation and forecasting components, became operational in June 2024 (ECCC, 2024).
-Henceforth, the GDPS experiment based on this configuration will be referred to as GDPS-CTL.
+The sequence of forecast integrations initialized from these dates using an atmosphere-only operational GDPS configuration is denoted GDPS-CTL.
+Predictions from the two systems, GDPS-CTL and GC, are first assessed for their ability to depict the full range of atmospheric circulations using variance spectra (section 4a).
+This analysis is aimed at ascertaining effective resolutions of the two models, which will help to fairly interpret their relative predictive skills when comparing against radiosonde observations later in section 4b.
 4.1 Verification with global power spectra
@@
 (2023)).
-The smoothing effect is well indicated by departures of 120-h GC forecasts’ global 250-hPa kinetic energy and 500-hPa temperature variance spectra from those of GDPS-CTL forecasts and the analyses (Fig.
+The smoothing effect is well indicated by departures of 120-hr GC forecasts’ global 250-hPa kinetic energy and 500-hPa temperature variance spectra from those of GDPS-CTL forecasts and the analyses (Fig.
 1).
-GraphCast’s spectra are steeper, leading to a notable variance deficiency at wavenumbers as low as 30, corresponding to lower synoptic and sub-synoptic scales ($<$1500 km).
-However, both GDPS-CTL forecasts and analyses much more closely follow the $k^{-3}$-law, pertaining to the 3-dimensional quasi-geostrophic flow (Charney, 1971; Skamarock et al., 2014), which indicates that GC slopes are not realistic.
-Hence, the effective resolution of GC, defined as the highest wave number at which the atmospheric variability is still adequately represented, is considerably lower than its nominal resolution.
-Conversely, the effective resolution of the classical NWP models is typically about 6$\sim$8 times their nominal resolution, i.e., grid spacing (Skamarock, 2004).
-With the current operational configuration, the nominal resolution for GDPS-CTL is 0.1375°$\degree$, leading to an effective resolution of around 90-100 km.
-Figure 1: Global (a) kinetic energy power spectrum of 250-hPa horizontal wind and (b) variance spectrum of 500-hPa temperature after 120 h of integration for the GDPS analyses (black), GDPS-CTL (red) and GC (red), averaged over 60 cases of boreal winter 2022.
+For the fine scales, GraphCast’s spectra are steeper, leading to a notable variance deficiency at wavenumbers as low as 30, corresponding to lower synoptic and sub-synoptic scales ($<$1500 km).
+More importantly, both GDPS-CTL forecasts and analyses much more closely follow the expected $k^{-3}$ spectral slope pertaining to the 3-dimensional quasi-geostrophic flow (Charney, 1971; Skamarock et al., 2014), which indicates that GC slopes are not realistic.
+Hence, the effective resolution of GC, defined as the highest wave number at which the atmospheric variability is still adequately represented, is considerably lower than its nominal resolution, i.e., grid spacing.
+Conversely, the effective resolution of the classical NWP models is typically about 6$\sim$8 times their nominal resolution (Skamarock, 2004).
+With the current operational configuration, the nominal resolution for GDPS-CTL is 0.1375°$\degree$, leading to an approximate effective resolution of 90$\sim$120 km (McTaggart-Cowan et al., 2024).
+Figure 1: Global (a) kinetic energy power spectrum of 250-hPa horizontal wind and (b) variance spectrum of 500-hPa temperature for the GDPS analyses (black) as well as 120-hr lead time of GDPS-CTL (blue) and GC (red), averaged over 60 cases of boreal winter 2022.
 The dashed line indicates the -3 slope.
+The 5%–95% confidence interval is presented with light shadings (not discernable for the given logarithmic scale).
 It is important to note that the stationary effects of orography, land-sea contrast and large latitudinal climatological variations tend to be much less affected by smoothing in GC than the less predictable transient anomalies.
 It is therefore useful to introduce the transient spectral amplitude ratio, defined as follows:
-$\gamma=\left(\frac{\overline{\sigma_{x^{\prime}x^{\prime}}}}{\overline{\sigma_{y^{\prime}y^{\prime}}}}\right)^{1/2},$ (1)
-where $\sigma_{x^{\prime}x^{\prime}}$ and $\sigma_{y^{\prime}y^{\prime}}$ are the power spectra of the forecast’ and analysis’ transient-eddy components, respectively, and the overbar denotes averaging over the cases for a given season.
+$\gamma(n)=\left(\frac{\overline{\sigma_{x^{\prime}x^{\prime}}(n)}}{\overline{\sigma_{y^{\prime}y^{\prime}}(n)}}\right)^{1/2},$ (1)
+where $\sigma_{x^{\prime}x^{\prime}}$ and $\sigma_{y^{\prime}y^{\prime}}$ are real-valued non-negative functions of the total spherical wavenumber $n$, representing global spectral variances of the transient-eddy components associated with the forecast and analysis, respectively.
+The overbar denotes averaging over the cases for a given season.
 The forecast transient eddies are defined as $x^{\prime}=x-\overline{x}$, and similarly for the analysis.
-Figure 2 shows the evolution of $\gamma$ with lead time for GC and GDPS-CTL forecasts, against the GDPS analysis.
-While $\gamma$ remains close to 1 for all lead times for GDPS-CTL, indicating that its effective resolution does not change during the integration, for GC it rapidly decreases, leading to an effective resolution as low as 1000 km for 24-h forecasts.
-The smoothing effect further reduces the effective resolution to 2500$\sim$2700 km before it saturates around forecast days 3–5.
-It is worth noting here that GC still resolves scales smaller than these limits, but only partly ($10<\gamma<$ 1).
-While designing any strategy to leverage GC predictions to improve GEM, it would be prudent to not force the GEM model to follow GC for these variance-deficient scales in order to not adversely affect GEM’s effective resolution.
-This will be revisited in Section 5, where the concept of spectral nudging and its implementation in GEM are discussed.
-Figure 2: Spectral (a, b) amplitude ratio $\gamma$ and (c, d) coherence $\rho$ for the global (a, c) kinetic energy spectrum of 500-hPa horizontal wind transient-eddy deviations and (b, d) variance spectrum of 500-hPa temperature transient-eddy deviations, or GDPS-CTL (blue) and for GC (red), after (dark to light) 24, 72, 120 and 240 h of integration.
+The optimal value of $\gamma$ is 1 for the range of scales that are adequately represented in the reference analysis.
+In general, for any given length scale (or wavenumber), a value of $\gamma$ larger than 1 would imply that the forecast has more variability than the analysis, whereas a value smaller than 1 would indicate variance deficiency in the forecast.
+Figure 2 shows the evolution of $\gamma$ with lead time for GC and GDPS-CTL forecasts against the GDPS analysis.
+While $\gamma$ remains close to 1 for all lead times for GDPS-CTL, indicating that its effective resolution does not change during the integration, for GC it rapidly decreases, leading to an effective resolution as low as 1000 km for 24-hr forecasts.
+The smoothing effect further reduces the effective resolution to 2500$\sim$2700 km before it saturates around forecast days 3$\sim$5.
+It is worth noting here that GC still resolves scales smaller than these limits, but only partly ($\gamma<1$).
+The impact of this variance-deficiency on optimal usage of GC guidance in a hybrid system will be discussed in more detail in section 5.
+Figure 2: Spectral (a, b) amplitude ratio $\gamma$ and (c, d) coherence $\rho$ for the global (a, c) kinetic energy spectrum of 500-hPa horizontal wind transient-eddy deviations and (b, d) variance spectrum of 500-hPa temperature transient-eddy deviations, for GDPS-CTL (blue) and for GC (red), after 24, 72, 120 and 240 hr of integration (darker to lighter shades of color).
 Spectra are averaged over 60 cases of boreal winter 2022.
+Light shadings indicate the 5%–95% confidence interval.
 Figure 2 also displays the spectral coherence defined as follows:
-$\rho=\frac{\overline{\sigma_{x^{\prime}y^{\prime}}}}{\left(\overline{\sigma_{x^{\prime}x^{\prime}}}\;\overline{\sigma_{y^{\prime}y^{\prime}}}\right)^{1/2}}\;,$ (2)
-where $\sigma_{x^{\prime}y^{\prime}}$ is the transient-eddy co-spectrum between the GDPS-CTL or GC forecasts and the GDPS analyses.
+$\rho(n)=\frac{\overline{\sigma_{x^{\prime}y^{\prime}}(n)}}{\left(\overline{\sigma_{x^{\prime}x^{\prime}}(n)}\;\overline{\sigma_{y^{\prime}y^{\prime}}(n)}\right)^{1/2}}\;,$ (2)
+where $\sigma_{x^{\prime}y^{\prime}}$ is a real-valued function of the total spherical wavenumber $n$, representing the transient-eddy covariance spectrum between the GDPS-CTL (or GC forecasts) and the GDPS analyses.
+Spectral coherence can be interpreted as a scale-dependent spatial correlation coefficient with an optimal value of 1, which would imply a perfect spatial correlation between forecast and analysis transient-eddy anomalies.
+Conversely, a value of 0 would imply a total de-correlation.
+By integrating the individual terms on the right-hand side of Eq.
+2 across entire spectrum of wavenumbers, one may obtain a quantity which is analogous to the anomaly correlation, but in space.
 Since the same GEM model is used to produce both the GDPS forecasts and backgrounds for the analysis steps, they may have common systematic errors, especially at shorter lead times.
-This may unfairly penalize GC, which is trained on ECMWF-ERA5 reanalyses.
-Despite that, Fig.
-2 clearly indicates an improved spectral coherence with the GDPS analyses for a large range of scales – from synoptic to planetary, especially at longer lead times.
-This suggests that forcing GEM toward GC has potential for improving the predictive skill of its transient-eddy anomalies.
+Even though this penalizes the ERA5-trained GC, Figs.
+2c-d indicate that GC long-lead inferences are superior to the equivalent GDPS predictions over a broad range of synoptic and planetary scales.
 4.2 Verification against radiosondes
@@
 (2021)).
-This also helps to get around the issue of inhomogeneity of the observing network.
-Since GC was primarily trained with the ERA5 reanalysis and then fine-tuned on the ECMWF-HRES analysis, this likely introduces some dependency on the ECMWF analysis systems.
-To ensure that the dataset used to verify the forecasts is completely independent from both models as well as the GDPS data assimilation system, it was opted to rely mainly on observations from the global radiosondes network.
-Observations of zonal wind ($U$), geopotential height ($Z$), temperature ($T$) and dewpoint depression ($T–T_{d}$) at mandatory pressure levels between 1000 and 100 hPa will therefore be used as reference for verification.
-Two flavors of the verification against radiosonde observations are performed: (1) verification at full resolution, with the GDPS-CTL and GC forecasts at their native resolution, and (2) verification at low resolution –- corresponding to the effective resolution of GC – wherein both the GDPS-CTL and GC forecasts are filtered to remove the variability at scales not adequately resolved by GC.
-The fact that GC has a much lower effective resolution than GDPS-CTL would otherwise provide an unfair advantage to GC by reducing the so-called “double penalty” for misplacing the predicted patterns.
-Low-resolution verification against radiosondes also allows the study of how suitable GC forecasts are for large-scale spectral nudging in GEM.
+Conversely, since GC was primarily trained with the ERA5 reanalysis and then fine-tuned on the ECMWF-HRES analysis, this likely introduces some dependency on the ECMWF analysis systems.
+To ensure that the dataset used to verify the forecasts is as independent as possible from both models as well as the GDPS data assimilation system, it was opted to rely mainly on observations from the global radiosondes network.
+Observations of zonal wind ($U$), geopotential height ($Z$), temperature ($T$) and dewpoint depression ($T–T_{d}$) at mandatory pressure levels between 1000 and 100 hPa are therefore used as reference for verification.
+Two flavors of verification against radiosonde observations are performed based on the results shown in the previous section: (1) verification at full resolution, with the GDPS-CTL and GC forecasts at their native resolution, and (2) verification at low resolution –- corresponding to the effective resolution of GC – wherein both the GDPS-CTL and GC forecasts are filtered to remove the variability at scales not adequately resolved by GC.
+The fact that GC has a much lower effective resolution than GDPS-CTL would otherwise provide an unfair advantage to GC by reducing the so-called “double penalty” for misplacing the predicted patterns (Mass et al., 2002).
 The filter applied here is a spectral filter based on spherical harmonics.
@@
 Note also that the same filter configuration is applied to GC and GDPS-CTL (more details on this filter can be found in Appendix A).
-Figure 3: Verification for 120-h forecasts from GC (red) and GDPS-CTL (blue) at their full (top row: a–e) and low (bottom row: e–h) resolutions against global radiosonde observations, averaged over 60 cases in boreal winter.
+Figure 3: Verification for 120-hr forecasts from GC (red) and GDPS-CTL (blue) at their full (top row: a–d) and low (bottom row: e–h) resolutions against global radiosonde observations, averaged over 60 cases in boreal winter.
 The error standard deviation (STDE; solid lines) and bias (mean forecast minus observation; dashed lines) are shown for (a,e) zonal wind, (b,f) geopotential heights, (c,g) temperature and (d,h) dewpoint depression.
@@
 4.2.1 At full resolution
-Comparing the forecasts at their native resolution to the radiosonde observations reveals that for most variables and lead times, and for levels between 1000 and 200 hPa, GC forecasts are much closer to the observations than GDPS-CTL.
-Vertical profiles of error averages (bias) oriented as forecast minus observation and error standard deviations (STDE), calculated over the global domain for 120-h forecasts for boreal winter, highlight the typical verification differences between GC and GDPS-CTL (Fig.
+Comparing the forecasts at their native resolution against radiosonde observations reveals that for most variables and lead times, and for levels between 1000 and 200 hPa, GC forecasts are much closer to the observations than GDPS-CTL.
+Vertical profiles of error averages (bias) oriented as forecast minus observation and error standard deviations (STDE), calculated over the global domain for 120-hr forecasts for boreal winter, highlight the typical verification differences between GC and GDPS-CTL (Fig.
 3a–d).
-The differences between GC and GDPS-CTL in terms of biases are mixed, except for temperature during boreal winter, where clear improvements can be seen with GC.
+The differences between GC and GDPS-CTL in terms of biases are mixed, except for temperature, where clear improvements can be seen with GC.
 In terms of STDE, however, GC clearly shows large improvements over GDPS-CTL at almost every level, and particularly for zonal wind and dewpoint depression.
@@
 Positive values of $f_{RMSE}$ thus imply improved forecasts from GC measured by the RMSE against the radiosonde observations, while negative values represent the opposite.
-The forecast quality index $f_{RMSE}$ was computed every 24 h for each selected variable and level.
+The forecast quality index $f_{RMSE}$ was computed every 24 hr for each selected variable and level.
 The results for each variable were then vertically averaged for three layers: low-level (from 1000 to 850 hPa), mid-level (from 700 to 300 hPa) and high-level layer (from 250 hPa to 100 hPa).
-Table 1: Latitude and longitude bounds for the 6 subdomains considered in Figures 5, 11 and 13.
+Table 1: Latitude and longitude bounds for the 6 subdomains considered in Figures 5, 10 and 12.
 […]
@@
 (2022)).
-The region-wise variations in relative forecast performances can be examined by computing, as above, the forecast quality index $f_{RMSE}$ averaged over all variables and lead times for each vertical layer but, this time, for the 6 subdomains listed in Table 1 with their respective latitude and longitude limits.
+The region-wise variations in relative forecast performances can be examined by computing, as before, the forecast quality index $f_{RMSE}$ averaged over all variables and lead times for each vertical layer but, this time, for the 6 subdomains listed in Table 1 with their respective latitude and longitude limits.
 In low- and mid-level layers, GC forecasts show smaller RMSE in every subdomain in both seasons (Figs.
 5a and c).
-In the upper-level layer, GDPS-CTL performs better during each hemisphere’s respective winter.
+In the high-level layer, GDPS-CTL performs better during each hemisphere’s respective winter.
 However, GC’s poor performance during winter over the northern hemisphere is primarily originating from Europe.
@@
 At low resolution, the forecast improvements in terms of RMSE from GC over GDPS-CTL are significantly reduced for all variables, levels, lead times, seasons and geographical domains (Figs.
-3e–h and 4).
-This reduction in improvements at low resolution is particularly striking in the mid-level layer over the first 72 h (compare Figs 4a and b), leading to growing forecast improvements with time from GC that saturate or peak near 120 h (Fig.
+3e–h and 4b).
+This reduction in improvements at low resolution is particularly striking in the mid-level layer over the first 72 hr (compare Figs 4a and b), leading to growing forecast improvements with time from GC that saturate or peak near 120 hr (Fig.
 5b) compared to a relatively steady improvement at all lead times found with the full resolution (Fig.
 4a).
-This suggests that the improvements with GC for the first 72 h at full resolution largely originate from partial smoothing of the less-accurately predicted fine scales that reduces the double penalty.
-Degradations in GC, seen at full resolution over Northern Hemisphere in boreal winter (mostly coming from Europe) and Southern hemisphere in boreal summer, are enhanced at low resolution (compare Figs 5a, c, and 5b, d).
+This suggests that the improvements with GC for the first 72 hr at full resolution largely originate from partial smoothing of the less-accurately predicted fine scales that reduces the double penalty.
+Degradations in GC, seen at full resolution over Northern Hemisphere during boreal winter (mostly coming from Europe) and Southern hemisphere during boreal summer, are enhanced at low resolution (compare Figs 5a,c, and 5b,d).
 There are also notable degradations in the mid- and low-level layers over Asia in summer (Fig.
-6d).
+5d).
 The average reduction factor for $f_{RMSE}$ from full- to low-resolution verification is roughly between 2 to 3 (compare the left and right columns in Fig.
@@
 Nevertheless, large scales of GC forecasts are still, on average, significantly closer to the radiosonde observations than those from GDPS-CTL in the mid- and low-level layers.
-This suggests that nudging these large-scale components of key GEM state variables toward GC could result in substantial improvements over GDPS-CTL.
+This suggests that the GC forecasts contain useful large-scale information that could substantially improve guidance if successfully incorporated into a hybrid NWP-AI system.
 5 Spectral nudging in GEM
 5.1 The concept
-Spectral nudging operates by gradually directing the model-predicted atmospheric large scales towards those found in some reference data which is considered to have a more accurate representation of those scales.
-This approach fundamentally differs from grid nudging or indiscriminate nudging, as it only targets a predefined range of scales.
+Spectral nudging works by directing the model-predicted atmospheric large scales toward a more skillful reference (von Storch et al., 2000; Husain et al., 2014).
+This approach fundamentally differs from grid nudging or indiscriminate nudging (Liu et al., 2012), as it only targets a predefined range of scales.
 Even though an NWP model may employ grid-based spatial discretizations, the nudging increments at a given model level are computed by decomposing the model predictions and the reference fields in a spectral space, retaining only the target scales.
@@
 5.2 Implementation in GEM
-With the split-adjustment-type dynamics-physics coupling in GEM, the adiabatic dynamical core first solves the prognostic dynamical equations to obtain an intermediate state of the atmosphere.
+With the split-type dynamics-physics coupling in GEM, the adiabatic dynamical core first solves the prognostic dynamical equations to obtain an intermediate state of the atmosphere.
 In the absence of spectral nudging, this interim dynamics solution serves as the input for the physical parameterization schemes, which compute the physics-induced tendencies.
@@
 However, when spectral nudging is applied, the large scales in the solutions for a selected set of prognostic variables from GEM dynamics are first nudged toward the reference (here, from GC).
-Subsequently, the nudged solutions are fed to the physics schemes to compute the physics-induced tendencies before the eventual dynamics-physics coupling, following the regular model procedure.
+Subsequently, the nudged solutions are fed to the physics schemes to compute the physics-induced tendencies before the eventual dynamics-physics coupling.
 In mathematical form, the nudging step at a given model vertical level, i.e., a constant-$\zeta$ surface, can be expressed as
 $F_{nudge}=F_{GEM}+\omega\left[F_{GC}-F_{GEM}\right]_{LS},$ (4)
-where $F_{GEM}$ is a prognostic variable predicted by GEM dynamics, $F_{GC}$ is the corresponding prediction from GC valid at the same time, $\omega$ is the nudging relaxation factor such that $(0\leq\omega\leq 1)$, and $F_{nudge}$ is the nudged solution.
-The subscript $LS$ in the above equation refers to some user-defined large scales targeted by the nudging mechanism.
+where $F_{GEM}$ is a prognostic variable predicted by GEM dynamics, $F_{GC}$ is the corresponding prediction from GC vertically interpolated to the $\zeta$ level valid at the same time, $\omega$ is the nudging relaxation factor such that $(0\leq\omega\leq 1)$, and $F_{nudge}$ is the nudged solution.
+The subscript $LS$ in the above equation refers to some user-defined large scales targeted by the nudging mechanism, and is discussed later in more detail.
 Through simple rearrangement of the terms on the right-hand-side of Eq.4, it can be shown that the large scales in $F_{nudge}$ are indeed a weighted average of those from $F_{GEM}$ and $F_{GC}$ (with $\omega$ being the weighting parameter), whereas the small scales from $F_{GEM}$ are retained entirely in $F_{nudge}$.
@@
 $\omega=\frac{\beta(\zeta)}{\tau(t)}\Delta t,$ (5)
-where $\beta(\zeta)$ defines the nudging vertical profile with respect the $\zeta$-coordinate, $\tau(t)$ represents the nudging relaxation time scale , and $\Delta t$ denotes the model time-step length (in the same unit as $\tau$).
+where $\beta(\zeta)$ defines the nudging vertical profile with respect to the $\zeta$-coordinate, $\tau(t)$ represents the nudging relaxation time scale , and $\Delta t$ denotes the model time-step length.
 A vertically-variable $\beta$ allows for different nudging strengths at different model levels.
@@
 To nudge a prognostic variable $F$ at any model level, it is first essential to spectrally decompose $(F_{GC}-F_{GEM})$ so that a filter can be applied in the spectral space to retain only the desirable scales.
-Spherical harmonics-based spectral decomposition would be the ideal choice for such a filter.
+For global simulations, such a filter should ideally employ a spherical harmonics-based spectral decomposition.
 However, this would require multiple transformations of model solutions between the model’s Yin-Yang grids and an intermediate global Gaussian grid, resulting in significant increase in the computational cost.
 As the present study is primarily of a proof-of-concept nature, in order to keep the approach simple, the spectral filter for nudging was instead chosen to be based on discrete cosine transform (DCT) (Denis et al., 2002), and the filtering is performed separately on the Yin and Yang grids.
-Additionally, nudging is also separately applied to the Yin and Yang grid solutions after every dynamics sub-step.
-Large scales that are retained with spectral nudging are defined by users through configurable parameters.
+Additionally, nudging is separately applied to the Yin and Yang grid solutions after every dynamics sub-step.
+It is important to note that, for the range of cut-off scales considered in this study, the response of the DCT-based filter (applied separately to the Yin-Yang grids) is found to be closely equivalent to the response of a spherical-harmonics-based filter applied to a global Gaussian field (see Appendix B).
+Large scales that are retained with spectral nudging are defined through configurable parameters.
 Instead of an abrupt cut-off for the target scales, which may lead to Gibbs phenomena (Sardeshmukh and Hoskins, 1984), a soft cut-off is employed within the DCT-based spectral filter.
@@
 A series of systematic sensitivity experiments was carried out to identify an optimal spectral nudging configuration for GDPS.
-All subsequent results presented in the following section are based on this optimal configuration which, from this point onward, will be referred to as GDPS-SN.
-It is important to note that spectral nudging is currently only applied within the forecasting component of GDPS, i.e., GDPS-CTL, GDPS-SN and GC are all initialized with the same analyses.
-Some key aspects of this optimal setup are discussed below.
+Results presented in section 6 are based on this optimal configuration which is referred to as GDPS-SN.
+Note that spectral nudging is currently only applied within the forecasting component of GDPS, i.e., GDPS-CTL, GDPS-SN and GC are all initialized with the same analyses.
+Some key aspects of the optimal nudging configuration are discussed below.
 5.3.1 Nudged variables
 Nudging is only applied to the $u$-$v$ components of wind and virtual temperature.
-While nudging specific humidity can improve the global bias of temperature and humidity in the boundary layer, the computational cost outweighs the benefits.
-Moreover, nudging specific humidity tends to negatively impact tropical cyclone intensity during the tested period, making it less desirable.
+While nudging specific humidity can improve the global bias of temperature and humidity in the boundary layer, the computational cost (see section 66.6) outweighs the benefits.
+Moreover, sensitivity experiments have revealed a negative impact of specific humidity nudging on tropical cyclone intensity (not shown), making it less desirable.
 5.3.2 Nudging vertical profile
@@
 This profile involves no nudging in the boundary layer (below 850 hPa) and the stratosphere (above 250 hPa).
-Relatively weaker performance of GC in the stratosphere – as has been mentioned earlier – is the reason for no stratospheric nudging.
-Within the boundary layer, nudging is avoided to prevent potential negative impacts arising from inconsistencies between the surface forcing in GDPS and GC training data (ERA5).
-Furthermore, GC-13 does not have sufficient vertical resolution in the boundary layer to provide a useful reference.
+Relatively weaker performance of GC in the stratosphere – as mentioned earlier – is the reason for no stratospheric nudging.
+Although nudging is generally avoided in the boundary layer to allow uninhibited generation of fine scales (e.g., Schubert-Frisius et al.
+2017), other studies have found the potential for significant near-surface skill improvement with nudging towards an accurate reference in the boundary layer (e.g., Husain et al.
+2014).
+However, in this study, nudging is avoided in the boundary layer for different reasons.
+First, the 13-level version of GC (used here) does not have sufficient vertical resolution in the boundary layer to provide a useful reference.
+Even in the absence of this issue, inconsistencies between the surface forcing in GDPS and GC training data (ERA5) may lead to potential negative impacts over regions with complex terrain.
+In addition, differences in mean boundary-layer state (moisture in particular) may result in adverse reactions from the physical parameterizations in GEM.
+Therefore, nudging in the boundary layer is not a viable option at this stage.
 […]
 5.3.3 Nudging length scales
-It has been already shown that the fine-scales in GC suffer from considerable smoothing, worsening at longer lead times (see Figs.
+It was shown earlier in section 4 that the fine-scales in GC suffer from considerable smoothing, worsening at longer lead times (see Figs.
 1 and 2).
 The implementation of spectral nudging in GEM allows for time-varying cut-off scales for nudging.
-However, the improvements with such an approach over a stationary cut-off scale are found to be negligible.
+However, improvements with such an approach over stationary cut-off scales are found to be negligible.
 Therefore, based on the comparison of spectra between GDPS and GC, the DCT-based filter for the optimal configuration utilized in this study is configured with $\lambda_{LS}$=2750 km and $\lambda_{SS}$=2250 km.
@@
 The scales in between are subjected to partial filtering (see Fig.
-20).
+19).
 This choice of nudging length scales corresponds to the effective resolution of GC beyond day 3 (see Figs.
@@
 However, selecting an optimal $\tau$ is not straightforward.
-First, it is important to note that GC inferences are only available every 6 hours and become increasingly inaccurate with longer lead times (as demonstrated by spectral coherence in Fig.
-2).
+First, it is important to note that GC inferences are only available every 6 hours.
 During model integration, when a valid GC inference is not available at certain time steps, an estimate is approximated through linear interpolation between two consecutive GC inferences around the time of interest.
 In this regard, Omrani et al.
-(2012) have argued that $\tau$ should not be less than $\tau_{a}$ which denotes the time interval between consecutive driving fields (here, GC inferences).
-Strong nudging with $\tau<\tau_{a}$ can lead to erroneous evolution of nudged fields.
-Furthermore, sensitivity experiments carried out during this study revealed considerable deterioration of model’s skill over complex terrain with $\tau<\tau_{a}$ (not shown).
-As a result, $\tau_{a}$ may safely be considered the lower bound for $\tau$.
+(2012) have argued that $\tau$ should not be less than $\tau_{a}$, which denotes the time interval between consecutive driving fields (here, GC inferences).
+Husain et al.
+(2014) showed that, with $\tau<\tau_{a}$, the impact of time interpolation error on the nudged fields may not be negligible.
+As a result, $\tau_{a}$ may be considered the lower bound for $\tau$.
 It is also crucial to ascertain an acceptable upper bound for $\tau$.
-Intuitively, as the accuracy of reference nudging data worsens with increasing forecast lead time, it might seem prudent to employ a large $\tau$.
-However, sensitivity experiments have revealed that too large of a nudging relaxation time ($\tau>24$ hr) can result in excessive undesirable smoothing in the nudged fields around the cut-off scales developing at day 3 or 4.
+Sensitivity experiments have revealed that too large of a nudging relaxation time ($\tau>24$ hr) can result in excessive undesirable smoothing in the nudged fields around the cut-off scales developing at day 3 or 4.
 Eventually, this smoothing affects all scales by day 10 of the forecast (not shown).
-This increased smoothing with overly large $\tau$, is caused by the averaging effect resulting from blending two forecasts, with $\tau=48h$ maximizing the effect.
-Based on the insights derived from the sensitivity tests, a nudging relaxation time of 12 hr is chosen to be optimal.
+This increased smoothing with overly large $\tau$, is caused by the averaging effect resulting from blending two forecasts, with $\tau=48$ hr maximizing the effect.
+Strong nudging with $\tau=\tau_{a}$ is also found to result in erroneous evolution of nudged fields emanating from inconsistencies between GEM and GC over complex terrain, as well as error originating from vertically interpolating GC to GEM levels.
+Therefore, based on the insights derived from the sensitivity tests, a nudging relaxation time of 12 hr is chosen to be optimal.
+An important point to consider in this context is that although GC inferences become increasingly inaccurate with respect to GDPS analyses at longer lead times, they still maintain higher spectral coherence (compared to GDPS-CTL) for the largest scales (see Fig.
+2c,d).
+Therefore, a constant value of $\tau$ during the entire integration is justified.
 6 Impact of large-scale spectral nudging
 6.1 Verification with global power spectra
-The power spectra of the GDPS-SN forecast variables remain approximately the same as in the GDPS-CTL (Fig.
-7), which essentially demonstrates that the new hybrid system can adequately resolve the full range of scales present in the GDPS-CTL, without significant smoothing.
+The power spectra of the GDPS-SN forecast variables is indistinguishable from GDPS-CTL (not shown), which essentially suggests that the new hybrid system can adequately resolve the full range of scales present in the GDPS-CTL, without significant smoothing.
 The resulting spectral amplitude ratio of transient-eddy anomalies, $\gamma$ (Eq.
-1), also has values very close to one for all prognostic variables, lead times and spatial scales for both GDPS-CTL and GDPS-SN.
-This provides a solid proof-of-concept for the spectral nudging approach (Fig.
-8).
-The only exception is a 10-percent reduction of $\gamma$ for scales between 2000 and 4000 km at lead times approaching day 10 (wavenumbers 10-20 in Figs.
-8a, b).
-This reduction is explained by GC having some variance deficiency at the corresponding range of scales (2000-2750 km) for longer lead times (beyond day 4).
-Sensitivity tests with time-varying cut-off length scales that more strictly follow GC’s time-evolving effective resolution, showed that this issue could be avoided, but at the expense of significantly reduced GDPS-SN forecast improvements with respect to the GDPS-CTL.
+1), also has values very close to one for all prognostic variables, lead times and spatial scales for both GDPS-CTL and GDPS-SN (Fig.
+7).
+This provides a solid proof-of-concept for the spectral nudging approach.
+The only exception is a 10-percent reduction of $\gamma$ for scales between 2000 and 4000 km at lead times approaching day 10 (wavenumbers 10–20 in Figs.
+7a, b).
+This reduction is explained by GC having some variance deficiency at the corresponding range of scales (2000–2750 km) for longer lead times (beyond day 4).
+Sensitivity tests with time-varying cut-off length scales that more strictly follow GC’s time-evolving effective resolution showed that this issue could be avoided, but at the expense of significantly reduced GDPS-SN forecast improvements with respect to the GDPS-CTL.
 This is caused by a weaker constraint being imposed on GDPS-SN towards GC at synoptic scales by such a configuration.
@@
 2) between the forecasts and GDPS analysis shows an improved skill of GDPS-SN over GDPS-CTL (Figs.
-8c, d).
+7c, d).
 These improvements are somewhat weaker than those seen with GC at shorter lead times (up to day 5), but become comparable beyond day 5 (Figs 2c, d).
 The scales relevant for improvements at shorter lead times are smaller than those included in the spectral nudging configuration.
-Nudging the smaller scales would therefore bring the coherence on par with GC but it would lead to variance deficiency, as discussed earlier.
-Figure 7: Same as in Fig. 1, but for the analysis, GDPS-CTL and GDPS-SN.
-Figure 8: Same as in Fig. 2, but for GDPS-SN and GDPS-CTL.
+Nudging the smaller scales would therefore bring the coherence on par with GC, particularly for short lead times, but it would result in variance deficiency, as discussed earlier.
+Figure 7: Same as in Fig. 2, but for GDPS-SN and GDPS-CTL. Additionally, for convenience of comparison, GC has been added in Fig. 7c and d.
 An illustrative example of the value added by the new hybrid system over GC is provided in Fig.
-9.
-Large-scale features of transient-eddy anomalies in GDPS-SN tend to be better (but not fully) aligned with GC than with GDPS-CTL, while containing similar level of fine-scale details as in GDPS-CTL.
-Figure 10 suggests that large-scale components of GDPS-SN forecast fields are preconditioned with the piloting AI model.
-Therefore, further improvements to its predictions have potential for similarly improving GDPS-SN.
-Figure 9: 10-day forecast of 700-hPa transient-eddy specific humidity anomalies, valid 22 February 2022 at 12z for (a) GDPS-CTL, (b) GC and (c) GDPS-SN.
+8.
+For temperature, which is explicitly nudged in GDPS-SN, the large-scale features of its transient-eddy anomalies tend to very closely resemble GC (Fig.
+8a–c).
+Even for specific humidity, a variable which is not directly constrained by nudging, the large scales are also found to be more aligned with GC than with GDPS-CTL (Fig.
+8d–f).
+This is however not surprising as it reinforces the findings of previous studies on spectral nudging (Husain et al., 2014).
+Regarding fine scales, Fig.
+8 clearly shows that GDPS-SN has similar level of variability as GDPS-CTL, which is not the case with GC as it clearly suffers from severe smoothing.
+Figure 8: 10-day forecast of 700-hPa transient-eddy anomalies for (a–c) temperature and (d–f) specific humidity, valid 22 February 2022 at 12z for GDPS-CTL (left), GC (middle) and GDPS-SN (right).
 6.2 Verification against radiosondes
 Comparison of forecasts at full resolution against the global radiosonde observations clearly indicates that GDPS-SN forecasts are closer to the observations than GDPS-CTL forecasts for basically all variables, lead times and levels, and over both seasons (Fig.
-10).
-The improvements gradually grow in the first 72 h and are remarkably constant with altitude, which contrasts with the strong vertical variations observed with GC in Section 4.
-This demonstrates that large-scale spectral nudging, despite being only applied in the mid-level layer for winds and temperature, is sufficient to impart similar improvements at both upper- and lower-level layers that are not constrained by spectral nudging.
+9).
+The improvements gradually grow in the first 72 hr and are remarkably constant with altitude, which contrasts with the strong vertical variations observed with GC in Section 4.
+This demonstrates that large-scale spectral nudging, despite being only applied in the mid-level layer for winds and temperature, is sufficient to impart similar improvements at both upper- and lower-level layers that are not directly constrained by spectral nudging.
 Compared to boreal winter (Fig.
-10a), the improvements from GDPS-SN are reduced by a factor of about 2 during boreal summer (Fig.
-10b), which is not surprising given that the benefits from GC over GDPS-CTL at the large scales are smaller by about the same factor for zonal wind and temperature in the mid-level (not shown).
+9a), the improvements from GDPS-SN are reduced by a factor of about 2 during boreal summer (Fig.
+9b), which is not surprising given that the large-scale improvements with GC over GDPS-CTL are smaller by about the same factor during boreal summer (not shown).
 In terms of region-wise performances, the GDPS-SN has smaller RMSE than GDPS-CTL in every subdomain except over Asia in the mid- and upper-levels during boreal summer (Figs.
-11c and d), which is relatively consistent with the degradations from GC over the GDPS-CTL at the large scales previously noted for the mid-levels over that part of globe during this season (see Fig.
+10c and d), which is consistent with the degradations from GC over the GDPS-CTL at the large scales in this region (Fig.
 5d).
@@
 10 and, again, in Fig.
-6).
+5).
 This does not necessarily imply that the improvements only originate from the large scales.
@@
 Figure 9: Same as in Fig.
-4, except that upward-pointing red triangles indicate a reduction of the RMSE in GDPS-SN with respect to the GDPS-CTL, whereas downward-pointing blue triangles indicate the opposite.
+4, except that upward-pointing red triangles indicate a reduction of the RMSE in GDPS-SN with respect to GDPS-CTL, whereas downward-pointing blue triangles indicate the opposite.
 Only verification using forecasts at full resolution is shown.
@@
 It is also interesting to compare GDPS-SN to GC.
-For the sake of fair comparison, this will only be done at low resolution.
-Predictions from GDPS-SN are closer to the radiosonde observations than GC in the upper and mid-level layers, except for dewpoint depression.
+For the sake of fair comparison, this is only done at low resolution.
+Predictions from GDPS-SN are closer to the radiosonde observations than GC in the high- and mid-level layers, except for dewpoint depression.
 Conversely, for the low-level layer, GC is closer to the observations, except for geopotential height (Fig.
-12).
+11).
 The degradations with respect to dewpoint depression, also observed in the mid-level layer where spectral nudging is active, suggest that further gains could likely be achieved by nudging the humidity field in GEM towards GC.
+Nudging of specific humidity, however, would only be viable once a fine-tuned version of GC that emulates GDPS analyses instead of ERA5 (and ECMWF-HRES) is available.
+This is expected to bring the mean moisture state of GC inferences closer to GDPS analyses and thereby reduce any error associated with potential adverse response from GEM physics.
 The results also suggest that extending spectral nudging to the boundary layer could be beneficial.
-This, however, would necessitate an AI model with more vertical levels in the boundary layer and fine-tuned to emulate the GDPS analyses rather than ECMWF-ERA5, for improved consistency with the GEM model’s lower-boundary forcings.
+This, however, would necessitate an AI model with more vertical levels in the boundary layer (as mentioned earlier) and fine-tuned to emulate the GDPS analyses rather than ECMWF-ERA5, for improved consistency with the GEM model’s lower-boundary forcings.
 The small but statistically significant GDPS-SN improvements over GC in the mid-level layer for temperature and zonal winds, two variables that are nudged toward GC, that grow with the integration time, are somewhat surprising (Fig.
-12).
-This could be due to a positive feedback mechanism between the large and small scales in GEM.
-The averaging effect caused by the blending of two forecasts due to a long relaxation time, as previously discussed in Section 5, may also be partly responsible for these improvements.
-Answering this question is, however, deemed outside the scope of this paper.
+11).
+The use of a nudging relaxation time $\tau=12$ hr implies that the large-scales in GDPS-SN are not too tightly constrained, which is visible in Fig.
+8.
+This likely allows for some large-scale improvements through upscale propagation of information from an improved representation of the fine scales.
+However, this choice of $\tau$ also leads to a small averaging effect caused by the blending of two forecasts, as previously discussed in Section 5.
+To a lesser extent, this may also have contributed to these improvements.
 Finally, the overall region-wise comparison of GDPS-SN and GC forecasts at low resolution (Fig.
-13) shows that the results that were observed over the globe hold also relatively well for each of the subdomains.
-The fact that spectral nudging is applied only in the mid-level layer targeting only the synoptic scales, it considerably limits the ability of GDPS-SN to fully leverage the strengths of GC for seasonal periods and over regions where mesoscale near-surface conditions play a major role in determining the weather.
-This may explain the degradations seen over the tropics.
+12) shows that the skill of GDPS-SN is either slightly better than or comparable to GC, with the tropical boundary layer being an exception.
+Fig.
+5 reveals that, compared to GC, GDPS-CTL large scales have substantially reduced skill in the tropics, which is slightly improved with GDPS-SN (see Fig.
+10).
+To leverage the full potential of GC in the tropics it appears to be necessary to extend nudging in the boundary layer, which is not viable at this stage for reasons explained earlier in this section.
 Figure 12: Same as in Fig.
@@
 6.3 Verification against ECMWF analyses
-The comparison of GDPS-SN and GDPS-CTL forecasts against ECMWF-HRES analyses leads to very much similar results as the comparison against radiosonde observations, with general improvements for all variables, lead times and pressure levels (not shown).
+The comparison of GDPS-SN and GDPS-CTL forecasts against ECMWF-HRES analyses leads to very similar results as the comparison against radiosonde observations, with general improvements for all variables, lead times and pressure levels (not shown).
 Figure 13 provides an example of improvements in anomaly correlation coefficient (ACC).
-For 500-hPa geopotential heights, the ACC improvements over the Northern Hemisphere region at forecast day 7 are equivalent to an 18-h increase in predictability during boreal winter and to an 8-h increase in summer.
-The corresponding ACC improvements over the Southern Hemisphere region are about 2 hours during austral winter and even exceed 24 hours during austral summer.
-There are also notable improvements in zonal wind ACC in the tropics that are equivalent to 24 hours or more (e.g., see Fig.
-14b,e).
+For 500-hPa geopotential heights, the ACC improvements over the Northern Hemisphere region at forecast day 7 are equivalent to an 13-hr increase in predictability during boreal winter and to an 8-hr increase in summer.
+The corresponding ACC improvements over the Southern Hemisphere region are about 3 hours during austral winter and 22 hours during austral summer.
+There are also notable improvements in zonal wind ACC in the tropics that are equivalent to 30 hours or more (e.g., see Fig.
+13b,e).
 The strength of GC in improving the ACC around day 5 (and beyond) is thus well leveraged by the hybrid system.
+Note that GC has not been included in this figure as its ACC is strongly influenced by the excessive fine-scale smoothing.
+To fairly compare GC against the two versions of GDPS, it is necessary to do that at low resolution (as in section 4b).
+However, based on spectral coherence presented in Fig.
+7c–d, low-resolution ACC of GC is expected to be comparable to GDPS-SN.
 Figure 13: Anomaly correlation coefficient against the ECMWF-HRES analyses with GDPS-CTL (blue) and GDPS-SN (red) for (a,d) northern hemisphere 500-hPa geopotential height, (b,e) tropical 850-hPa zonal wind, and (c,f) southern hemisphere 500-hPa geopotential height.
 Results are presented for 60 cases of boreal winter 2022 (left column: a–c) and 62 cases of boreal summer 2022 (right column: d–f).
+Red and blue markers denote statistically significant differences in favor of GDPS-SN and GDPS-CTL, respectively.
 No marker at a pressure level implies that the null hypothesis, stating that the statistics of the two samples are the same, cannot be rejected based on the 95th percentile.
+Grey arrows with printed numbers depict gains in predictability in terms of forecast hours.
 6.4 Verification against surface observations
 For verification at the surface level, combined SYNOP, METAR and Surface Weather and marine OBservations (SWOB; available only over Canada and distributed by ECCC) data were used for surface pressure, screen-level temperature, dewpoint depression and anemometer-level wind speed.
-For precipitation, 24-h accumulation ground observations are subject to collection and quality control by the CAnadian Precipitation Analysis system (CaPA; Lespinas et al.
-(2015)).
+For precipitation, ground observations of 24-hr accumulation, subject to collection and quality control by the CAnadian Precipitation Analysis system (CaPA; Lespinas et al.
+2015), are used.
 Any data pertaining to stations with an altitude difference larger than 100 m with respect to GDPS orography, along with any observed wind speed below 1.5 $\text{m s}^{-1}$, are excluded.
@@
 Diamond symbols denote selected percentiles with their values printed.
-Shadings denote 95% confidence bands for the inverse cumulative distribution functions, based on the Kolmogorov-Smirnov statistic.
+Light shadings indicate the 5%–95% confidence interval for the inverse cumulative distribution functions, based on the Kolmogorov-Smirnov statistic.
 Comparison of GDPS-SN and GDPS-CTL using the forecast quality index (Eq.
 3) is summarized in Fig.
-15.
+14.
 Guidance by GDPS-SN is generally improved for all variables and regions.
@@
 The ability of the GPDS-SN to represent high-impact weather, such as high wind speeds and extreme temperatures, is not significantly different from the GDPS-CTL but, at the same time, it is substantially better than GC, which provides a strong example of the value added by the hybrid system over GC (Fig.
-16).
-For example, the distribution of late-afternoon wind speeds over North America provided by GC shows a considerable shift toward lower values, which increases at high percentiles of the distribution, resulting in a -4 $\text{m s}^{-1}$ bias at the 99.9th percentile (Figs.
+15).
+For example, the distribution of late-afternoon wind speeds over North America provided by GC shows a considerable shift toward lower values, which increases at high percentiles of the distribution, resulting in a -3 $\text{m s}^{-1}$ bias around ${}^{\text{th}}$ percentile (Figs.
 15 a, b).
-GDPS-CTL also has a negative bias but affecting only higher percentiles and, more importantly, GDPS-SN closely follows GDPS-CTL, with little or no sign of deceleration at extreme wind speeds for both seasons.
+GDPS-CTL also has a negative bias but affecting only high percentiles and, more importantly, GDPS-SN closely follows GDPS-CTL, with little or no sign of deceleration at extreme wind speeds for both seasons.
 GraphCast also suffers a 2-4 K cold bias at mid-to-high temperature percentiles over North America in summer, which has no impact on the GDPS-SN; its distribution aligns very well with the observations.
@@
 Spectral nudging also significantly improves boreal winter precipitation forecast guidance over North America for a large range of thresholds from 1 to 50 mm day-1, mostly by reducing the false alarm ratio, with no reduction of probability of detection, which results in a reduced moist frequency bias and an improved equitable threat score (Fig.
-17).
+16).
 These improvements seem to be mostly generated via an improved prediction of synoptic-scale flow.
-There is, however, not much of an impact of spectral nudging on summer precipitation over North America.
-Improving precipitation in summer, when weak synoptic forcings are prevalent, may necessitate the application of spectral nudging in the boundary layer.
-Figure 17: Quantitative precipitation forecast verification for day 5 assessed by 24-hr precipitation threshold (accumulations between 108 and 132 h of integration) with the (a) equitable threat score, (b) frequency bias, (c) probability of detection, and (d) false alarm ratio.
+There is, however, not much of an impact of spectral nudging on summer precipitation over North America (not shown).
+During summer, when weak synoptic forcings are prevalent, nudging only large scales is not expected to considerably improve precipitation forecasts, although extending nudging to the boundary layer may lead to some improvements by mitigating model biases.
+Figure 16: Quantitative precipitation forecast verification for day 5 assessed by 24-hr precipitation threshold (accumulations between 108 and 132 hr of integration) with (a) equitable threat score, (b) frequency bias, (c) probability of detection, and (d) false alarm ratio.
 Results from the 60 GDPS-CTL (blue) and GDPS-SN (red) cases of boreal winter are compared to ground observations over North America used by CaPA.
@@
 For the results presented in this section, GDPS TC forecasts are matched with observations, by excluding from consideration any storm that is not detected by both GDPS and IBTrACS systems.
-Spectral nudging improves TC trajectories by reducing the errors resulting from GDPS-CTL’s tendency of predicting storms that move too slowly and veer too much to the right (with respect to its observed track), especially at lead times beyond day 5 (Figs.
-18a and b).
-Conversely, there is little or no significant impact of spectral nudging on the TC intensity, as measured by the minimum central pressure and maximum sustained wind speed (Figs.
-18c and d).
-Both GDPS-CTL and GDPS-SN have a similar week intensity bias, which is a well-known weakness of the GEM model (see McTaggart-Cowan et al.
+Spectral nudging leads to an overall reduction in position error in the predicted TC trajectories (Fig.
+17a).
+Notably, GDPS-CTL’s tendency of predicting storms that move too slowly is improved with GDPS-SN, especially at lead times beyond day 5 (Fig.
+17b).
+There is also an indication of improvement with respect to GDPS-CTL’s tendency of predicting TCs that veer too much to the right from their observed tracks (Fig.
+17c).
+Conversely, there is little or no significant impact of spectral nudging on the TC intensity, as measured by the minimum central pressure (not shown) and maximum sustained wind speed (Figs.
+17d).
+Both GDPS-CTL and GDPS-SN have a similar weak-intensity bias, which is a well-known weakness of the GEM model (see McTaggart-Cowan et al.
 (2024), for more details).
@@
 Furthermore, these results suggest that the weak-intensity bias for TCs could be reduced in future by focusing on improving the physics-based NWP component of the hybrid system.
-Figure 18: Tropical cyclone tracking performance comparison of the GDPS-CTL (blue) and GDPS-SN (red) against IBTrACS data in three northern hemisphere basins (West Atlantic, East and West Pacific) for 62 cases of boreal summer 2022: (a) along-track and (b) cross-track mean position errors, (c) mean central pressure error and (d) mean maximum wind speed error.
+Figure 17: Tropical cyclone tracking performance comparison of the GDPS-CTL (blue) and GDPS-SN (red) against IBTrACS data in three northern hemisphere basins (West Atlantic, East and West Pacific) for 62 cases of boreal summer 2022: (a) mean position error, (b) mean along-track position error, (c) mean cross-track position error, and (d) mean maximum wind speed error.
 Negative along-track errors imply cyclones move too slowly and positive cross-track errors imply cyclones veer to much to the right with respect to the observed trajectory.
+Red and blue markers denote statistically significant differences in favor of GDPS-SN and GDPS-CTL, respectively.
+No marker at a pressure level implies that the null hypothesis, stating that the statistics of the two samples are the same, cannot be rejected based on the 90th percentile.
+The number of cases for different forecast lead times are indicated in 17c.
 6.6 Computational cost of spectral nudging
 The current implementation of spectral nudging in GEM with the presented optimal configuration leads to an increase of computational cost by approximately 25%.
+Applying nudging to specific humidity increases the computational cost by another 10% approximately.
 Due to the proof-of-concept nature of this study, no optimization of the informatics code has been performed.
-Therefore, the cost could be considerably reduced through future code optimization, especially for the spectral filter.
+Therefore, the cost could be considerably reduced through future code optimization for the spectral filter and implementation of an asynchronous input server.
 Further improvements could be attained by considering time-varying nudging increments, where nudging would only be applied at time-steps close to the inference times of GC.
@@
 Naturally, serious questions about the future role of AI and NWP models in operational forecasting are arising.
-For instance, are AI models capable of adding any reliable and meaningful value for weather services, or will NWP models eventually disappear from the forecasting landscape and be replaced by AI models?
+For instance, are AI models capable of adding any reliable and meaningful value for weather services, or is it possible to improve guidance from operational NWP models by leveraging the AI models?
 This study attempts to address these pressing questions by choosing the GEM and GC models to represent the two forecasting approaches.
@@
 Skill verification against observations and GDPS analyses reveals that GC suffers from excessive fine-scale smoothing, affecting scales as large as 2500 km beyond day 3 of a 10-day forecast.
-However, despite smoothing, the results demonstrate that large-scale features (wavelength$>$2500 km) in GC exhibit better skill compared to GEM, which is evidenced by increased spectral coherence relative to the GDPS analyses and reduced STDE against radiosonde observations.
+However, despite smoothing, large-scale features (wavelength$>$2500 km) in GC exhibit better skill compared to GEM, which is evidenced by increased spectral coherence relative to the GDPS analyses and reduced STDE against radiosonde observations.
 Inspired by the improved large-scale skill of GC, a hybrid NWP-AI system – namely, GDPS-SN – has been developed to produce global forecasts where GEM-predicted large scales in horizontal wind and temperature are spectrally nudged toward inferences from GC.
@@
 For example, RMSE of the 500-hPa geopotential height is reduced by 5-10%, with an overall predictability improvement of 12-18 hours, peaking at day 7 of a 10-day forecast.
-Notably, trajectories of tropical cyclones are predicted with enhanced accuracy without significant changes in intensity.
-Regarding precipitation, statistically significant improvements are obtained in equitable threat score, frequency bias and false alarm ratio, particularly in non-summer regimes.
-Skill enhancements with GDPS-SN are generally found to be more remarkable in winter when large scales apparently play a more critical role in determining weather at the regional level.
-In addition to enhancing the skill and predictability of GEM, the proposed hybrid NWP-AI system is shown to generate the full spectra of fine scales that is missing in GC inferences.
-Furthermore, the new system is capable of producing all the usual forecast variables that operational meteorologists rely on for forecasting weather in general and high-impact events in particular.
+Notable improvements are also obtained in the predicted precipitation and tropical cyclone trajectories.
+Furthermore, predictions from GDPS-SN are expected to be more physically consistent than GC as, at its core, GDPS-SN employs an NWP model based on the first principles.
+In addition to enhancing the skill and predictability of GEM, the proposed hybrid NWP-AI system is shown to generate the full spectra of fine scales that are missing from GC inferences.
 Comparisons against near-surface observations have revealed that – unlike GC, which has a limited ability to capture the extremes in wind speed and temperature – weather extremes predicted by GDPS-SN closely follow the operational GDPS-CTL.
-This implies that the fine scales predicted by GPDS-SN not only possess similar level of energy as GDPS-CTL, they are also comparable in terms of skill.
-Other aspects of near-surface predictions from GDPS-SN, however, show relatively modest improvement (e.g., screen-level scores).
-Lack of vertical resolution in the boundary layer in GC-13 and inconsistencies in surface forcing (such as orography) between ERA5 and GDPS analyses were the reasons why nudging in the boundary layer has not been considered in this study.
-This largely explains the less-than-remarkable near-surface performance of GDPS-SN.
-However, work is currently in progress at ECCC to fine-tune GC to emulate the operational GDPS analyses.
+Operational NWP models are generally required to track hundreds of prognostic variables while generating forecasts with high vertical resolution and temporal frequency.
+These capabilities are essential for meteorologists to forecast weather patterns, particularly those relevant for high-impact events.
+The proposed hybrid system maintains these necessary capabilities, albeit with a reasonable increase in computational cost.
+In this regard, it is important to recognize that similar capability requirements would substantially diminish the computational cost advantage of the AI models.
+Additionally, training such AI models would be prohibitively expensive.
+Work is currently in progress at ECCC to fine-tune GC to emulate the operational GDPS analyses.
 This fine-tuned version is expected to enhance GC’s skill with the 37-level version, particularly in the stratosphere and the boundary layer.
 Any such improvement could be easily leveraged by the hybrid system in the near future.
-Importantly, a fine-tuned GC would allow for nudging in the boundary layer which is expected to result in further enhancement of near-surface skill with GDPS-SN.
-Finally, results from the proposed hybrid NWP-AI system imply that, rather than viewing AI and NWP models as two competing paradigms, it is likely more prudent to consider these forecasting methods as complementary.
+Importantly, a fine-tuned 37-level version of GC would allow for nudging in the boundary layer which is expected to result in further enhancement of near-surface skill with GDPS-SN, which with the current optimal configuration, shows less remarkable improvements over GDPS-CTL.
+In the future, GDPS-SN could be utilized as the nudging reference for other downstream higher-resolution deterministic GEM-based operational NWP systems.
+Additionally, spectral nudging could be investigated for global ensemble prediction using a probabilistic AI model, such as GenCast (Price et al., 2024).
+Finally, results from this study imply that, rather than viewing AI and NWP models as two competing paradigms, it is likely more prudent to consider these forecasting methods as complementary.
 A well-designed fusion of these two approaches can significantly mitigate their individual limitations while allowing for harnessing their respective strengths to provide better meteorological guidance.
@@
 As shown in Fig.
-19, the resulting filtered fields fully retain amplitudes for scales associated with approximately $n<20$ (wavelength$>$2000 km) and fully removes scales corresponding to $n>40$ (wavelength$<$1000 km).
+18, the resulting filtered fields fully retain amplitudes for scales associated with approximately $n<20$ (wavelength$>$2000 km) and fully removes scales corresponding to $n>40$ (wavelength$<$1000 km).
 The physical fields reconstructed from the filtered spectra leads to the desired filtered fields.
@@
 where $\Delta$ denotes the model grid spacing, $\alpha$ is the normalized two-dimensional wavenumber given by $\alpha=\sqrt{\frac{m^{2}}{N_{i}^{2}}+\frac{n^{2}}{N_{j}^{2}}}$ associated with each of the Yin-Yang sub-domains of size $(N_{i}\times N_{j})$, and $\hat{\alpha}=\alpha\lambda_{LS}/(2\Delta)$.
-Figures 20a and 20b illustrate the response of this DCT-based filter for a configuration based on $\lambda_{LS}=1500$ km and $\lambda_{SS}=1000$ km.
+Figures 19a and 19b illustrate the response of this DCT-based filter for the optimal configuration based on $\lambda_{LS}=2750$ km and $\lambda_{SS}=2250$ km.
 The figures demonstrate that the DCT-based filter is capable of targeting the desired scales for the individual Yin-Yang sub-domains.
@@
 The corresponding results are presented in Fig.
-20c.
+19c.
 The figure reveals that, even though DCT is primarily designed for LAMs where curvature effects of spherical surfaces are neglected, its application on global Yin-Yang grid does not result in any major issue as the filter is found to result in an acceptable response.

### v3

@@
 Operational meteorological forecasting has long relied on physics-based numerical weather prediction (NWP) models.
-Recently, this landscape is facing disruption by the advent of data-driven artificial intelligence (AI)-based weather models, which offer tremendous computational performance and competitive forecasting skill.
+Recently, this landscape has faced disruption by the advent of data-driven artificial intelligence (AI)-based weather models, which offer tremendous computational performance and competitive forecasting accuracy.
 However, data-driven models for medium-range forecasting generally suffer from major limitations, including low effective resolution and a narrow range of predicted variables.
-This study illustrates the relative strengths and weaknesses of these competing paradigms using the GEM (Global Environmental Multiscale) and GraphCast models to represent physics-based and AI-based approaches, respectively.
-By analyzing global predictions from these two models against observations and analyses in both physical and spectral spaces, this study demonstrates that GraphCast-predicted large scales outperform GEM, particularly for longer lead times.
-Building on this insight, a hybrid NWP-AI system is proposed, wherein GEM-predicted large-scale state variables are spectrally nudged toward GraphCast predictions, while allowing GEM to freely generate fine-scale details critical for weather extremes.
-Results indicate that this hybrid approach is capable of leveraging the strengths of GraphCast to enhance the prediction skill of the GEM model.
-Importantly, trajectories of tropical cyclones are predicted with enhanced accuracy without significant changes in intensity.
-Furthermore, this new hybrid system ensures that meteorologists have access to a complete set of forecast variables, including those relevant for high-impact weather events.
+This study illustrates the relative strengths and weaknesses of these competing paradigms using the physics-based GEM (Global Environmental Multiscale) and the AI-based GraphCast models.
+Analyses of their respective global predictions in physical and spectral space reveal that GraphCast-predicted large scales outperform GEM, particularly for longer lead times, even though fine scales predicted by GraphCast suffer from excessive smoothing.
+Building on this insight, a hybrid NWP-AI system is proposed, wherein temperature and horizontal wind components predicted by GEM are spectrally nudged toward GraphCast predictions at large scales, while GEM itself freely generates the fine-scale details critical for local predictability and weather extremes.
+This hybrid approach is capable of leveraging the strengths of GraphCast to enhance the prediction skill of the GEM model while generating a full suite of physically consistent forecast fields with a full power spectrum.
+Additionally, trajectories of tropical cyclones are predicted with enhanced accuracy without significant changes in intensity.
+Work is in progress for operationalization of this hybrid system at the Canadian Meteorological Centre.
 1 Introduction
-State-of-the-art numerical weather prediction (NWP) models generally include some form of a dynamical core that solves the nonlinear partial differential equations governing the resolved-scale dynamical and thermodynamical processes within the atmosphere.
-Solutions from the dynamical core are subsequently combined with tendencies from parameterized physics schemes, representing diabatic, frictional and subgrid-scale processes that are not explicitly accounted for by the dynamical equations.
-Although statistical approaches have been explored in the past — particularly for downscaling purposes (Yu et al., 2006; Busuioc et al., 2008; Cheng et al., 2014) — physics-based NWP models have long been the foundational approach for operational meteorological forecasting.
+State-of-the-art physics-based NWP models include some form of a dynamical core that solves the atmospheric governing equations, and are coupled to a suite of parameterization schemes to represent diabatic, frictional, and subgrid-scale processes that are not explicitly accounted for by the dynamical equations.
+Although statistical alternatives have been explored in the past — particularly for downscaling purposes (Yu et al., 2006; Busuioc et al., 2008; Cheng et al., 2014) — physics-based models have long been the foundational approach for operational meteorological forecasting.
 The recent emergence of data-driven models inspired by artificial intelligence (AI) has, however, started to seriously challenge this well established paradigm.
 Keisler (2022) presented an important breakthrough demonstrating considerable potential for weather forecasting with data-driven models.
-This was quickly followed by Pangu-Weather from Huawei (Bi et al., 2022, 2023), GraphCast from Google DeepMind (Lam et al., 2022, 2023) and several other models (e.g., Pathak et al.
+This was quickly followed by Pangu-Weather from Huawei (Bi et al., 2022, 2023), GraphCast from Google DeepMind (Lam et al., 2022, 2023), and several other models (e.g., Pathak et al.
 2022; Chen et al.
@@
 In general, these models rely on some form of deep neural network architecture.
-Like any other application of AI, data-driven weather models necessitate a substantial volume of high quality training data.
+Like any other application of AI, data-driven weather models require a substantial volume of high quality training data.
 As a result, all currently available AI-based deterministic global weather simulators are trained on the ERA5 reanalyses (Hersbach et al., 2020) from the European Centre for Medium-Range Weather Forecasts (ECMWF), which is undoubtedly the most comprehensive resource available.
-The neural network weights within these models are specifically trained to make predictions that closely emulate ERA5.
-One important aspect of these AI-based weather emulators that set them apart from the classical NWP models is their computational performance.
-They can speed up the production of multi-day forecasts by orders of magnitude while using a fraction of the computational resources.
-Furthermore, many of these models are not only comparable to the NWP models with respect to their prediction accuracy, in terms of a number of the performance metrics, they can even surpass the skill of the Integrated Forecasting System (IFS; Lang et al.
-2023) from ECMWF, which is widely accepted as the benchmark within the NWP community (Bi et al., 2023; Lam et al., 2023; Chen et al., 2023a).
-Despite their advantages, AI-based weather emulators also have limitations.
+The neural network weights within these data-driven models are specifically trained to make inferences (forecasts) that closely emulate ERA5.
+Although training of AI-based weather emulators is computationally expensive, their exceptional computational performance during inference accelerates production times by orders of magnitude while using a fraction of the computational resources usually devoted to the physics-based models.
+In addition to their efficiency, standard headline scores suggest that data-driven systems generate predictions that are more accurate even than those of the Integrated Forecasting System (IFS; Lang et al.
+2023) from ECMWF, the model used to generate the ERA reanalyses on which data-driven models were trained.
+Despite these advantages, AI-based weather emulators also have limitations.
 A major weakness of most of the currently available data-driven models is considerable fine-scale smoothing (Bi et al., 2022; Lam et al., 2022).
-Data-driven models based on generative AI, e.g., diffusion-based GenCast model from Google DeepMind (Price et al., 2024), can address the fine-scale smoothing problem to some extent.
-However, as of now, such approaches generally require considerable increase in computational cost, thereby diminishing one of the principal advantages of the AI models.
-Improving the nominal horizontal resolution of the AI model inferences (i.e., forecasts) is also challenging.
-The general construction of these models requires higher-resolution training data to enable these models to make inferences at higher resolutions.
-Since the ERA5 reanalysis is only available at 0.25°$\degree$ resolution, this restriction means that current global AI models cannot be trained for kilometer-scale forecasting.
-Another common limitation of the current generation of AI models is the limited range of predicted variables.
-At present, these models can only be trained to emulate fields that are present in the analyses dataset.
-However, training to predict non-analyzed variables present in the analyses (e.g., precipitation in ERA5) would generally lead to suboptimal inference.
-Furthermore, adding the full suite of physics-related prognostic variables required for operational NWP, such as categorization of cloud and precipitation types, would be cost-prohibitive with current AI model architectures.
-Time resolution of inferences is another constraint for most of the current AI models, which make predictions at 6-hour intervals, far coarser than physics-based NWP models.
-Pangu-Weather (Bi et al., 2022) provides separate models for 1-, 3-, 6-, and 24-hour predictions, but the models with shorter forecasting steps suffer from larger error accumulation.
-In summary, despite their strengths, the aforementioned limitations make it implausible for the AI models to replace the classical NWP approach in operational weather forecasting.
-The ultimate goal of this study is to develop a hybrid NWP-AI system for real-time forecasting applications that combines the strengths of the physics-based NWP approach with those of the AI-based approach while addressing the inherent limitations of the AI models.
-The physics-based model chosen for this study is the GEM (Global Environmental Multiscale) model (Girard et al., 2014), which is employed by Environment and Climate Change Canada (ECCC) for operational weather forecasting.
-The AI model in this study is represented by GraphCast from Google DeepMind.
-Based on internal testing at ECCC, among the different freely available open-source AI models, GraphCast demonstrates the best accuracy with respect to most of the metrics.
-The study aims to determine whether large-scale spectral nudging of GEM predictions toward inferences from a state-of-the-art AI-based weather emulator, such as GraphCast, can improve GEM’s predictability and offer a feasible approach for developing a suitable hybrid NWP-AI system.
-Spectral nudging can be seen as an indirect suboptimal data assimilation method (von Storch et al., 2000).
-Although the concept of spectral nudging was first proposed to control spatial computational modes in a limited-area model (LAM; Waldron et al.
-1996), it is primarily used to improve dynamical downscaling of coarse-resolution atmospheric model simulations over high-resolution limited area domains (Leduc and Laprise, 2009; Husain et al., 2014).
+Furthermore, these models are unable to accurately represent the fundamental dynamical balances in the atmosphere, leading to smoothing that is inconsistent across physically-related variables (Bonavita, 2024).
+It is claimed that smoothing can largely be addressed by employing a diffusion architecture (Price et al., 2024); however, such approaches currently increase computational cost considerably, thereby diminishing one of the principal advantages of the AI paradigm.
+Increasing the nominal horizontal resolution of AI inferences will also be challenging, as it would require higher resolution training data.
+Since ERA5 reanalyses are only available on a 0.25°$\degree$ grid, current global AI models cannot be trained for kilometer-scale forecasting.
+Current data-driven systems are also only capable of predicting a limited number of variables, a subset of those available in the training dataset.
+Training to predict non-analyzed variables (e.g., precipitation in ERA5) generally leads to suboptimal inference (Lam et al., 2022, 2023).
+Predicting the full suite of physics-related prognostic variables required of operational NWP, such as categorization of cloud and precipitation types, would also substantially increase the computational cost.
+Time resolution of inferences is another constraint for most of the current AI models, whose 6-hour prediction intervals are far coarser than physics-based equivalents.
+Pangu-Weather (Bi et al., 2022) provides separate models for 1-, 3-, 6-, and 24-hour inferences, but the models with shorter forecasting steps suffer from larger error accumulation.
+These limitations make it impossible for current AI-based models to completely replace operational physics-based NWP systems.
+The ultimate goal of this study is to develop a hybrid NWP-AI system for real-time global forecasting applications that combines the strengths of each paradigm while overcoming their individual limitations.
+Although the techniques developed here are generally applicable, the implementation presented in this paper is based on Environment and Climate Change Canada’s (ECCC’s) physics-based Global Environmental Multiscale (GEM) model (Girard et al., 2014) and the AI-based GraphCast model from Google DeepMind.
+The hybridization involves large-scale spectral nudging of GEM predictions toward GraphCast inferences.
+Although the concept of spectral nudging was first proposed to control spatial computational modes in a limited-area model (Waldron et al.
+1996), it is primarily used to improve dynamical downscaling over high-resolution sub-domains for regional climate modelling (Leduc and Laprise, 2009; Husain et al., 2014).
+In general, it can be interpreted as an indirect suboptimal data assimilation method (von Storch et al., 2000).
 Spectral nudging has also been explored for global atmospheric hindcasting (Schubert-Frisius et al., 2017).
-Spectral nudging has so far been unfeasible for real-time forecasting applications due to the absence of timely and accurate reference data.
-The emergence of AI models, like GraphCast, has unlocked the potential for applying spectral nudging in real-time operational weather forecasting.
-In this context, application of spectral nudging for enhancing real-time NWP represents an important novelty proposed by this study.
-Although the hybridization concept is illustrated with a specific combination of AI and NWP models, the proposed strategy is expected to be broadly applicable to any other combination, provided that the AI model offers added value over the NWP model.
-To achieve the most optimal fusion of GraphCast and GEM, it is essential to ascertain the relative strengths and limitations of these models through objective evaluation against observations and analyses.
-A detailed comparative analysis of these models is also critical for devising the most optimal spectral nudging configuration, and therefore, constitutes a significant component of this study.
-This article is organized as follows.
-Sections 2 and 3 provide some important background information regarding the GEM and GraphCast models, respectively.
-Details pertaining to the relative performances of these models are documented in section 4.
-Section 5 covers the concept of spectral nudging, its implementation in GEM as well as the optimal nudging configuration used in this study.
-Detailed evaluations of the spectral nudging-based hybrid NWP-AI system are presented in 6.
-Finally, section 7 summarizes the main conclusions of this study and outlines potential directions for future work.
-2 The NWP System
+Spectral nudging has so far been unfeasible for real-time global forecasting applications due to the lack of a timely and accurate reference data source.
+However, the emergence of AI models, with their rapid inference speed and improved large-scale skill, as shown in this study, has unlocked the potential for applying spectral nudging in real-time operational global weather forecasting.
+The development of the hybrid NWP-AI system presented here, begins with background information on the GEM and GraphCast models in section 2.
+The relative performance of these models is assessed in section 3.
+Section 4 introduces spectral nudging in GEM and documents an optimal nudging configuration.
+Detailed evaluations of the hybrid system are presented in section 5, followed by a summary of the study’s main conclusions and recommendations for future work in section 6.
+2 Model descriptions
 2.1 The GEM Model
@@
 The global domain is represented as a pair of overlapping limited-area grids in a Yin-Yang configuration (Qaddouri and Lee, 2011).
-Solutions from the GEM dynamical core are combined with tendencies from parameterized physics schemes (McTaggart-Cowan et al., 2019) that represent the diabatic, frictional and subgrid-scale processes.
-The parameterization schemes are executed in a sequential manner.
-Sequential coupling (Gross et al., 2018), also referred to as split-type coupling (Husain et al., 2019), is employed to combine the parameterized physics tendencies with the resolved dynamics to obtain the complete solution for a model time step.
-2.2 Global Deterministic Prediction System (GDPS)
-The version of the GDPS used in this study is based on its current operational configuration (McTaggart-Cowan et al., 2019; ECCC, 2024).
-It has a horizontal grid spacing of approximately 15 km along with a time-step length of 450 s.
-In the vertical, GDPS employs 84 prognostic levels with the top approximately at 0.1 hPa.
+Solutions from the GEM dynamical core are combined with tendencies from parameterized physics schemes (McTaggart-Cowan et al., 2019) that represent the diabatic, frictional, and subgrid-scale processes.
+The resulting tendencies augment the dry-dynamical solution via sequential/split coupling (Gross et al., 2018; Husain et al., 2019) to produce the complete solution for a model time step.
+The GEM configuration used here is that of the current operational GDPS (Global Deterministic Prediction System; McTaggart-Cowan et al.
+2019; ECCC 2024), with a horizontal grid spacing of approximately 15 km and 450 s time step.
+The GDPS has 84 prognostic vertical levels with the top approximately at 0.1 hPa.
 Although operational GDPS is coupled with the NEMO ocean model (Smith et al., 2018), all the experiments for this study were conducted with atmosphere-only configurations.
 This simplification reduces both complexity and computational cost, and in the past, has been found to provide reliable guidance for the behavior of the full operational system.
-3 The AI-based Prediction System
 2.2 The GraphCast Model
-The GraphCast (GC) model from Google DeepMind is based on AI, and like other global AI weather models, has been trained to emulate ECMWF’s ERA5 reanalyses (Lam et al., 2022, 2023).
-It requires the state of the atmosphere at times $t$ and $t-\Delta t_{GC}$ to produce a forecast valid at $t+\Delta t_{GC}$, where $t$ is the initialization time and $\Delta t_{GC}$=6 hr.
-Although the model can be employed auto-regressively $N$ times to generate predictions valid at $t+N\Delta t_{GC}$, it is trained to minimize the mean squared error (weighted by vertical levels) for up to 12 recursive forecast steps (from 6 hr to 3 days) by considering ERA5 as the ground truth.
-The objective of such a training is to produce forecasts with reasonable skill for up to day 10.
-At its core, GC employs graph neural networks (GNNs) with an “encoder-processor-decoder” configuration (Lam et al., 2022, 2023).
-The input and output states of GC are represented over a 0.25°$\degree$ latitude-longitude global grid.
+The AI-based GraphCast model from Google DeepMind has been trained to emulate ECMWF’s ERA5 reanalyses available with a horizontal grid spacing of 0.25°$\degree$ (Lam et al., 2022, 2023).
+It is trained to minimize the mean squared error (weighted by vertical levels) for up to 12 recursive forecast steps (from 6 hr to 3 days), and is capable of producing forecasts with reasonable accuracy up to day 10.
+The atmospheric state predicted by GraphCast is represented by 6 variables defined on multiple pressure levels (temperature, $u-v$ components of wind, geopotential, specific humidity, and vertical wind speed) and 5 surface variables (2-m temperature, 10-m wind components, mean sea-level pressure, and total precipitation).
+At its core, GraphCast employs graph neural networks (GNNs) with an “encoder-processor-decoder” configuration (Lam et al., 2022, 2023).
+The input and output states of GraphCast are represented over a 0.25°$\degree$ latitude-longitude global grid.
 A major advantage of GNNs is the possibility of having arbitrary range of spatial interactions.
 GraphCast takes advantage of this GNN feature by introducing a multi-mesh architecture based on iteratively-refined icosahedral grids (from level-0 to level-6 refinements) within the processor component.
-The encoder of GC maps data from the input 0.25°$\degree$ latitude-longitude grid to the internal multi-mesh of the processor, whereas the decoder brings back information from the processor multi-mesh to the output grid.
-The internal multi-mesh of GC processor is spatially homogeneous.
-The highest resolution (level-6) mesh has a nominal resolution of 1°$\degree$ with 40,962 nodes, while the nodes of the coarse-resolution meshes are subsets of these nodes.
-This facilitates multi-scale graph edges that connect the highest-resolution mesh nodes to the nodes associated with the coarse-resolution meshes to achieve multi-mesh message passing.
-Such an approach, therefore, leads to multi-scale interactions ranging from synoptic to mesoscales.
-The atmospheric state predicted by GC is represented by 6 variables defined on multiple pressure levels (temperature, $u-v$ components of wind, geopotential, specific humidity, and vertical wind speed) and 5 surface variables (2-m temperature, 10-m wind components, mean sea-level pressure, and total precipitation).
-In addition to the predicted variables, the input system of GC requires time-dependent forcing in the form of total incident solar radiation at the top of the atmosphere along with information regarding the local time of the day and progression of year.
-These time-dependent forcing terms are computed analytically.
-Furthermore, the input system requires some stationary input features, such as land-sea mask, surface geopotential as well as information on the local longitude and latitude of its input grid.
-Pre-trained weights of GC (learned features of the GNNs) are available from Google DeepMind for two different sets of pressure levels.
-The version with higher vertical resolution has 37 pressure levels with the top at 1 hPa, while the coarser-resolution version has only 13 levels with the top at 50 hPa (Lam et al., 2022).
-It is important to note that, in addition to training on ERA5, the 13-level GC was also fine-tuned by DeepMind on initial conditions for ECMWF’s operational IFS system, from 2016 to 2021.
-3.2 GraphCast-based Global Prediction System
-An experimental GC-based global prediction system has been developed at ECCC.
-Both the 37-level and 13-level versions of GC are available.
-This system takes the GDPS operational analyses (Caron and Buehner, 2022) as its input.
-However, the GNN weights have not been subjected to any fine-tuning at ECCC.
-All GC predictions utilized in this study are based on this experimental system.
-Based on the evaluation of GC predictions with the GDPS analyses as input, the 13-level version of GC outperforms the 37-level version for most variables, at most levels and over most regions of the world.
-Although a detailed analysis of this behavior is beyond the scope of this work, it is believed to be related to systematic differences in stratospheric moisture values between the initializing GDPS analyses and ERA5 reanalyses used to train GC.
-The previously mentioned additional fine-tuning of GC-13L by DeepMind could also have contributed to its superior performance.
-This study is therefore, entirely based on the 13-level version of the system and the acronym ‘GC’ is henceforth used to refer to this version.
-4 Relative performances of the GDPS and GC forecasts
-Ten-day forecasts from GC and GDPS are initialized using the GDPS analysis every 36 h, leading to 60 cases for boreal winter 2022 (from 01 December 2021 at $0$ UTC to 27 February 2022 at 1200 UTC) and 62 cases for boreal summer 2022 (from 01 June 2022 at $0$ UTC to 31 August 2022 at 1200 UTC).
+The encoder of GraphCast maps data from the input 0.25°$\degree$ latitude-longitude grid to the internal multi-mesh of the processor, whereas the decoder brings back information from the processor multi-mesh to the output grid.
+This study employs the 13 pressure-level version of GraphCast with pre-trained weights (learned features of the GNNs) that are available from Google DeepMind.
+Although a 37-level version is available, only the 13-level variant has been subjected to additional fine-tuning with ECMWF’s operational analyses (2016–2021), making it more skillful than the 37-level version.
+In this study, GraphCast is initialized with operational GDPS analyses (Caron and Buehner, 2022) without any fine-tuning applied to its weights.
+3 Relative performances of the GDPS and GraphCast forecasts
+Ten-day forecasts from GDPS and GraphCast are initialized every 36 h, leading to 60 cases for boreal winter 2022 (from 01 January 2022 at 1200 UTC to 31 March 2022 at $0$ UTC) and 62 cases for boreal summer 2022 (from 01 June 2022 at $0$ UTC to 31 August 2022 at 1200 UTC).
 The sequence of forecast integrations initialized from these dates using an atmosphere-only operational GDPS configuration is denoted GDPS-CTL.
-Predictions from the two systems, GDPS-CTL and GC, are first assessed for their ability to depict the full range of atmospheric circulations using variance spectra (section 4a).
+Predictions from the two systems (GDPS-CTL and GraphCast) are first assessed for their ability to depict the full range of atmospheric circulations using variance spectra (section 3a).
 This analysis is aimed at ascertaining effective resolutions of the two models, which will help to fairly interpret their relative predictive skills when comparing against radiosonde observations later in section 3b.
 3.1 Verification with global power spectra
-Although GC employs a 0.25°$\degree$ latitude-longitude resolution grid, its predictions are subject to a smoothing effect, which increases with lead time.
-This is generally assumed to be associated with learning to minimize the mean square error (MSE) (e.g., Keisler (2022), Lam et al.
+Although GraphCast employs a 0.25°$\degree$ latitude-longitude resolution grid, its predictions are subject to a smoothing effect that increases with lead time.
+This is believed to be associated with learning to minimize the mean square error (MSE) (e.g., Keisler (2022), Lam et al.
 (2023)).
-The smoothing effect is well indicated by departures of 120-hr GC forecasts’ global 250-hPa kinetic energy and 500-hPa temperature variance spectra from those of GDPS-CTL forecasts and the analyses (Fig.
+The smoothing effect is evidenced by departures of 120-hr GraphCast forecasts’ global 250-hPa kinetic energy and 500-hPa temperature variance spectra from those of GDPS-CTL forecasts, as well as GDPS and ECMWF analyses (Fig.
 1).
 For the fine scales, GraphCast’s spectra are steeper, leading to a notable variance deficiency at wavenumbers as low as 30, corresponding to lower synoptic and sub-synoptic scales ($<$1500 km).
-More importantly, both GDPS-CTL forecasts and analyses much more closely follow the expected $k^{-3}$ spectral slope pertaining to the 3-dimensional quasi-geostrophic flow (Charney, 1971; Skamarock et al., 2014), which indicates that GC slopes are not realistic.
-Hence, the effective resolution of GC, defined as the highest wave number at which the atmospheric variability is still adequately represented, is considerably lower than its nominal resolution, i.e., grid spacing.
+Unlike GraphCast, both GDPS-CTL forecasts and the two analyses closely follow the expected $k^{-3}$ spectral slope for 3D quasi-geostrophic flow (Charney, 1971; Skamarock et al., 2014).
+The fact that the spectral variance of ECMWF analyses – on which GraphCast has been fine-tuned – closely matches that of both GDPS-CTL forecasts and GDPS analyses demonstrates that GraphCast’s fine-scale smoothing does not originate from any lack of fine-scale variance in the training data.
+Overall, the effective resolution, defined as the highest wave number at which the atmospheric variability is still adequately represented, is considerably lower for GraphCast than its nominal resolution, i.e., grid spacing.
 Conversely, the effective resolution of the classical NWP models is typically about 6$\sim$8 times their nominal resolution (Skamarock, 2004).
 With the current operational configuration, the nominal resolution for GDPS-CTL is 0.1375°$\degree$, leading to an approximate effective resolution of 90$\sim$120 km (McTaggart-Cowan et al., 2024).
-Figure 1: Global (a) kinetic energy power spectrum of 250-hPa horizontal wind and (b) variance spectrum of 500-hPa temperature for the GDPS analyses (black) as well as 120-hr lead time of GDPS-CTL (blue) and GC (red), averaged over 60 cases of boreal winter 2022.
+Figure 1: Global (a) kinetic energy power spectrum of 250-hPa horizontal wind and (b) variance spectrum of 500-hPa temperature for the GDPS analyses (black) and ECMWF analyses (dotted black) as well as 120-hr lead time of GDPS-CTL (blue) and GraphCast (red), averaged over 60 cases of boreal winter 2022.
 The dashed line indicates the -3 slope.
-The 5%–95% confidence interval is presented with light shadings (not discernable for the given logarithmic scale).
-It is important to note that the stationary effects of orography, land-sea contrast and large latitudinal climatological variations tend to be much less affected by smoothing in GC than the less predictable transient anomalies.
-It is therefore useful to introduce the transient spectral amplitude ratio, defined as follows:
-$\gamma(n)=\left(\frac{\overline{\sigma_{x^{\prime}x^{\prime}}(n)}}{\overline{\sigma_{y^{\prime}y^{\prime}}(n)}}\right)^{1/2},$ (1)
-where $\sigma_{x^{\prime}x^{\prime}}$ and $\sigma_{y^{\prime}y^{\prime}}$ are real-valued non-negative functions of the total spherical wavenumber $n$, representing global spectral variances of the transient-eddy components associated with the forecast and analysis, respectively.
-The overbar denotes averaging over the cases for a given season.
-The forecast transient eddies are defined as $x^{\prime}=x-\overline{x}$, and similarly for the analysis.
+The 5%–95% confidence interval is presented with light shadings (not discernible for the given logarithmic scale).
+The small- and large-scale cut-offs for the global filter (appendix B) for fair comparison at low resolution (section 3b2) are approximately 1000 km and 2500 km, respectively, as indicated with the shaded area.
+It is important to note that the stationary effects of orography, land-sea contrast and latitudinal climatological variations tend to be much less affected by smoothing in GraphCast than the less-predictable transient anomalies.
+Let $x$ and $y$ denote the forecast and analysis, respectively.
+The transient anomalies – which indicate deviations of individual forecasts (or analyses) from the seasonal mean state – are then defined as $x^{\prime}=x-\overline{x}$ and $y^{\prime}=y-\overline{y}$, where the overbar denotes an average over all cases within the season of interest.
+It is then useful to introduce the transient spectral amplitude ratio, defined as follows:
+$\gamma(n)=\left(\frac{\overline{\sigma_{x^{\prime},x^{\prime}}(n)}}{\overline{\sigma_{y^{\prime},y^{\prime}}(n)}}\right)^{1/2},$ (1)
+where $\sigma_{x^{\prime},x^{\prime}}$ and $\sigma_{y^{\prime},y^{\prime}}$ are real-valued non-negative functions of the total spherical wavenumber $n$, representing global spectral variances of the transient-eddy components associated with the forecast and analysis, respectively.
 The optimal value of $\gamma$ is 1 for the range of scales that are adequately represented in the reference analysis.
 In general, for any given length scale (or wavenumber), a value of $\gamma$ larger than 1 would imply that the forecast has more variability than the analysis, whereas a value smaller than 1 would indicate variance deficiency in the forecast.
-Figure 2 shows the evolution of $\gamma$ with lead time for GC and GDPS-CTL forecasts against the GDPS analysis.
-While $\gamma$ remains close to 1 for all lead times for GDPS-CTL, indicating that its effective resolution does not change during the integration, for GC it rapidly decreases, leading to an effective resolution as low as 1000 km for 24-hr forecasts.
+By replacing the seasonal means with climatology and integrating the individual terms on the right-hand side of Eq.
+1 across the entire spectrum of wavenumbers, one may obtain the ratio of forecast to analysis activity.
+Further information on forecast activity and its spectral decomposition is provided in appendix A.
+Figures 2a,b show the evolution of $\gamma$ with lead time for GraphCast and GDPS-CTL forecasts against the GDPS analysis.
+A model’s effective resolution can be obtained by identifying the smallest wavenumber at which it suffers from a considerable drop in $\gamma$.
+A value of $\gamma<0.9$, implies an amplitude damping of 10%, and is assumed to be the threshold for determining a model’s effective resolution in this study.
+While $\gamma$ remains close to 1 for all lead times for GDPS-CTL, indicating that its effective resolution does not change during the integration, for GraphCast it rapidly decreases, leading to an effective resolution as low as 1000 km for 24-hr forecasts.
 The smoothing effect further reduces the effective resolution to 2500$\sim$2700 km before it saturates around forecast days 3$\sim$5.
-It is worth noting here that GC still resolves scales smaller than these limits, but only partly ($\gamma<1$).
-The impact of this variance-deficiency on optimal usage of GC guidance in a hybrid system will be discussed in more detail in section 5.
-Figure 2: Spectral (a, b) amplitude ratio $\gamma$ and (c, d) coherence $\rho$ for the global (a, c) kinetic energy spectrum of 500-hPa horizontal wind transient-eddy deviations and (b, d) variance spectrum of 500-hPa temperature transient-eddy deviations, for GDPS-CTL (blue) and for GC (red), after 24, 72, 120 and 240 hr of integration (darker to lighter shades of color).
+It is worth noting here that GraphCast still resolves scales smaller than these limits, but only partly ($\gamma<0.9$).
+The impact of this variance-deficiency on optimal usage of GraphCast guidance in a hybrid system will be discussed in more detail in section 4.
+Figure 2: Spectral (a, b) amplitude ratio $\gamma$, and (c, d) coherence $\rho$ for the global (a, c) kinetic energy spectrum of 500-hPa horizontal wind transient-eddy deviations, and (b, d) variance spectrum of 500-hPa temperature transient-eddy deviations, for GDPS-CTL (blue) and for GraphCast (red), after 24, 72, 120, and 240 hr of integration (darker to lighter shades of color).
 Spectra are averaged over 60 cases of boreal winter 2022.
 Light shadings indicate the 5%–95% confidence interval.
-Figure 2 also displays the spectral coherence defined as follows:
-$\rho(n)=\frac{\overline{\sigma_{x^{\prime}y^{\prime}}(n)}}{\left(\overline{\sigma_{x^{\prime}x^{\prime}}(n)}\;\overline{\sigma_{y^{\prime}y^{\prime}}(n)}\right)^{1/2}}\;,$ (2)
-where $\sigma_{x^{\prime}y^{\prime}}$ is a real-valued function of the total spherical wavenumber $n$, representing the transient-eddy covariance spectrum between the GDPS-CTL (or GC forecasts) and the GDPS analyses.
+The dashed horizontal lines in Figs.
+2a,b indicate $\gamma$=0.9.
+The small- and large-scale cut-offs for the global filter (appendix B) for fair comparison at low resolution (section 3b2) are approximately 1000 km and 2500 km, respectively, as indicated with the shaded area.
+Figures 2c,d display the spectral coherence defined as follows:
+$\rho(n)=\frac{\overline{\sigma_{x^{\prime},y^{\prime}}(n)}}{\left(\overline{\sigma_{x^{\prime},x^{\prime}}(n)}\;\overline{\sigma_{y^{\prime},y^{\prime}}(n)}\right)^{1/2}}\;,$ (2)
+where $\sigma_{x^{\prime},y^{\prime}}$ is a real-valued function of the total spherical wavenumber $n$, representing the transient-eddy covariance spectrum between the GDPS-CTL (or GraphCast forecasts) and the GDPS analyses.
 Spectral coherence can be interpreted as a scale-dependent spatial correlation coefficient with an optimal value of 1, which would imply a perfect spatial correlation between forecast and analysis transient-eddy anomalies.
 Conversely, a value of 0 would imply a total de-correlation.
-By integrating the individual terms on the right-hand side of Eq.
-2 across entire spectrum of wavenumbers, one may obtain a quantity which is analogous to the anomaly correlation, but in space.
+By replacing the seasonal means with climatology and integrating the individual terms on the right-hand side of Eq.
+2 across the entire spectrum of wavenumbers, one may obtain a quantity which is analogous to the anomaly correlation.
+Further information in this regard is presented in appendix A.
 Since the same GEM model is used to produce both the GDPS forecasts and backgrounds for the analysis steps, they may have common systematic errors, especially at shorter lead times.
-Even though this penalizes the ERA5-trained GC, Figs.
-2c-d indicate that GC long-lead inferences are superior to the equivalent GDPS predictions over a broad range of synoptic and planetary scales.
+Even though this penalizes the ERA5-trained GraphCast, Figs.
+2c-d indicate that GraphCast long-lead inferences are more skillful compared to the equivalent GDPS predictions over a broad range of synoptic and planetary scales.
 3.2 Verification against radiosondes
-At ECCC, model developers often rely on forecast verification against ECMWF-HRES analysis (Lang et al., 2023) to avoid the problem associated with own-analysis verification (see, e.g., Casati et al.
+At ECCC, model developers often rely on forecast verification against ECMWF analysis (Lang et al., 2023) to avoid the problem associated with own-analysis verification (see, e.g., Casati et al.
 (2008), Privé et al.
 (2021)).
-Conversely, since GC was primarily trained with the ERA5 reanalysis and then fine-tuned on the ECMWF-HRES analysis, this likely introduces some dependency on the ECMWF analysis systems.
-To ensure that the dataset used to verify the forecasts is as independent as possible from both models as well as the GDPS data assimilation system, it was opted to rely mainly on observations from the global radiosondes network.
-Observations of zonal wind ($U$), geopotential height ($Z$), temperature ($T$) and dewpoint depression ($T–T_{d}$) at mandatory pressure levels between 1000 and 100 hPa are therefore used as reference for verification.
-Two flavors of verification against radiosonde observations are performed based on the results shown in the previous section: (1) verification at full resolution, with the GDPS-CTL and GC forecasts at their native resolution, and (2) verification at low resolution –- corresponding to the effective resolution of GC – wherein both the GDPS-CTL and GC forecasts are filtered to remove the variability at scales not adequately resolved by GC.
-The fact that GC has a much lower effective resolution than GDPS-CTL would otherwise provide an unfair advantage to GC by reducing the so-called “double penalty” for misplacing the predicted patterns (Mass et al., 2002).
-The filter applied here is a spectral filter based on spherical harmonics.
-It approximately removes scales smaller than 1000 km (total wavenumber 40) and fully preserves scales larger than 2000 km (wavenumber 20).
-These cutoff wavenumbers well reflect the effective resolution of GC at forecast days 3 to 5 (See Figs.
-2a and b).
-Note also that the same filter configuration is applied to GC and GDPS-CTL (more details on this filter can be found in Appendix A).
-Figure 3: Verification for 120-hr forecasts from GC (red) and GDPS-CTL (blue) at their full (top row: a–d) and low (bottom row: e–h) resolutions against global radiosonde observations, averaged over 60 cases in boreal winter.
-The error standard deviation (STDE; solid lines) and bias (mean forecast minus observation; dashed lines) are shown for (a,e) zonal wind, (b,f) geopotential heights, (c,g) temperature and (d,h) dewpoint depression.
-Red and blue markers denote statistically significant differences in favor of GC and GDPS-CTL, respectively.
+Conversely, GraphCast’s training likely introduces some dependency on the ECMWF analysis system.
+To ensure that the dataset used to verify the forecasts is as independent as possible from both models as well as the GDPS data assimilation system, it was opted to mainly rely on observations from the global radiosondes network for zonal wind ($U$), wind modulus ($|\mathbf{V_{H}}|$), geopotential height ($Z$), temperature ($T$), and dewpoint depression ($T–T_{d}$; where $T_{d}$ denotes dewpoint temperature) at mandatory pressure levels between 1000 and 100 hPa.
+Two flavors of verification against radiosonde observations are performed based on the results shown in section 3a: (1) verification at full resolution, with the GDPS-CTL and GraphCast forecasts at their native resolutions, and (2) verification at a low resolution –- corresponding to the effective resolution of GraphCast (around day 3) – wherein both the GDPS-CTL and GraphCast forecasts are filtered to remove the variability at scales not adequately resolved by GraphCast.
+The latter is required because GraphCast’s smoothing artificially reduces the “double penalty” for misplaced patterns in the nominal resolution evaluation (Mass et al., 2002).
+Low resolution fields are computed using a spectral filter based on spherical harmonics.
+The response function of the filter reflects the effective resolution of GraphCast at forecast days 3 to 5 (Figs.
+2a and b; details in appendix B).
+Figure 3: Verification for 120-hr forecasts from GraphCast (red) and GDPS-CTL (blue) at their full (top row: a–d) and low (bottom row: e–h) resolutions against global radiosonde observations, averaged over 60 cases in boreal winter.
+The error standard deviation (STDE; solid lines) and bias (mean forecast-minus-observation; dashed lines) are shown for (a,e) wind modulus, (b,f) geopotential heights, (c,g) temperature, and (d,h) dewpoint depression.
+Red and blue markers denote statistically significant differences in favor of GraphCast and GDPS-CTL, respectively.
 Significance is computed using the $t$-test for bias and the $F$-test for STDE.
 No marker at a pressure level implies that the null hypothesis, stating that the statistics of the two samples are the same, cannot be rejected based on the 90th percentile.
-A low pass filter (Appendix A) is applied on both GC and GDPS-CTL to obtain the low-resolution fields (e–h).
+A low pass filter (appendix B) is applied on both GraphCast and GDPS-CTL to obtain the low-resolution fields (e–h).
 3.2.1 At full resolution
-Comparing the forecasts at their native resolution against radiosonde observations reveals that for most variables and lead times, and for levels between 1000 and 200 hPa, GC forecasts are much closer to the observations than GDPS-CTL.
-Vertical profiles of error averages (bias) oriented as forecast minus observation and error standard deviations (STDE), calculated over the global domain for 120-hr forecasts for boreal winter, highlight the typical verification differences between GC and GDPS-CTL (Fig.
-3a–d).
-The differences between GC and GDPS-CTL in terms of biases are mixed, except for temperature, where clear improvements can be seen with GC.
-In terms of STDE, however, GC clearly shows large improvements over GDPS-CTL at almost every level, and particularly for zonal wind and dewpoint depression.
-Above 200 hPa, however, GC suffers from increased errors, with biases and STDE being comparable or worse than those in GDPS-CTL.
-To summarize the scores at different lead times, variables and vertical levels a single forecast quality index based on the fractional root-mean-square error (RMSE) change is defined as follows:
-$f_{RMSE}=\left[\frac{RMSE(x_{1})}{RMSE(x_{2})}-1\right]\times 100\%,$ (3)
-with $x_{1}$ and $x_{2}$ here representing GDPS-CTL and GC, respectively.
-Positive values of $f_{RMSE}$ thus imply improved forecasts from GC measured by the RMSE against the radiosonde observations, while negative values represent the opposite.
-The forecast quality index $f_{RMSE}$ was computed every 24 hr for each selected variable and level.
-The results for each variable were then vertically averaged for three layers: low-level (from 1000 to 850 hPa), mid-level (from 700 to 300 hPa) and high-level layer (from 250 hPa to 100 hPa).
-Table 1: Latitude and longitude bounds for the 6 subdomains considered in Figures 5, 10 and 12.
+Comparing the forecasts at their native resolutions against radiosonde observations reveals that GraphCast forecasts are generally closer to the observations than GDPS-CTL in the troposphere.
+Figures 3a–d show the corresponding results at day 5.
+While GraphCast shows clear improvements in temperature biases (dashed lines), it suffers from some noticeable deceleration within the troposphere, leading to deteriorations in wind modulus biases.
+In terms of error standard deviation (STDE, solid lines), however, GraphCast clearly shows large improvements over GDPS-CTL for wind modulus and dewpoint depression, although suffers from some deterioration in the mid-to-upper troposphere geopotential height.
+Above 200 hPa, GraphCast suffers from increased errors, with biases and STDE being comparable or worse than those in GDPS-CTL.
+To summarize the scores at different lead times, variables, and vertical levels, a single forecast quality index based on the change in root-mean-square error (RMSE) is defined as
+$f_{RMSE}=\left[\frac{RMSE(x_{1})}{RMSE(x_{2})}-1\right]\times 100\%$ (3)
+where $x_{1}$ and $x_{2}$ represent GDPS-CTL and GraphCast, respectively.
+Positive values of $f_{RMSE}$ indicate improved forecasts from GraphCast, while negative values imply degradations.
+A $f_{RMSE}$ was computed every 24 hr for each selected variable and level, with the results vertically averaged over three layers: low level (from 1000 to 850 hPa), mid level (from 700 to 300 hPa), and upper level (from 250 hPa to 100 hPa).
+Table 1: Latitude and longitude bounds for the 6 subdomains considered in Figures 5, 11 and 13.
 […]
@@
 […]
-Results for boreal winter (see Fig.
-4a) show that $f_{RMSE}$ is quite constant with lead times for each variable and layer.
-The largest improvements from GC are obtained in the low-level layer for all variables and decrease with altitude, turning into deteriorations in the high-level layer for geopotential heights and temperatures.
-GraphCast geopotential heights are also degraded with respect to GDPS-CTL in the mid-level layer at short lead times.
+Results for boreal winter show that $f_{RMSE}$ is quite constant with lead times for most variables and layers (Fig.
+4a).
+The largest improvements from GraphCast are obtained in the low-level layer for all variables and decrease with altitude, turning into deteriorations in the upper-level layer for geopotential heights and temperatures.
+GraphCast geopotential heights are also degraded with respect to GDPS-CTL in the mid-level layer at short lead times, which is consistent with the error profile presented in Fig.
+3b.
 Results for boreal summer are roughly similar (not shown).
-The pressure weighting approach adopted for the training of GC (see Fig.
+The pressure weighting approach adopted for the training of GraphCast (see Fig.
 6 in Lam et al.
-(2022)) is likely responsible for this vertical variation in performances.
-The differences in the upper-level layer between ECMWF-ERA5 and GDPS analyses used for GC initial conditions (e.g., different bias correction approaches for radiances, different data assimilated, distinct biases in the IFS and GEM model, etc.) could also explain some of the degradations observed in the upper-level layer when GC is initialized with the GDPS analysis.
-However, GC initialized with ECMWF-ERA5 has also been shown to increase the RMSE with respect to ECMWF-HRES above the tropopause when compared to analyses (see, e.g., Fig.
+2022) is likely responsible for this vertical variation in performances.
+Systematic differences in the weakly constrained upper-level layer between ECMWF (used for fine-tuning) and GDPS analyses (used for initializing GraphCast) – e.g., different bias correction approaches for radiances, different data assimilated, distinct biases in the IFS and GEM model, etc.
+– could also explain some of the degradations observed in the upper-level layer.
+However, GraphCast initialized with ECMWF-ERA5 has also been shown to increase stratospheric RMSE with respect to ECMWF analyses (see, e.g., Fig.
 19 in Lam et al.
-(2022)).
-The region-wise variations in relative forecast performances can be examined by computing, as before, the forecast quality index $f_{RMSE}$ averaged over all variables and lead times for each vertical layer but, this time, for the 6 subdomains listed in Table 1 with their respective latitude and longitude limits.
-In low- and mid-level layers, GC forecasts show smaller RMSE in every subdomain in both seasons (Figs.
-5a and c).
-In the high-level layer, GDPS-CTL performs better during each hemisphere’s respective winter.
-However, GC’s poor performance during winter over the northern hemisphere is primarily originating from Europe.
-The largest improvements from GC in the mid- and high-level layers are found over the Tropics, which is consistent with the results of Lam et al.
+2022).
+Regional variations in relative forecast performance can be examined by computing $f_{RMSE}$ on the six subdomains listed in Table 1.
+In low- and mid-level layers, GraphCast forecasts show smaller RMSE in every subdomain in both seasons (Figs.
+5a and b).
+At upper levels, GDPS-CTL performs better during each hemisphere’s respective winter and over Asia during boreal summer.
+However, GraphCast’s poor performance during winter over the Northern Hemisphere is primarily originating from Europe and Asia.
+The largest improvements from GraphCast in the mid- and upper-level layers are found over the Tropics, as reported by Lam et al.
 (2022) (see their Fig.
 19).
+This consistency in GraphCast’s regional performance with respect to its initialization with different analyses (ECMWF and GDPS analyses) is reassuring.
 Figure 4: Changes in the forecast quality index (Eq.
-3) against global radiosonde observations, averaged over 3 layers and 60 cases during boreal winter 2022, corresponding to forecasts at (a) high resolution and (b) low resolution.
-Upward-pointing red triangles indicate a reduction of the RMSE in GC forecasts with respect to GDPS-CTL, whereas downward-pointing blue triangles indicate the opposite.
+3) against global radiosonde observations, averaged over 3 layers and 60 cases during boreal winter 2022, corresponding to forecasts at (a) full resolution and (b) low resolution.
+Upward-pointing red triangles indicate a reduction of RMSE by GraphCast with respect to GDPS-CTL, whereas downward-pointing blue triangles indicate the opposite.
 The size of the triangles varies linearly up to a value of 5.0%.
@@
 Figure 5: Heatmap of the changes in the forecast quality index (Eq.
-3) against global radiosonde observations for various geographical domains and atmospheric layers for (a, b) 60 cases of boreal winter 2022 and (c, d) 62 cases of boreal summer 2022.
-Verifications are presented for forecasts (a, c) at their full resolution and (b, d) after applying a low pass filter (Appendix A) on both GC and GDPS-CTL.
-Red shadings indicate a reduction of the RMSE by GC with respect to GDPS-CTL, while blue shadings indicate the opposite.
+3) against global radiosonde observations for various geographical domains and atmospheric layers averaged over all lead times for (a, c) 60 cases of boreal winter 2022, and (b, d) 62 cases of boreal summer 2022.
+Verifications are presented for forecasts (a, b) at their full resolution, and (c, d) after applying a low pass filter (appendix B) on both GraphCast and GDPS-CTL.
+Red shadings indicate a reduction of the RMSE by GraphCast with respect to GDPS-CTL, while blue shadings indicate the opposite.
 3.2.2 At low resolution
-At low resolution, the forecast improvements in terms of RMSE from GC over GDPS-CTL are significantly reduced for all variables, levels, lead times, seasons and geographical domains (Figs.
-3e–h and 4b).
-This reduction in improvements at low resolution is particularly striking in the mid-level layer over the first 72 hr (compare Figs 4a and b), leading to growing forecast improvements with time from GC that saturate or peak near 120 hr (Fig.
-5b) compared to a relatively steady improvement at all lead times found with the full resolution (Fig.
-4a).
-This suggests that the improvements with GC for the first 72 hr at full resolution largely originate from partial smoothing of the less-accurately predicted fine scales that reduces the double penalty.
-Degradations in GC, seen at full resolution over Northern Hemisphere during boreal winter (mostly coming from Europe) and Southern hemisphere during boreal summer, are enhanced at low resolution (compare Figs 5a,c, and 5b,d).
-There are also notable degradations in the mid- and low-level layers over Asia in summer (Fig.
-5d).
-The average reduction factor for $f_{RMSE}$ from full- to low-resolution verification is roughly between 2 to 3 (compare the left and right columns in Fig.
-5), which confirms that the lower effective resolution of GC provides a significant advantage over the GDPS-CTL.
-Nevertheless, large scales of GC forecasts are still, on average, significantly closer to the radiosonde observations than those from GDPS-CTL in the mid- and low-level layers.
-This suggests that the GC forecasts contain useful large-scale information that could substantially improve guidance if successfully incorporated into a hybrid NWP-AI system.
+At low resolution, GraphCast forecast improvements over GDPS-CTL are reduced for all variables, levels, lead times, seasons, and geographical domains (Figs.
+3e–h, 4b, and 5c,d).
+The change is particularly striking in the mid-level layer over the first 72 hr (compare Figs.
+4a and b), with considerably smaller GraphCast forecast improvements than in the full-resolution case.
+The forecast improvements in low- and full-resolution context, however, become more aligned around 120 hr (Fig.
+4b).
+This stands in contrast with a relatively steady improvement at all lead times found at full resolution (Fig.
+4a), implying that these apparent short-range GraphCast forecast improvements, with respect to GDPS-CTL, originate primarily from partial smoothing of the less-accurately predicted fine scales that reduces the double penalty.
+At low resolution, as expected, both GraphCast and GDPS-CTL show a worsening wind speed bias resulting from the suppression of extremes (Fig.
+3e); however, there is no statistically significant difference between their predictions.
+This suggests that the degradation observed in GraphCast at full resolution, relative to to GDPS-CTL (Fig.
+3a), is primarily due to fine-scale smoothing.
+After filtering, regional deteriorations in GraphCast predictions are accentuated and regional improvements are significantly reduced, with the average reduction factor for $f_{RMSE}$ from full- to low-resolution verification being roughly between 3 to 4% lower than the values obtained at full resolution (c.f.
+the columns in Fig.
+5).
+This further confirms that the lower effective resolution of GraphCast provides a significant advantage over the GDPS-CTL.
+Nevertheless, over most regions, and for both seasonal periods, the large scales in GraphCast forecasts are still, on average, considerably closer to tropospheric radiosonde observations than those from GDPS-CTL, implying that they contain useful large-scale information that could substantially improve guidance if successfully incorporated into a hybrid NWP-AI system.
 4 Spectral nudging in GEM
@@
 For global simulations, such a filter should ideally employ a spherical harmonics-based spectral decomposition.
-However, this would require multiple transformations of model solutions between the model’s Yin-Yang grids and an intermediate global Gaussian grid, resulting in significant increase in the computational cost.
-As the present study is primarily of a proof-of-concept nature, in order to keep the approach simple, the spectral filter for nudging was instead chosen to be based on discrete cosine transform (DCT) (Denis et al., 2002), and the filtering is performed separately on the Yin and Yang grids.
-Additionally, nudging is separately applied to the Yin and Yang grid solutions after every dynamics sub-step.
-It is important to note that, for the range of cut-off scales considered in this study, the response of the DCT-based filter (applied separately to the Yin-Yang grids) is found to be closely equivalent to the response of a spherical-harmonics-based filter applied to a global Gaussian field (see Appendix B).
-Large scales that are retained with spectral nudging are defined through configurable parameters.
-Instead of an abrupt cut-off for the target scales, which may lead to Gibbs phenomena (Sardeshmukh and Hoskins, 1984), a soft cut-off is employed within the DCT-based spectral filter.
-With this method, the spectral filter relies on two different cut-off scale parameters denoted by $\lambda_{LS}$ and $\lambda_{SS}$ such that scales larger than $\lambda_{LS}$ are fully retained while those smaller than $\lambda_{SS}$ are completely eliminated.
-A partial filtering, similar to the one utilized by Husain et al.
-(2014), is applied for the scales in between $\lambda_{LS}$ and $\lambda_{SS}$.
-Further information on the DCT-based filter and its response is provided in Appendix B.
+However, this would require multiple transformations of model solutions between the model’s overlapping limited-area Yin-Yang grids and an intermediate global Gaussian grid, resulting in significant increase in the computational cost.
+As the present study is primarily of a proof-of-concept nature, the spectral filter for nudging was instead chosen to be based on discrete cosine transform (DCT) (Denis et al., 2002), and the filtering is performed separately on the Yin and Yang grids.
+The DCT-based spectral filter employs two cut-off wavelengths for large and small scales, denoted by $\lambda_{LS}$ and $\lambda_{SS}$, respectively.
+Scales larger than $\lambda_{LS}$ are fully retained by the filter, whereas scales smaller than $\lambda_{SS}$ are entirely removed, with a partial filtering of the scales in between (appendix C).
+This partial filtering, often referred to as soft cut-off (Husain et al., 2014), helps to minimize Gibbs oscillations (Sardeshmukh and Hoskins, 1984).
+The values of $\lambda_{LS}$ and $\lambda_{SS}$ are selected to produce a qualitatively acceptable response of the filter when tested over a global Gaussian grid of equivalent resolution (appendix C).
+Overall, the evaluation of $\left[F_{GC}-F_{GEM}\right]_{LS}$ for the purpose of spectral nudging involves computing the spectral coefficients of $(F_{GC}-F_{GEM})$ using a DCT, followed by the application of the spectral filter described in appendix C.
+An inverse DCT is then applied to obtain the large-scale differences in physical space.
 4.3 Optimal configuration
-A series of systematic sensitivity experiments was carried out to identify an optimal spectral nudging configuration for GDPS.
-Results presented in section 6 are based on this optimal configuration which is referred to as GDPS-SN.
-Note that spectral nudging is currently only applied within the forecasting component of GDPS, i.e., GDPS-CTL, GDPS-SN and GC are all initialized with the same analyses.
+A series of systematic sensitivity experiments was carried out to identify an optimal spectral nudging configuration for the hybrid GDPS forecasts, referred to as GDPS-SN.
+Note that spectral nudging is currently only applied within the forecasting component of GDPS, i.e., GDPS-CTL, GDPS-SN, and GraphCast are all initialized with the same analyses.
 Some key aspects of the optimal nudging configuration are discussed below.
@@
 The scales in between are subjected to partial filtering (see Fig.
-19).
+20).
+Over a global Gaussian grid of comparable resolution, the filter response is approximately equivalent to a soft cut-off between 5500-km and 2000-km scales (i.e., $\lambda_{LS}$=5500 km and $\lambda_{SS}$=2000 km), which provides a more relevant context for interpreting the results.
 This choice of nudging length scales corresponds to the effective resolution of GC beyond day 3 (see Figs.
@@
 5.1 Verification with global power spectra
-The power spectra of the GDPS-SN forecast variables is indistinguishable from GDPS-CTL (not shown), which essentially suggests that the new hybrid system can adequately resolve the full range of scales present in the GDPS-CTL, without significant smoothing.
+The power spectra of the GDPS-SN forecast variables are indistinguishable from GDPS-CTL (not shown), which essentially suggests that the new hybrid system can resolve the full range of scales present in the GDPS-CTL without significant smoothing.
 The resulting spectral amplitude ratio of transient-eddy anomalies, $\gamma$ (Eq.
-1), also has values very close to one for all prognostic variables, lead times and spatial scales for both GDPS-CTL and GDPS-SN (Fig.
-7).
-This provides a solid proof-of-concept for the spectral nudging approach.
-The only exception is a 10-percent reduction of $\gamma$ for scales between 2000 and 4000 km at lead times approaching day 10 (wavenumbers 10–20 in Figs.
+1), for GDPS-SN also has values very close to the desired value of 1 for all prognostic variables, lead times, and spatial scales (Figs.
 7a, b).
-This reduction is explained by GC having some variance deficiency at the corresponding range of scales (2000–2750 km) for longer lead times (beyond day 4).
-Sensitivity tests with time-varying cut-off length scales that more strictly follow GC’s time-evolving effective resolution showed that this issue could be avoided, but at the expense of significantly reduced GDPS-SN forecast improvements with respect to the GDPS-CTL.
-This is caused by a weaker constraint being imposed on GDPS-SN towards GC at synoptic scales by such a configuration.
+The only exception is a 10% reduction of $\gamma$ for scales between 2000 km and 4000 km at lead times approaching day 10 (wavenumbers 10–20 in Figs.
+7a, b).
+This reduction is explained by GraphCast having some variance deficiency at the corresponding range of scales (2000–2750 km) for longer lead times (beyond day 4; Figs.
+2a, b).
+Sensitivity tests with time-varying cut-off length scales ($\lambda_{LS}$ and $\lambda_{SS}$) that more strictly follow GraphCast’s time-evolving effective resolution showed that this issue could be avoided, but at the expense of significantly reduced GDPS-SN forecast improvements with respect to the GDPS-CTL.
+This is because larger cut-offs imply a weaker GraphCast-based constraint at synoptic scales in the hybrid system.
 Comparison of the spectral coherence (Eq.
-2) between the forecasts and GDPS analysis shows an improved skill of GDPS-SN over GDPS-CTL (Figs.
+2) between the forecasts and GDPS analysis shows improved skill in GDPS-SN over GDPS-CTL (Figs.
 7c, d).
-These improvements are somewhat weaker than those seen with GC at shorter lead times (up to day 5), but become comparable beyond day 5 (Figs 2c, d).
-The scales relevant for improvements at shorter lead times are smaller than those included in the spectral nudging configuration.
-Nudging the smaller scales would therefore bring the coherence on par with GC, particularly for short lead times, but it would result in variance deficiency, as discussed earlier.
-Figure 7: Same as in Fig. 2, but for GDPS-SN and GDPS-CTL. Additionally, for convenience of comparison, GC has been added in Fig. 7c and d.
-An illustrative example of the value added by the new hybrid system over GC is provided in Fig.
+The improvements are somewhat smaller than those seen with GraphCast at shorter lead times (up to day 3), but become comparable thereafter (Figs 7c, d).
+The scales relevant for short-range improvements are smaller than those targeted by the chosen spectral nudging configuration.
+Inclusion of finer scales for nudging would improve short-range coherence, but it would result in problematic variance deficiency at longer leads.
+Figure 7: Same as in Fig.
+2, but for GDPS-SN and GDPS-CTL.
+Additionally, for convenience of comparison, GraphCast has been added in Figs.
+7c and d.
+The small- and large-scale cut-offs for the DCT-based filter over a global Gaussian grid (appendix C) are approximately 2000 km and 5500 km, respectively, as indicated with the shaded area.
+The impact of spectral nudging in the physical space is shown using maps of transient-eddy anomalies of temperature and specific humidity in Fig.
 8.
-For temperature, which is explicitly nudged in GDPS-SN, the large-scale features of its transient-eddy anomalies tend to very closely resemble GC (Fig.
+Nudging ensures that the large-scale temperature anomalies closely resemble GraphCast (Figs.
 8a–c).
-Even for specific humidity, a variable which is not directly constrained by nudging, the large scales are also found to be more aligned with GC than with GDPS-CTL (Fig.
+Even for specific humidity, a variable which is not directly constrained by nudging, the large scales are found to be more aligned with GraphCast than with GDPS-CTL (Figs.
 8d–f).
-This is however not surprising as it reinforces the findings of previous studies on spectral nudging (Husain et al., 2014).
+This reinforces the findings of previous studies on spectral nudging (Husain et al., 2014).
 Regarding fine scales, Fig.
-8 clearly shows that GDPS-SN has similar level of variability as GDPS-CTL, which is not the case with GC as it clearly suffers from severe smoothing.
-Figure 8: 10-day forecast of 700-hPa transient-eddy anomalies for (a–c) temperature and (d–f) specific humidity, valid 22 February 2022 at 12z for GDPS-CTL (left), GC (middle) and GDPS-SN (right).
+8 shows that GDPS-SN has similar level of variability as GDPS-CTL in contrast to the heavily smoothed GraphCast fields.
+Figure 8: 10-day forecast of 700-hPa transient-eddy anomalies for (a–c) temperature (°$\degree$°C), and (d–f) specific humidity (g kg-1), valid at $0$ UTC, 18 February 2022 for GDPS-CTL (left), GraphCast (middle), and GDPS-SN (right).
+The black rectangles in Figs.
+8 d-f indicate the region that is later used for presenting results in Fig.
+9.
+An illustrative example of the value added by the proposed hybrid system is presented in Fig.
+9, which shows the 5.5-day forecasts by GDPS-CTL, GraphCast, and GDPS-SN, as well as GDPS analysis, valid at $0$ UTC on 18 February 2022.
+GDPS analysis shows the presence of the extratropical winter storm Eunice, which is also predicted by GDPS-SN, whereas both GDPS-CTL and GraphCast fail to predict the storm for this lead time.
+It is important to note that all three models are capable of predicting the storm at the next initialization time ($0$ UTC on 14 February 2022; not shown).
+These results imply that, for a lead time of 5.5 days, although GraphCast may predict large scales that are favorable for the formation of Eunice, it fails to develop the storm, presumably due to excessive smoothing at scales smaller than 2000 km.
+Since GDPS-SN does not suffer from similar smoothing, it can effectively leverage GraphCast’s more accurate large-scale information, resulting in a significant gain in predictability for this specific event, compared to both GraphCast and GDPS-CTL.
+More importantly, these results demonstrate that large-scale nudging of GDPS-SN toward GraphCast can lead to mesoscale features that are substantially different from both GDPS-CTL and GraphCast.
+In other words, the spectral nudging-based hybrid system has the potential to add substantial value beyond what is attainable through any offline post-processing method that combines NWP and AI models.
+Figure 9: Maps of mean sea-level pressure (hPa) and 10-m wind speed (m s-1), valid at $0$ UTC, 18 February 2022 from: (a) GDPS analyses, and 5.5-day forecasts from (b) GDPS-CTL, (c) GraphCast, and (d) GDPS-SN.
+Results are presented over a domain identified with a black rectangle in Fig.
+8d–f.
 5.2 Verification against radiosondes
-Comparison of forecasts at full resolution against the global radiosonde observations clearly indicates that GDPS-SN forecasts are closer to the observations than GDPS-CTL forecasts for basically all variables, lead times and levels, and over both seasons (Fig.
-9).
-The improvements gradually grow in the first 72 hr and are remarkably constant with altitude, which contrasts with the strong vertical variations observed with GC in Section 4.
-This demonstrates that large-scale spectral nudging, despite being only applied in the mid-level layer for winds and temperature, is sufficient to impart similar improvements at both upper- and lower-level layers that are not directly constrained by spectral nudging.
+Comparison of forecasts at full resolution against the global radiosonde observations clearly indicates that GDPS-SN forecasts are closer to the observations than GDPS-CTL forecasts across variables, lead times, and levels for both seasons (Fig.
+10).
+The improvements grow gradually over the first 72 hr and are remarkably constant with altitude, which contrasts with the strong vertical variations observed with GraphCast in Section 3.
+This demonstrates that large-scale spectral nudging, despite being only applied to free-tropospheric winds and temperature, is sufficient to impart considerable improvements to the unconstrained upper- and lower-level layers.
 Compared to boreal winter (Fig.
-9a), the improvements from GDPS-SN are reduced by a factor of about 2 during boreal summer (Fig.
-9b), which is not surprising given that the large-scale improvements with GC over GDPS-CTL are smaller by about the same factor during boreal summer (not shown).
-In terms of region-wise performances, the GDPS-SN has smaller RMSE than GDPS-CTL in every subdomain except over Asia in the mid- and upper-levels during boreal summer (Figs.
-10c and d), which is consistent with the degradations from GC over the GDPS-CTL at the large scales in this region (Fig.
+10a), the improvements from GDPS-SN are generally reduced during boreal summer (Fig.
+10b), which is consistent with the relative performance of the component systems (not shown).
+Regionally, GDPS-SN has smaller RMSE than GDPS-CTL in every subdomain except over Asia in the mid- and upper-levels during boreal summer (Figs.
+11b and d), which is consistent with the degradations noted for GraphCast large scales in this region (Fig.
 5d).
-Unlike GC, the improvements from GDPS-SN over GDPS-CTL are virtually identical when only the large-scale components of the forecast are compared to the radiosonde observations (compare the left and right columns in Fig.
-10 and, again, in Fig.
-5).
-This does not necessarily imply that the improvements only originate from the large scales.
-It rather implies that the relative RMSE reductions with GDPS-SN are of similar level for both the large and small scales.
-In other words, as the small scales are generally conditioned by the large scales, these results suggest that the fine scales predicted by GDPS-SN benefit from fractional improvements that are on par with its large-scale skill enhancement.
+Unlike GraphCast, the forecast improvements of GDPS-SN over GDPS-CTL, when measured against radiosonde observations, are comparable at both full resolution (left column of Fig.
+11) and for large scales only (right column of Fig.
+11, compare Fig.
+5), showing only a small degradation under the low-pass filter.
+This implies that the fine scales predicted by GDPS-SN benefit from fractional improvements at large scales, which is supported by the improvements in fine-scale spectral coherence with GDPS-SN over GDPS-CTL, particularly beyond day 3 (Fig.
+7c-d).
 Figure 10: Same as in Fig.
-4, except that upward-pointing red triangles indicate a reduction of the RMSE in GDPS-SN with respect to GDPS-CTL, whereas downward-pointing blue triangles indicate the opposite.
+4, except that upward-pointing red triangles indicate a reduction of RMSE by GDPS-SN with respect to GDPS-CTL, whereas downward-pointing blue triangles indicate the opposite.
 Only verification using forecasts at full resolution is shown.
-Figure 10: Same as in Fig. 5, except that red shadings indicate a reduction of the RMSE in GDPS-SN with respect to GDPS-CTL, while blue shadings indicate the opposite.
+Figure 11: Same as in Fig. 5, except that red shadings indicate a reduction of RMSE by GDPS-SN with respect to GDPS-CTL, while blue shadings indicate the opposite.
 Figure 12: Same as in Fig.
-4, except that upward-pointing red triangles indicate a reduction of the RMSE in GDPS-SN with respect to GC, whereas downward-pointing blue triangles indicate the opposite.
-Only verification using forecasts at low resolution is shown but for both seasons.
-It is also interesting to compare GDPS-SN to GC.
-For the sake of fair comparison, this is only done at low resolution.
-Predictions from GDPS-SN are closer to the radiosonde observations than GC in the high- and mid-level layers, except for dewpoint depression.
-Conversely, for the low-level layer, GC is closer to the observations, except for geopotential height (Fig.
+4, except that upward-pointing red triangles indicate a reduction of RMSE by GDPS-SN with respect to GraphCast, whereas downward-pointing blue triangles indicate the opposite.
+Only verification using forecasts at low resolution is shown, but for both seasons.
+Forecasts from GDPS-SN must also be compared to those from GraphCast to ensure that the skill of the AI-based system is not lost during hybridization.
+For the sake of fair comparison, this is only done at low resolution (section 3b2).
+Predictions from GDPS-SN are better than GraphCast in the upper- and mid-level layers, except for dewpoint depression (Fig.
+12).
+Conversely, GraphCast forecasts are closer to observations at lower levels, except for geopotential height.
+The dewpoint depression degradations suggest that further gains could be achieved by spectrally nudging the humidity field.
+However, this will only be viable once a fine-tuned version of GraphCast becomes available because of systematic differences between the moisture state estimates of ERA5 and the initializing GDPS analyses.
+Fine-tuning will bring the mean moisture state of GraphCast inferences closer to GDPS analyses and is expected to reduce errors associated with any adverse response from the GEM physics suite.
+The lower-level results also suggest that extending spectral nudging to the boundary layer could be beneficial.
+However, this would require an AI model trained on GEM’s terrain-following vertical coordinate with more vertical levels in the boundary layer, fine-tuned to emulate the GDPS analyses for improved consistency with the GEM model’s lower-boundary forcings.
+The small but growing GDPS-SN improvements over GraphCast in the mid-level layer for spectrally-nudged temperature and zonal wind are somewhat surprising (Fig.
+12).
+The nudging relaxation time of $\tau=12$ hr implies that the large scales in GDPS-SN are not too tightly constrained, which is visible in Fig.
+8.
+This may allow for some large-scale improvements through upscale propagation of information from an improved fine-scale representation.
+However, this choice of $\tau$ also leads to a small averaging effect caused by the blending of two forecasts.
+To a lesser extent, smoothing attributable to this effect may also have contributed to the noted improvements.
+Finally, the regional comparison of GDPS-SN and GraphCast forecasts at low resolution (Fig.
+13) shows that the prediction accuracy of GDPS-SN is typically slightly better than or comparable to GraphCast, with the tropical boundary layer being one of the exceptions.
+Fig.
+5 reveals that, compared to GraphCast, GDPS-CTL large scales have substantially reduced accuracy in the tropics, which is slightly improved with GDPS-SN (Fig.
 11).
-The degradations with respect to dewpoint depression, also observed in the mid-level layer where spectral nudging is active, suggest that further gains could likely be achieved by nudging the humidity field in GEM towards GC.
-Nudging of specific humidity, however, would only be viable once a fine-tuned version of GC that emulates GDPS analyses instead of ERA5 (and ECMWF-HRES) is available.
-This is expected to bring the mean moisture state of GC inferences closer to GDPS analyses and thereby reduce any error associated with potential adverse response from GEM physics.
-The results also suggest that extending spectral nudging to the boundary layer could be beneficial.
-This, however, would necessitate an AI model with more vertical levels in the boundary layer (as mentioned earlier) and fine-tuned to emulate the GDPS analyses rather than ECMWF-ERA5, for improved consistency with the GEM model’s lower-boundary forcings.
-The small but statistically significant GDPS-SN improvements over GC in the mid-level layer for temperature and zonal winds, two variables that are nudged toward GC, that grow with the integration time, are somewhat surprising (Fig.
-11).
-The use of a nudging relaxation time $\tau=12$ hr implies that the large-scales in GDPS-SN are not too tightly constrained, which is visible in Fig.
-8.
-This likely allows for some large-scale improvements through upscale propagation of information from an improved representation of the fine scales.
-However, this choice of $\tau$ also leads to a small averaging effect caused by the blending of two forecasts, as previously discussed in Section 5.
-To a lesser extent, this may also have contributed to these improvements.
-Finally, the overall region-wise comparison of GDPS-SN and GC forecasts at low resolution (Fig.
-12) shows that the skill of GDPS-SN is either slightly better than or comparable to GC, with the tropical boundary layer being an exception.
-Fig.
-5 reveals that, compared to GC, GDPS-CTL large scales have substantially reduced skill in the tropics, which is slightly improved with GDPS-SN (see Fig.
-10).
-To leverage the full potential of GC in the tropics it appears to be necessary to extend nudging in the boundary layer, which is not viable at this stage for reasons explained earlier in this section.
+To leverage the full potential of GraphCast in the tropics it will likely be necessary to extend nudging in the boundary layer, once fine-tuning to GDPS analyses is complete.
 Figure 13: Same as in Fig.
-5, except that red shadings indicate a reduction of the RMSE in GDPS-SN with respect to GC, whereas blue shadings indicate the opposite.
-Only verification using forecasts at low resolution is shown.
+5, except that red shadings indicate a reduction of RMSE by GDPS-SN with respect to GraphCast, whereas blue shadings indicate the opposite.
+Only verification using forecasts at low resolution is shown, but for both seasons.
 5.3 Verification against ECMWF analyses
-The comparison of GDPS-SN and GDPS-CTL forecasts against ECMWF-HRES analyses leads to very similar results as the comparison against radiosonde observations, with general improvements for all variables, lead times and pressure levels (not shown).
-Figure 13 provides an example of improvements in anomaly correlation coefficient (ACC).
-For 500-hPa geopotential heights, the ACC improvements over the Northern Hemisphere region at forecast day 7 are equivalent to an 13-hr increase in predictability during boreal winter and to an 8-hr increase in summer.
-The corresponding ACC improvements over the Southern Hemisphere region are about 3 hours during austral winter and 22 hours during austral summer.
-There are also notable improvements in zonal wind ACC in the tropics that are equivalent to 30 hours or more (e.g., see Fig.
-13b,e).
-The strength of GC in improving the ACC around day 5 (and beyond) is thus well leveraged by the hybrid system.
-Note that GC has not been included in this figure as its ACC is strongly influenced by the excessive fine-scale smoothing.
-To fairly compare GC against the two versions of GDPS, it is necessary to do that at low resolution (as in section 4b).
-However, based on spectral coherence presented in Fig.
-7c–d, low-resolution ACC of GC is expected to be comparable to GDPS-SN.
-Figure 13: Anomaly correlation coefficient against the ECMWF-HRES analyses with GDPS-CTL (blue) and GDPS-SN (red) for (a,d) northern hemisphere 500-hPa geopotential height, (b,e) tropical 850-hPa zonal wind, and (c,f) southern hemisphere 500-hPa geopotential height.
+The comparison of GDPS-SN and GDPS-CTL forecasts against ECMWF analyses – in terms of $f_{RMSE}$ – leads to conclusions that are similar to those of the radiosonde assessment: general improvements for all variables, lead times, and pressure levels (not shown).
+Figure 14 provides an example of changes in anomaly correlation coefficient (ACC).
+For 500-hPa geopotential heights, the ACC improvements with GDPS-SN (relative to GDPS-CTL) in day-7 forecasts over the Northern Hemisphere region are equivalent to a 11-hr increase in predictability during boreal winter and an 8-hr increase in summer (Figs.
+14a,d).
+The corresponding ACC improvements over the Southern Hemisphere region are about 3 hours during boreal summer and 13 hours during boreal winter (Figs.
+14c,f).
+There are also notable improvements in zonal wind ACC in the Tropics that are equivalent to 26 hours or more (Figs.
+14b,e).
+The strength of GraphCast in improving the ACC around day 5 (and beyond) is thus well leveraged by the hybrid system.
+GraphCast has not been included in this figure because its ACC is strongly influenced by the excessive fine-scale smoothing.
+To fairly compare GraphCast against the two versions of GDPS, it is necessary to do so at low resolution (section 3b); however, based on spectral coherence presented in Figs.
+7c–d, low-resolution ACC of GraphCast is expected to be comparable to GDPS-SN.
+Figure 14: Anomaly correlation coefficient (ACC) against the ECMWF analyses with GDPS-CTL (blue) and GDPS-SN (red) for (a,d) northern hemisphere 500-hPa geopotential height, (b,e) tropical 850-hPa zonal wind, and (c,f) southern hemisphere 500-hPa geopotential height.
 Results are presented for 60 cases of boreal winter 2022 (left column: a–c) and 62 cases of boreal summer 2022 (right column: d–f).
+ACC is computed using climatology of ERA-Interim reanalyses.
 Red and blue markers denote statistically significant differences in favor of GDPS-SN and GDPS-CTL, respectively.
-No marker at a pressure level implies that the null hypothesis, stating that the statistics of the two samples are the same, cannot be rejected based on the 95th percentile.
+No marker for a forecast lead time implies that the null hypothesis, stating that the statistics of the two samples are the same, cannot be rejected based on the 95th percentile.
 Grey arrows with printed numbers depict gains in predictability in terms of forecast hours.
-6.4 Verification against surface observations
-For verification at the surface level, combined SYNOP, METAR and Surface Weather and marine OBservations (SWOB; available only over Canada and distributed by ECCC) data were used for surface pressure, screen-level temperature, dewpoint depression and anemometer-level wind speed.
+5.4 Evaluation against surface observations
+For near-surface evaluation, combined SYNOP, METAR, and Surface Weather and marine OBservations (SWOB; available only over Canada and distributed by ECCC) data were used for surface pressure, screen-level temperature and dewpoint temperature, and anemometer-level wind speed.
 For precipitation, ground observations of 24-hr accumulation, subject to collection and quality control by the CAnadian Precipitation Analysis system (CaPA; Lespinas et al.
 2015), are used.
-Any data pertaining to stations with an altitude difference larger than 100 m with respect to GDPS orography, along with any observed wind speed below 1.5 $\text{m s}^{-1}$, are excluded.
+Any data originating from stations with an altitude difference larger than 100 m with respect to GDPS orography is excluded, whereas any observed or forecast wind speed below 1.5 $\text{m s}^{-1}$ is set to 0.
 Figure 15: Heatmap of the changes in the forecast quality index (Eq.
-3) against combined SYNOP, SWOB and METAR observations for surface pressure ($P_{s}$), screen-level temperature ($T$), dewpoint depression ($T-T_{d}$) and anemometer-level wind speed ($W$) in various geographical domains for (a) 60 cases of boreal winter 2022 and (b) 62 cases of boreal summer 2022.
+3) against combined SYNOP, SWOB, and METAR observations for surface pressure ($P_{s}$), screen-level temperature ($T$), dewpoint temperature ($T_{d}$), and anemometer-level wind speed ($W$) in various geographical domains for (a) 60 cases of boreal winter 2022 and (b) 62 cases of boreal summer 2022.
+Results are computed at full resolution.
 Red shadings indicate a reduction of the RMSE in GDPS-SN with respect to GDPS-CTL, while blue shadings indicate the opposite.
-Figure 15: Q-Q plots of day-5 GC (black), GDPS-CTL (blue) and GDPS-SN (red) forecasts against combined SYNOP, SWOB and METAR for: (a, b) anemometer-level wind speeds and (c, d) screen-level temperatures over North America for (a, c) 30 cases of boreal winter valid at (a) 00 and (c) 12 UTC, and (b, d) 31 cases of boreal summer valid 00 UTC.
+Figure 16: Q-Q plots of day-5 GraphCast (black), GDPS-CTL (blue), and GDPS-SN (red) forecasts against combined SYNOP, SWOB, and METAR for: (a, b) anemometer-level wind speeds, and (c, d) screen-level temperatures over North America for (a, c) 30 cases of boreal winter valid at (a) 00 and (c) 12 UTC, and (b, d) 31 cases of boreal summer valid 00 UTC.
 Local standard time over North America is mostly between UTC-5 and UTC-8 h.
+Results correspond to full resolution of the different models.
 Diamond symbols denote selected percentiles with their values printed.
 Light shadings indicate the 5%–95% confidence interval for the inverse cumulative distribution functions, based on the Kolmogorov-Smirnov statistic.
-Comparison of GDPS-SN and GDPS-CTL using the forecast quality index (Eq.
-3) is summarized in Fig.
-14.
-Guidance by GDPS-SN is generally improved for all variables and regions.
-The improvements are, however, notably smaller than in the upper air, which is particularly the case for the tropics, where achieving considerable improvements appears to be difficult without introducing spectral nudging in the boundary layer.
-The same could be concluded for the screen-level thermodynamic variables and wind.
-On the other hand, the spectral nudging in the mid-level layer has the largest positive impact on the surface pressure via redistribution of upper-air mass.
-The improvements in boreal summer are notably smaller than in boreal winter, consistent with the upper-air verification.
-The ability of the GPDS-SN to represent high-impact weather, such as high wind speeds and extreme temperatures, is not significantly different from the GDPS-CTL but, at the same time, it is substantially better than GC, which provides a strong example of the value added by the hybrid system over GC (Fig.
+Guidance from GDPS-SN is generally improved over GDPS-CTL for all variables and regions (Fig.
 15).
-For example, the distribution of late-afternoon wind speeds over North America provided by GC shows a considerable shift toward lower values, which increases at high percentiles of the distribution, resulting in a -3 $\text{m s}^{-1}$ bias around ${}^{\text{th}}$ percentile (Figs.
+The improvements are, however, smaller than in the upper air (Fig.
+11).
+Achieving considerable improvements near the surface appears to be difficult without introducing spectral nudging in the boundary layer.
+However, spectral nudging in the mid-level layer does have a considerable positive impact on the surface pressure via redistribution of upper-air mass.
+The improvements during boreal summer are reduced, consistent with the upper-air verification discussed above.
+The ability of the GDPS-SN to represent high-impact weather (high wind speeds and extreme temperatures) is not significantly different from the GDPS-CTL (Fig.
+16).
+Both are substantially better than GraphCast, which is unable to predict the tails of the climatological distributions.
+For example, the distribution of late-afternoon wind speeds over North America provided by GraphCast shows a considerable shift toward lower values, which increases at high percentiles of the distribution, resulting in a -3 $\text{m s}^{-1}$ bias near ${}^{\text{th}}$ percentile (Figs.
 16 a, b).
-GDPS-CTL also has a negative bias but affecting only high percentiles and, more importantly, GDPS-SN closely follows GDPS-CTL, with little or no sign of deceleration at extreme wind speeds for both seasons.
-GraphCast also suffers a 2-4 K cold bias at mid-to-high temperature percentiles over North America in summer, which has no impact on the GDPS-SN; its distribution aligns very well with the observations.
-GraphCast’s poor performance with respect to the extremes can be partly explained by its lack of well-resolved fine scales that are generally critical for representing the extremes.
-Furthermore, inconsistencies in the surface forcing between ERA5 and GDPS analyses may also be a contributing factor for the observed underestimation of temperature extremes by GC.
-As the spectral nudging configuration presented in this study only targets synoptic scales and the boundary layer is allowed to evolve freely, GDPS-SN is not adversely affected by GC’s limitations regarding the extremes.
-Spectral nudging also significantly improves boreal winter precipitation forecast guidance over North America for a large range of thresholds from 1 to 50 mm day-1, mostly by reducing the false alarm ratio, with no reduction of probability of detection, which results in a reduced moist frequency bias and an improved equitable threat score (Fig.
-16).
-These improvements seem to be mostly generated via an improved prediction of synoptic-scale flow.
-There is, however, not much of an impact of spectral nudging on summer precipitation over North America (not shown).
+Although GraphCast does not show any discernible bias for low temperature percentiles over North America in winter (Fig.
+16 c), it suffers from a 2$\sim$4 °$\degree$°C cold bias at mid-to-high temperature percentiles in summer (Fig.
+16 d), a problem that does not affect GDPS-SN.
+At high temperature percentiles, GraphCast’s poor performance with respect to the extremes can likely be explained by its lack of well-resolved fine scales (Iles et al., 2020) combined with inconsistencies in the surface forcing between ERA5 and GDPS analyses.
+As the spectral nudging configuration presented in this study only targets synoptic scales, the GEM component of the hybrid system is able to fill in small scales associated with local forcings.
+GDPS-SN is, therefore, not adversely affected by GraphCast’s limitations regarding the extremes.
+Spectral nudging also significantly improves boreal winter precipitation guidance over North America primarily through a reduction in false alarm (Fig.
+17).
+These improvements are likely the direct result of an improved representation of the synoptic-scale flow in strongly forced winter conditions.
+The impact of spectral nudging on summer precipitation over North America is, however, negligible (not shown).
 During summer, when weak synoptic forcings are prevalent, nudging only large scales is not expected to considerably improve precipitation forecasts, although extending nudging to the boundary layer may lead to some improvements by mitigating model biases.
-Figure 16: Quantitative precipitation forecast verification for day 5 assessed by 24-hr precipitation threshold (accumulations between 108 and 132 hr of integration) with (a) equitable threat score, (b) frequency bias, (c) probability of detection, and (d) false alarm ratio.
+Figure 17: Quantitative precipitation forecast verification for day 5 assessed by 24-hr precipitation threshold (accumulations between 108 and 132 hr of integration) with (a) equitable threat score (ETS), (b) frequency bias index (FBI), (c) probability of detection (POD), and (d) false alarm ratio (FAR).
 Results from the 60 GDPS-CTL (blue) and GDPS-SN (red) cases of boreal winter are compared to ground observations over North America used by CaPA.
 Red and blue line markers denote statistically significant differences in favor of GDPS-SN and GDPS-CTL, respectively.
-Significance is computed by bootstrapping 3-day data blocks consisting of forecast-observation pairs from all stations of a given region.
+Significance is computed by bootstrapping 3-day data blocks consisting of forecast-observation pairs from all stations.
 No marker at a threshold level implies that the null hypothesis, stating that the statistics of the two samples are the same, cannot be rejected based on the 90th percentile.
 5.5 Tropical cyclone evaluation
-Evaluation of the impact of spectral nudging on the tropical cyclone (TC) forecast guidance is performed using the data from the International Best Track Archive for Climate Stewardship (IBTrACS; Knapp et al.
-2010).
-For the results presented in this section, GDPS TC forecasts are matched with observations, by excluding from consideration any storm that is not detected by both GDPS and IBTrACS systems.
-Spectral nudging leads to an overall reduction in position error in the predicted TC trajectories (Fig.
-17a).
-Notably, GDPS-CTL’s tendency of predicting storms that move too slowly is improved with GDPS-SN, especially at lead times beyond day 5 (Fig.
-17b).
-There is also an indication of improvement with respect to GDPS-CTL’s tendency of predicting TCs that veer too much to the right from their observed tracks (Fig.
-17c).
-Conversely, there is little or no significant impact of spectral nudging on the TC intensity, as measured by the minimum central pressure (not shown) and maximum sustained wind speed (Figs.
-17d).
-Both GDPS-CTL and GDPS-SN have a similar weak-intensity bias, which is a well-known weakness of the GEM model (see McTaggart-Cowan et al.
-(2024), for more details).
-Severe TC weak-intensity biases are typical of state-of-the-art deterministic AI-based forecast models (e.g., Bouallègue et al.
-(2024)) due to the smoothing effect caused by learning to minimize the MSE.
-However, AI-based models, including GC, are generally capable of predicting the TC trajectories with enhanced accuracy (Lam et al., 2022).
-Therefore, the improvements in the predicted trajectories by GDPS-SN show that the proposed hybrid system is capable of leveraging the well-documented strength of the AI-based models, without sacrificing TC intensity.
-Furthermore, these results suggest that the weak-intensity bias for TCs could be reduced in future by focusing on improving the physics-based NWP component of the hybrid system.
-Figure 17: Tropical cyclone tracking performance comparison of the GDPS-CTL (blue) and GDPS-SN (red) against IBTrACS data in three northern hemisphere basins (West Atlantic, East and West Pacific) for 62 cases of boreal summer 2022: (a) mean position error, (b) mean along-track position error, (c) mean cross-track position error, and (d) mean maximum wind speed error.
-Negative along-track errors imply cyclones move too slowly and positive cross-track errors imply cyclones veer to much to the right with respect to the observed trajectory.
+Evaluation of the impact of spectral nudging on the tropical cyclone (TC) guidance is performed using the data from the International Best Track Archive for Climate Stewardship (IBTrACS; Knapp et al.
+2010) following the methodology described by McTaggart-Cowan et al.
+(2024).
+AI-based models, including GraphCast, are generally capable of predicting the TC trajectories with enhanced accuracy (Lam et al., 2022).
+Results presented in Fig.
+18a show that spectral nudging allows GDPS-SN to leverage GraphCast’s enhanced TC steering accuracy and leads to an overall reduction in position error in the predicted TC trajectories.
+Notably, GDPS-CTL’s tendency of predicting storms that move too slowly – indicated by negative along-track error – is improved with GDPS-SN, especially at lead times beyond day 5 (Fig.
+18b).
+There is also an indication of improvement with respect to GDPS-CTL’s tendency of predicting TCs that veer too much to the right from their observed tracks, which is indicated by positive cross-track error (Fig.
+18c).
+Conversely, there is little or no significant impact of spectral nudging on TC intensity, as measured by the maximum sustained wind speed (Figs.
+18d).
+Although both GDPS-CTL and GDPS-SN have a similar weak-intensity bias, which is a well-known weakness of the GEM model (see McTaggart-Cowan et al.
+(2024), for more details), both versions of GDPS nevertheless predict higher intensity TCs compared to GraphCast.
+Severe TC weak-intensity biases are typical of state-of-the-art deterministic AI-based forecast models due to the double-penalty effect caused by learning to minimize the MSE (Bouallègue et al., 2024).
+Combined with the small scales of the TC vortex that make direct nudging of the pertinent scales infeasible, this means that the potential for improvements in intensity prediction lies primarily in the physics-based NWP component of the hybrid system.
+Figure 18: Tropical cyclone tracking performance comparison of the GDPS-CTL (blue) and GDPS-SN (red) against IBTrACS data in three northern hemisphere basins (West Atlantic, East, and West Pacific) for 62 cases of boreal summer 2022: (a) mean absolute position error, (b) mean along-track position error, (c) mean cross-track position error, and (d) mean maximum wind speed error.
+Performance of GraphCast (dashed black) is presented for reference.
+Negative along-track errors imply cyclones move too slowly and positive cross-track errors imply cyclones veer too much to the right with respect to the observed trajectory.
 Red and blue markers denote statistically significant differences in favor of GDPS-SN and GDPS-CTL, respectively.
-No marker at a pressure level implies that the null hypothesis, stating that the statistics of the two samples are the same, cannot be rejected based on the 90th percentile.
-The number of cases for different forecast lead times are indicated in 17c.
+No marker at a lead time implies that the null hypothesis, stating that the statistics of the two samples are the same, cannot be rejected based on the 90th percentile.
+The number of cases for different forecast lead times are indicated in 18c.
 5.6 Computational cost of spectral nudging
 The current implementation of spectral nudging in GEM with the presented optimal configuration leads to an increase of computational cost by approximately 25%.
-Applying nudging to specific humidity increases the computational cost by another 10% approximately.
+This is in addition to the relatively insignificant cost of generating the GraphCast inferences.
+The nudging overhead scales linearly with the number of fields and levels to which the technique is applied.
 Due to the proof-of-concept nature of this study, no optimization of the informatics code has been performed.
-Therefore, the cost could be considerably reduced through future code optimization for the spectral filter and implementation of an asynchronous input server.
-Further improvements could be attained by considering time-varying nudging increments, where nudging would only be applied at time-steps close to the inference times of GC.
-These considerations however lie beyond the scope of this study.
+In the future, this cost could be considerably reduced through code optimization for the spectral filter and implementation of an asynchronous input server.
+Further improvements could be attained by implementing time-varying nudging, where $\tau$ is allowed to change over time following a cosine-bell profile as proposed by Husain et al.
+(2014).
+With this approach, the application of spectral nudging could be restricted to model time steps close to the inference times of GraphCast, potentially leading to significant computational cost savings.
+However, these considerations lie beyond the scope of this study.
 6 Summary and Future Work
-The recent emergence of AI models has begun to disrupt the operational weather forecasting paradigm which has long been dominated by the physics-based NWP models.
-Naturally, serious questions about the future role of AI and NWP models in operational forecasting are arising.
-For instance, are AI models capable of adding any reliable and meaningful value for weather services, or is it possible to improve guidance from operational NWP models by leveraging the AI models?
-This study attempts to address these pressing questions by choosing the GEM and GC models to represent the two forecasting approaches.
-First, by comparing and contrasting the GEM and GC models, their relative strengths and weaknesses are identified.
-Skill verification against observations and GDPS analyses reveals that GC suffers from excessive fine-scale smoothing, affecting scales as large as 2500 km beyond day 3 of a 10-day forecast.
-However, despite smoothing, large-scale features (wavelength$>$2500 km) in GC exhibit better skill compared to GEM, which is evidenced by increased spectral coherence relative to the GDPS analyses and reduced STDE against radiosonde observations.
-Inspired by the improved large-scale skill of GC, a hybrid NWP-AI system – namely, GDPS-SN – has been developed to produce global forecasts where GEM-predicted large scales in horizontal wind and temperature are spectrally nudged toward inferences from GC.
-This hybrid system is capable of generating real-time forecasts with accuracy that surpasses ECCC’s operational GPDS.
-For example, RMSE of the 500-hPa geopotential height is reduced by 5-10%, with an overall predictability improvement of 12-18 hours, peaking at day 7 of a 10-day forecast.
-Notable improvements are also obtained in the predicted precipitation and tropical cyclone trajectories.
-Furthermore, predictions from GDPS-SN are expected to be more physically consistent than GC as, at its core, GDPS-SN employs an NWP model based on the first principles.
-In addition to enhancing the skill and predictability of GEM, the proposed hybrid NWP-AI system is shown to generate the full spectra of fine scales that are missing from GC inferences.
-Comparisons against near-surface observations have revealed that – unlike GC, which has a limited ability to capture the extremes in wind speed and temperature – weather extremes predicted by GDPS-SN closely follow the operational GDPS-CTL.
-Operational NWP models are generally required to track hundreds of prognostic variables while generating forecasts with high vertical resolution and temporal frequency.
-These capabilities are essential for meteorologists to forecast weather patterns, particularly those relevant for high-impact events.
-The proposed hybrid system maintains these necessary capabilities, albeit with a reasonable increase in computational cost.
-In this regard, it is important to recognize that similar capability requirements would substantially diminish the computational cost advantage of the AI models.
-Additionally, training such AI models would be prohibitively expensive.
-Work is currently in progress at ECCC to fine-tune GC to emulate the operational GDPS analyses.
-This fine-tuned version is expected to enhance GC’s skill with the 37-level version, particularly in the stratosphere and the boundary layer.
-Any such improvement could be easily leveraged by the hybrid system in the near future.
-Importantly, a fine-tuned 37-level version of GC would allow for nudging in the boundary layer which is expected to result in further enhancement of near-surface skill with GDPS-SN, which with the current optimal configuration, shows less remarkable improvements over GDPS-CTL.
-In the future, GDPS-SN could be utilized as the nudging reference for other downstream higher-resolution deterministic GEM-based operational NWP systems.
-Additionally, spectral nudging could be investigated for global ensemble prediction using a probabilistic AI model, such as GenCast (Price et al., 2024).
-Finally, results from this study imply that, rather than viewing AI and NWP models as two competing paradigms, it is likely more prudent to consider these forecasting methods as complementary.
+The emergence of AI-based weather prediction models has disrupted the operational paradigm long dominated by the physics-based systems.
+The pressing question that motivates this study is whether it is possible to improve guidance from operational NWP models by leveraging the predictive skill of AI inferences.
+A careful comparison of physics-based GEM and AI-based GraphCast predictions reveals that the latter suffers from excessive smoothing up to synoptic scales.
+At larger scales, however, GraphCast predictions are found to be highly skillful.
+Inspired by the improved large-scale skill of GraphCast, a hybrid NWP-AI system – namely, GDPS-SN – has been developed to produce global forecasts in which GEM’s large-scale state is spectrally nudged toward larges scales of GraphCast inferences.
+This hybrid system is capable of generating real-time forecasts with accuracy that significantly surpasses ECCC’s operational GPDS.
+The RMSE of the 500-hPa geopotential height is reduced by 5-10%, with the largest predictability improvements attained around day 7 that exceed 24 hours over the tropics.
+This accurate prediction of large-scale circulation improves tropical cyclone steering estimates and the associated track predictions.
+Although these results are achievable directly with AI-based systems, the hybrid model also generates the full spectra of fine scales that represents the tails of the climatology.
+This allows GDPS-SN to predict weather extremes that are challenging for the current generation of purely “data-driven” models.
+Operational NWP models generate hundreds of internally and physically consistent forecast fields at high vertical resolution and temporal frequency.
+These outputs serve as essential guidance for operational meteorologists, especially for forecasts of high-impact events.
+The proposed hybrid system maintains these necessary capabilities, with a reasonable increase in computational cost.
+Increasing global AI model resolution and expanding the number of predicted variables represent real computational challenges for both training and inference.
+Moreover, the procedures needed to create data-driven predictions of the multitude of unanalyzed variables in a physically consistent way has not yet been devised.
+In this regard, the proposed hybrid system compensates for the perceived weaknesses of NWP models while addressing the limitations of current AI models.
+The first version of GDPS-SN became operational (with an experimental status) at ECCC in March 2025.
+It is based on a GraphCast version that has been retrained and fine-tuned at ECCC.
+Work is also in progress to fine-tune GraphCast to emulate operational GDPS analyses.
+This fine-tuned version is expected to enhance GraphCast’s skill with the 37 pressure-level version, particularly in the stratosphere and the boundary layer.
+Any such improvement will quickly be integrated into the hybrid GDPS-SN system.
+Although this study used the GEM and GraphCast models for hybridization, the protocol established here could be applied to any pair of systems, provided that the physical model supports spectral nudging.
+Hybrid systems like the one described here may represent an optimal blending of the individual advantages of the physics- and data-based approaches to weather prediction in the foreseeable future.
+The results from this study imply that, rather than viewing AI and NWP models as two competing paradigms, it is likely more prudent to consider these forecasting methods as complementary.
 A well-designed fusion of these two approaches can significantly mitigate their individual limitations while allowing for harnessing their respective strengths to provide better meteorological guidance.
@@
 […]
+$A_{x}$ $=\sqrt{S_{x^{\prime},x^{\prime}}},$ (6)
+$A_{y}$ $=\sqrt{S_{y^{\prime},y^{\prime}}},$ (7)
+where, given any two 2-D scalar fields $u$ and $v$, the operator $S$ is defined as
+$S_{u,v}\equiv E\bigg{[}\Big{\langle}{(u-\langle{u}\rangle)(v-\langle{v}\rangle)}\Big{\rangle}\bigg{]}.$ (8)
+Here, the angle brackets denote spatial averaging, including latitudinal weighting.
+The activity ratio of the forecast, normalized by the analysis activity, can then be obtained as follows:
+$\Gamma=\sqrt{\frac{S_{x^{\prime},x^{\prime}}}{S_{y^{\prime},y^{\prime}}}},$ (9)
+and the anomaly correlation coefficient is given as
+$P=\frac{S_{x^{\prime},y^{\prime}}}{\sqrt{S_{x^{\prime},x^{\prime}}S_{y^{\prime},y^{\prime}}}}.$ (10)
+These two coefficients quantify, respectively, the amplitude and phase errors of forecast anomalies with respect to the climatological mean.
+By expressing the forecast and analysis fields as truncated spherical harmonic expansions, the covariance can be decomposed into contributions from individual harmonic components.
+Applying Parseval’s identity and grouping terms by zonal wavenumber, $S$ can be written as a sum over total spherical wavenumbers $n$ as follows:
+$S_{u,v}=\sum_{n=1}^{N_{trunc}}E\big{[}\sigma_{u,v}(n)\big{]},$ (11)
+where $\sigma_{u,v}(n)$ denotes the cross spectral density between $u$ and $v$ at wavenumber $n$.
+When $u=v$, $\sigma_{u,u}$ corresponds to the power spectral density of $u$.
+This allows for introducing a spectral activity ratio as a function of wavenumber $n$ as
+$\gamma(n)=\sqrt{\frac{E\big{[}\sigma_{x^{\prime},x^{\prime}}(n)\big{]}}{E\big{[}\sigma_{y^{\prime},y^{\prime}}(n)\big{]}}},$ (12)
+and a spectral anomaly correlation coefficient, hereafter referred to as spectral coherence, as
+$\rho(n)=\frac{E\big{[}\sigma_{x^{\prime},y^{\prime}}(n)\big{]}}{\sqrt{E\big{[}\sigma_{x^{\prime},x^{\prime}}(n)\big{]}\,E\big{[}\sigma_{y^{\prime},y^{\prime}}(n)\big{]}}}.$ (13)
+Estimating model climatology $x_{c}$ requires running hindcasts, which may be impractical outside of an operational NWP context.
+The conventional workaround is to use a common climatology database, setting $x_{c}=y_{c}=c$.
+The results presented in Figs.
+19a–i correspond to this approach, where anomalies are calculated relative to a common ERA-Interim reanalysis climatology $c$ (Dee et al., 2011).
+Alternatively, approximate climatological means $x_{c}$ and $y_{c}$ can be obtained as $x_{c}=\overline{x}$ and $y_{c}=\overline{y}$, where the overbar denotes average over a set of cases for a given season.
+This is the approach adopted throughout this paper.
+Therefore, Eqs.
+1 and 2 are equivalent to Eqs.
+12 and 13, respectively, where expectations are estimated by averaging spectral densities over a set of cases for a given season.
+Both approaches deliver qualitatively similar insights regarding model performance at different scales (not shown).
+Figure 19: Normalized activity (left column), spectral activity ratio (middle column) at day 5, and scale-separated normalized activity ratio (right column) for temperature at 850 hPa (top row) and 250 hPa (middle row), and geopotential at 500 hPa (bottom row) for GDPS-CTL (blue), GDPS-SN (red), and GraphCast (grey).
+Forecast activity is calculated using ECMWF analysis climatology, whereas normalization is done using GDPS analyses.
+Results correspond to boreal winter.
+Figure 19a illustrates the normalized activity ratio for temperature at 850 hPa across different lead times for predictions from GDPS-CTL, GDPS-SN, and GraphCast.
+By day 5, GraphCast’s activity is suppressed by about 4% compared to GDPS analyses.
+GDPS-CTL and GDPS-SN show similar activity, with GDPS-CTL being slightly overactive (up to 2%) beyond day 5.
+The spectral activity ratio at day 5 (Fig.
+19b) shows increasingly suppressed activity for GraphCast at large wavenumbers, starting from wavenumber 16 (approx.
+2500 km).
+For scales below 100 km, activity reduction is nearly 50%.
+However, this significant reduction in fine-scale activity is not reflected in the total activity ratio (Fig.
+19a).
+Separating normalized activity for scales larger and smaller than 1000 km (wavenumber $<$40 and $>$40) in Fig.
+19c shows that total activity ratio is mainly reflective of scales larger than 1000 km.
+This inference is based on the fact that the total activity ratios (dashed lines in Fig.
+19c) almost match large-scale activity ratios (solid lines in Fig.
+19a), while fine-scale activity ratios (dotted lines in Fig.
+19c) indicate significant suppression in GraphCast.
+Figures 19d–f further demonstrate the limitation of total activity ratio in identifying fine-scale smoothing effects.
+For temperature at 250 hPa, GraphCast’s total activity is the highest among models (Fig.
+19d) despite a 60% reduction in activity for scales below 500 km (Fig.
+19e).
+It turns out that the total activity metric rewards GraphCast for over-predicting activity at the planetary scales (wavenumber$<$4) and is insensitive to an overall 30% deficiency in activity for scales smaller than 1000 km (Fig.
+19f).
+Figures 19g–-i illustrate, for 500 hPa geopotential height, that activity as a metric can also be insensitive to spurious fine-scale variance.
+Specifically, both total activity (Fig.
+19g) and fine-scale activity (Fig.
+19i) for GraphCast fail to capture the unphysical increase in fine-scale variance concentrated near wavenumber 400 (Fig.
+19h)—a known issue with GraphCast, as previously reported by Lam et al.
+(2023).
+Overall, these results highlight the limitations of total activity for assessing mesoscale issues in model predictions (smoothing or spurious variance) and underscores the importance of its spectral decomposition for proper assessment of model performance.
+Finally, for all cases shown in Fig.
+19, GDPS-SN’s fine-scale activity is close to GDPS-CTL, exhibiting only slight suppression depending on the specific variable and level.
+This behavior is consistent with the spectral activity and amplitude ratios shown in Figs.
+19 and 7, respectively.
+[…]
 $f_{n}=\exp\left[-\left(\frac{n(n+1)}{n_{o}(n_{o}+1)}\right)^{r\;}\right],$ (14)
-where $n$ denotes the total wavenumber, $n_{o}$ is the cut-off total wavenumber and the exponent $r$ is a non-dimensional parameter related to the sharpness of the filter response.
-Based on the spectral comparison of the GDPS and GC predictions, the filter is configured by setting $n_{0}=30$ and $r=4$.
+where $n$ denotes the total wavenumber, $n_{o}$ is the cut-off total wavenumber, and the exponent $r$ is a non-dimensional parameter related to the sharpness of the filter response.
+Based on the spectral comparison of the GDPS and GraphCast predictions, the filter is configured by setting $n_{0}=30$ and $r=4$.
 As shown in Fig.
-18, the resulting filtered fields fully retain amplitudes for scales associated with approximately $n<20$ (wavelength$>$2000 km) and fully removes scales corresponding to $n>40$ (wavelength$<$1000 km).
+20, the resulting filtered fields fully retain amplitudes for scales associated with approximately $n<20$ (wavelength$>$2000 km) and fully removes scales corresponding to $n>40$ (wavelength$<$1000 km).
 The physical fields reconstructed from the filtered spectra leads to the desired filtered fields.
@@
 where $\Delta$ denotes the model grid spacing, $\alpha$ is the normalized two-dimensional wavenumber given by $\alpha=\sqrt{\frac{m^{2}}{N_{i}^{2}}+\frac{n^{2}}{N_{j}^{2}}}$ associated with each of the Yin-Yang sub-domains of size $(N_{i}\times N_{j})$, and $\hat{\alpha}=\alpha\lambda_{LS}/(2\Delta)$.
-Figures 19a and 19b illustrate the response of this DCT-based filter for the optimal configuration based on $\lambda_{LS}=2750$ km and $\lambda_{SS}=2250$ km.
+Figures 21a and 21b illustrate the response of this DCT-based filter for the optimal configuration based on $\lambda_{LS}=2750$ km and $\lambda_{SS}=2250$ km.
 The figures demonstrate that the DCT-based filter is capable of targeting the desired scales for the individual Yin-Yang sub-domains.
-The response of the filter over the globe was determined by inspecting the spectral variance ratio of the filtered and unfiltered fields over a global Gaussian grid with respect to a spherical harmonics-based decomposition.
-The corresponding results are presented in Fig.
-19c.
-The figure reveals that, even though DCT is primarily designed for LAMs where curvature effects of spherical surfaces are neglected, its application on global Yin-Yang grid does not result in any major issue as the filter is found to result in an acceptable response.
-Figure 19: Response of the DCT-based filter (see Eq. 7) with respect to the (a) Yin grid, (b) Yang grid, and (c) a global Gaussian grid.
+The global response of the filter was determined by first interpolating a field (e.g., temperature) from the Yin-Yang grid to a global Gaussian grid before and after applying the DCT-based filter.
+The spectral variance ratio between the filtered and unfiltered fields was then computed using a spherical harmonics-based decomposition.
+The corresponding results for temperature at 500 hPa are presented in Fig.
+21c for illustration.
+Figure 21: Response of the DCT-based filter based on $\lambda_{LS}=2750$ km and $\lambda_{SS}=2250$ km (see Eq.
+15) with respect to the (a) Yin grid, (b) Yang grid, and (c) a global Gaussian grid for temperature at 500 hPa.
+Over the global Gaussian grid, the small- and large-scale cut-offs for the optimal configuration are approximately 2000 km and 5500 km, respectively, as indicated with the shaded area.